### v2

# A Weighted Loss Approach to Robust Federated Learning under Data Heterogeneity

Abstract
Federated learning (FL) is a machine learning paradigm that enables multiple data holders to collaboratively train a machine learning model without sharing their training data with external parties. In this paradigm, workers locally update a model and share with a central server their updated gradients (or model parameters). While FL seems appealing from a privacy perspective, it opens a number of threats from a security perspective as (Byzantine) participants can contribute poisonous gradients (or model parameters) harming model convergence. Byzantine-resilient FL addresses this issue by ensuring that the training proceeds as if Byzantine participants were absent. Towards this purpose, common strategies ignore outlier gradients during model aggregation, assuming that Byzantine gradients deviate more from honest gradients than honest gradients do from each other. However, in heterogeneous settings, honest gradients may differ significantly, making it difficult to distinguish honest outliers from Byzantine ones. In this paper, we introduce the Worker Label Alignement Loss (WoLA), a weighted loss that aligns honest worker gradients despite data heterogeneity, which facilitates the identification of Byzantines’ gradients. This approach significantly outperforms state-of-the-art methods in heterogeneous settings. In this paper, we provide both theoretical insights and empirical evidence of its effectiveness.
1 Introduction
Federated learning (FL) [19, 18, 41] enables a group of machines (workers) to collaboratively train a model without sharing their private data. In this paradigm, a central server coordinates the process by sending the current model to all workers. Each worker then performs a training step on its local data and sends an update back to the server. Typically, this update is the gradient computed from a mini-batch [19, 3]. The server aggregates the gradients, updates the model, and sends it back to the workers, repeating the cycle.
In an ideal setup, the server aggregates gradients by simply averaging them. However, some workers may fail or behave maliciously by sending corrupted updates (also referred to as Byzantine behaviour). Robust training seeks to achieve model performance comparable to that achieved in the absence of Byzantine workers. To this end, a common strategy is to replace simple averaging with robust aggregation methods on the server side [19, 18]. These methods are designed to reduce the impact of outliers, which are expected to include Byzantine updates. However, their effectiveness strongly depends on the assumption that honest workers hold homogeneous data distributions [18, 3, 33, 4]. In practice, this assumption may not hold [3, 24, 29, 15], as each worker typically has access to only a small, potentially unrepresentative portion of the dataset.
With heterogeneous data, the divergence among honest gradients increases, making it easier for Byzantine machines to mask their updates [3, 19, 24]. In this setting, a robust aggregator alone is not enough to resist Byzantine attacks. A common approach is to add a pre-aggregation step, such as Bucketing [24] or Nearest Neighbor Mean (NNM) [24, 3, 19, 18]. While effective under weak adversarial conditions (i.e., when the proportion of Byzantine updates is low), we show that the performance of these solutions drops significantly in stronger adversarial scenarios.
We present a novel solution for robust FL in heterogeneous and highly adversarial settings. Specifically, we address label distribution skew [29, 35, 46]—a form of data heterogeneity where label distributions vary across workers, while the feature distribution for each class remains consistent across them. This scenario is common in practice [29, 35, 46]. For instance, in a mobile image recognition app, users of different ages may capture different subjects, resulting in variations in label distributions across devices. In another case, hospitals may specialize in certain types of diseases, leading to an overrepresentation of specific labels in their local patient records.
Label skewness leads to inconsistent training across workers, as each worker optimizes the model based on its own label distribution [24, 46, 4]. Local training tends to overemphasize majority classes while neglecting minority ones [46, 4]. Since these class imbalances vary across workers, local objectives become misaligned, leading to highly inconsistent updates.
Rather than attempting to reduce variance of updates through pre-aggregation, we address the root cause: misaligned worker objectives. We introduce the Worker Label Alignment (WoLA) loss, a weighted loss function that aligns local training objectives. WoLA reweights point-wise errors for each worker to generate gradients as if all local datasets were drawn from the same label distribution. By steering honest gradients in a common direction, WoLA reduces variance and improves Byzantine resilience.
WoLA is model-agnostic, requires no hyperparameter tuning, no server-side dataset, no knowledge of the number of Byzantine machines, and adds no computational cost. It is also straightforward to implement and can be combined with other robustness strategies.
The only requirement is that a common training objective—represented by a probability distribution—is fixed and shared by the server once before training begins. For instance, this objective could be the global label distribution (i.e., the distribution obtained by merging all local datasets), derived either from prior knowledge or from local label distributions shared by the workers with the server. Using this information, the server estimates and broadcasts the global distribution. Despite the fact that Byzantine nodes can lie in their data distributions ahead of training, we show in this paper that their influence remains limited and that this approach still brings a significant gain in robustness compared to state-of-the-art approaches.
Our contributions can be summarized as follows:
  • We introduce WoLA, a weighted loss function designed to address label skewness in highly adversarial settings (Section 5).
  • We provide theoretical insights for the effectiveness of WoLA, with all proofs included in the appendix (Section 6)
  • We empirically show that WoLA significantly outperforms state-of-the-art methods through extensive experiments involving various models, datasets, aggregators, and Byzantine attacks. Our code is publicly available [^1] https://anonymous.4open.science/r/WoLA-E4CB to ensure reproducibility (Section 8).
The paper is organized as follows. Section 2 reviews the background and related work. Section 4 introduces the motivation behind the WoLA. Section 5 formally defines the loss and explores its properties. Section 6 presents theoretical justifications for its design. Section 7 describes the experimental setup, and Section 8 presents the experimental results. Finally, Section 9 concludes the paper and outlines directions for future work.
2 Background
This section outlines key FL concepts and defines our threat model. All notations are summarized in Table 1.
Table 1: Notation—In (c) and (d), superscript indices denote classes, while subscripts denote workers.
Notation Meaning
$[k]$ Set of integers from $1$ to $k$, with $k\in\mathbb{N}$
##\## Cardinality of a set
$\mathbbm{1}$ Indicator function. For example, $\mathbbm{1}(x\leq 0)=1$ if $x\leq 0$, and 00 otherwise.
$\nabla$ Gradient
[…]
$\mathbb{E}[X]$ Expected value of the random variable $X$
$\mathbb{E}[X|A]$ Conditional expectation of the random variable $X$ given the event $A$
i.i.d. Independent and identically distributed
a.s. Almost surely is a probabilistic concept meaning that an event happens with probability $1$.
[…]
Notation Meaning
$\Phi$ Model
$n$ Number of workers
$\mathcal{H}$ Set of honest workers, with $\mathcal{H}\subset[n]$
$f$ Number of Byzantine workers, with $f=n-\#\mathcal{H}$
Notation Meaning
$C$ Number of classes
$\mathcal{D}$ Global training set, composed of pairs $(x,y)\in\mathcal{X}\times[C]$, where $x$ is an observation, and $y$ its corresponding label.
$\mathcal{D}_{i}$ Local dataset of worker $i$, with $\mathcal{D}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$
$N$ Size of the global dataset: $N=\#\mathcal{D}$
$N_{i}$ Size of worker $i$’s local dataset: $N_{i}=\#\mathcal{D}_{i}$
$N^{c}$ Number of class-$c$ instances in the global dataset, with $N=\sum_{c=1}^{C}N^{c}$
$N_{i}^{c}$ Number of class-$c$ instances in worker $i$’s dataset, with $N_{i}=\sum_{c=1}^{C}N_{i}^{c}$
$p$ Label distribution of $\mathcal{D}$: $p=(p^{1},p^{2},\ldots,p^{C})$ with $p^{c}=N^{c}/N$
$p_{i}$ Label distribution of $\mathcal{D}_{i}$: $p_{i}=(p_{i}^{1},p_{i}^{2},\ldots,p_{i}^{C})$ with $p^{c}_{i}=N_{i}^{c}/N_{i}$
Notation Meaning
$\mathcal{L}$ Loss
$\ell(y,\Phi(x))$ Pointwise loss of prediction $\Phi(x)$ and label $y$
$\mathcal{L}_{i}$ Loss of worker $i$: $\mathcal{L}_{i}=\sum_{(x,y)\in\mathcal{D}_{i}}\ell\Big{(}y,\Phi(x)\Big{)}/N_{i}$, with $\sum_{i\in\mathcal{H}}\mathcal{L}_{i}\ N_{i}/N=\mathcal{L}$
$\mu^{c}$ Class-wise gradient of class $c$ in $\mathcal{D}$, i.e., $\sum_{(x,y)\in\mathcal{D}}\mathbbm{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}$
$\mu^{c}_{i}$ Class-wise gradient of class $c$ in $\mathcal{D}_{i}$, i.e., $\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}_{i}$
$\mathcal{WL}$ Worker Label Alignement Loss (WoLA)
$\mathcal{WL}_{i}$ WoLA of worker $i$, with $\sum_{i\in\mathcal{H}}\mathcal{WL}_{i}\ N_{i}/N=\mathcal{WL}$
$q$ Shared training objective across workers: a probability distribution $q=(q^{1},q^{2},\ldots,q^{C})$, where $q^{i}\in[0,1]$ and $\sum_{i\in[C]}q^{i}=1$.
2.1 Byzantine-free FL
We consider a distributed architecture with $n$ workers and a central server, aimed at solving a classification task with $C$ classes. The workers hold local datasets $\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{n}$. Each worker $i$ holds $N_{i}$ data points $(x,y)\in\mathcal{X}\times[C]$, where $\mathcal{X}$ is the input space, $x$ is an observation, and $y$ is its label. Given a model $\Phi$, the standard local loss function (not WoLA) for worker $i$ is:
$\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\ell\big{(}y,\Phi(x)\big{)},$
where $\ell$ is a point-wise loss function, assumed to be differentiable with respect to the model parameters.
In each communication round, the server broadcasts the model parameters to all workers. Each worker $i$ is expected to compute the gradient using its own data $\mathcal{D}_{i}$ and send back the resulting gradient $\nabla\mathcal{L}_{i}$.
2.2 Threat Model
In this paper, we focus on handling Byzantine workers under the common assumption of a trusted FL server [3, 24, 4]. We consider an adversarial setting where fewer than half the workers are Byzantine, i.e., $f<n/2$. These malicious workers can send arbitrary updates to the server and have full knowledge of all local updates, following the threat model in [3, 24, 4]. The server’s objective is to identify a stationary point of the average loss over the honest workers. That is, to find a point where $\nabla\mathcal{L}=0$, with
$\mathcal{L}:=\sum_{i\in\mathcal{H}}\mathcal{L}_{i}\frac{N_{i}}{N},$
and $N=\sum_{i\in\mathcal{H}}N_{i}$. Assuming such a point exists, it can be approximated via gradient descent and robustness strategies.
2.3 Global Dataset, Local Datasets & Label Distribution Skew
To model data heterogeneity, we consider the following label skew scenario: all data points are drawn from independent and identically distributed (i.i.d) random variables, forming the global dataset $\mathcal{D}$. These points are then randomly assigned to honest workers, resulting in local datasets $\mathcal{D}_{i}$ for $i\in\mathcal{H}$, such that $\mathcal{D}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$. Each worker receives data with a different label distribution, introducing heterogeneity. Although Byzantine workers may also have their own data, we exclude them from the global dataset as they play no role in our analysis.
The global label distribution refers to the distribution of labels in the full dataset $\mathcal{D}$, while the local label distributions refer to label distributions in a worker’s datasets $\mathcal{D}_{i}$, for $i\in\mathcal{H}$.
2.4 Gradient Dissimilarity
Karimireddy et al. [24] introduce Gradient Dissimilarity as a measure of system robustness, lower gradient dissimilarity implies stronger robustness. It quantifies how much individual gradients deviate from the average of the honest gradients:
Gradient $\text{Gradient Dissimilarity}=\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\|\nabla\mathcal{L}_{i}-\nabla\mathcal{L}\|^{2},$
Data heterogeneity tends to increase gradient dissimilarity [19, 18, 24].
3 Related Work
This section reviews state-of-the-art approaches relevant to our context.
3.1 Robustness Strategies with Homogeneous Data
To mitigate Byzantine behavior during aggregation, a common approach is to replace simple updates averaging with robust aggregation methods [34, 19]. Examples include the coordinate-wise median (CWMed) [45], geometric median (GM) [38, 34], coordinate-wise trimmed mean (CwTM) [45], Krum [7], Multi-Krum (MKrum) [20], Minimum Diameter Averaging [20], and Comparative Gradient Elimination [22]. These methods heavily rely on the assumption that honest workers have homogeneous data. Their robustness under data heterogeneity remains unclear [18, 3, 33].
Besides robust aggregators, additional strategies can enhance robustness in homogeneous settings. These approaches rely on a clean dataset held by the server to evaluate and filter worker updates [44, 9, 14, 36].
3.2 Robustness Strategies Under Data Heterogeneity
We begin by defining Gradient Dissimilarity, a robustness measure commonly used in studies addressing Byzantine resilience under data heterogeneity [19, 18, 24]. Next, we review state-of-the-art approaches in Byzantine FL, followed by methods developed for the Byzantine-free setting. Finally, we highlight the advantages of our proposed method.
3.2.1 Byzantine FL
A common strategy to reduce gradient variance is to apply a pre-aggregation step, such as Bucketing [24], Nearest-Neighbor Mixing (NNM) [3], or FoundationFL [15]. These methods act before the main aggregation rule, resulting in a two-step aggregation process—for example, applying NNM before Krum. Other methods like BOBA [4] and RAGE [12] use different strategies and do not rely on pre-aggregation.
Bucketing [24] randomly partitions inputs into groups (buckets), averages the inputs in each bucket, and feeds these averages into the aggregation rule. This random partitioning reduces the empirical variance of honest updates in expectation (i.e., on average). However, in some iterations, it may fail to reduce heterogeneity, giving Byzantine workers a better chance to disrupt learning [18].
In contrast, NNM [3] is a deterministic pre-aggregation method that replaces each gradient with the average of its $n-f$ nearest neighbors (including itself). NNM introduces additional computational cost, which prevents its use in large-scale distributed learning tasks [18].
FoundationFL [15] synthesizes additional updates at each training step. The server assigns a score to each update and replicates the highest-scoring gradient $m$ times, where $m$ is a hyperparameter set by the authors to $m=n/2$. A critical limitation of FoundationFL is its underlying assumption [^2] The score is based on the distance to the coordinate-wise minimum and maximum. Given updates $g_{1},\dots,g_{n}$, define $g_{\text{max}}=\max(g_{1},\dots,g_{n})$ and $g_{\text{min}}=\min(g_{1},\dots,g_{n})$ (coordinate-wise). The score $s_{i}$ for worker $i$ is then $s_{i}=\min(\|g_{\text{max}}-g_{i}\|,\|g_{\text{min}}-g_{i}\|)$. A high score means the update is far from the extremes.: extreme values correspond to Byzantine updates. However, this assumption is not reliable under data heterogeneity, particularly against the A Little Is Enough (ALIE) attack [5].
BOBA [4] differs from standard robustness strategies. It projects client gradients onto a $C-1$ dimensional subspace using truncated Singular Value Decomposition (SVD), estimates the $C-1$ simplex from a server-side dataset, reconstructs each client’s label distribution based on their position relative to the simplex, and discards those with abnormal values. In addition to requiring server-side data, at each round, BOBA incurs extra cost from computing the truncated SVD of the gradient matrix, which scales with both the number of clients and the gradient dimension [^3] The computational complexity of BOBA is $\mathcal{O}(kCnd)$, where $k$ is the number of truncated SVDs computed per round (typically between 3 and 5 in thei experiments), $C$ is the number of classes, $n$ the number of workers, and $d$ the gradient dimension..
RAGE [12] performs a weighted average using confidence scores assigned to gradients. These scores are computed iteratively, downweighting gradients with large projections on the principal eigenvector of the empirical covariance matrix of gradients. The covariance is recomputed with updated scores until the largest eigenvalue is sufficiently small. The computational complexity of RAGE makes it impractical for large-scale optimization [24].
3.2.2 Byzantine-free FL
In a Byzantine-free setting, non-IID data remains a major obstacle to model convergence. Several techniques have been proposed to address this challenge, including knowledge distillation [^4] Knowledge distillation encourages a student model to approximate the teacher model’s output logits, enabling it to mimic the teacher’s behavior. and logit calibration.
Some methods rely on knowledge distillation and require access to an additional test set [28, 11, 10, 30]. In contrast, [47] and [39] avoid relying on proxy datasets. They use a generator trained on the server or client to enable data-free knowledge distillation. However, these approaches require significant communication and coordination, and introduce substantial computational overhead [40, 21].
Other approaches introduce personalized loss functions based on local data distributions, including DMFL [35], FedLC [46], and FedDisk [32]. DMFL and FedLC, inspired by [8], adjust logits [^5] Logits are the outputs of the final classification layer and inputs to the softmax. before applying softmax cross-entropy, increasing margins for minority classes. These methods require the tuning of a hyperparameter with no bounded range. FedDisk uses loss with sample-level weights, learned from a global and multiple local density estimators, independent of class labels.
All these propositions either require an additional tes set, introduce significant overhead [40, 21, 32], or rely on hyperparameter tuning—a particularly difficult task in our setting: (i) without a test set, local tuning is possible, but data heterogeneity may cause inconsistent values; and (ii) since these methods assume a Byzantine-free setting, no robust protocol is provided to handle malicious behavior.
3.2.3 Our proposition
To the best of our knowledge, using a loss function to improve Byzantine resilience under data heterogeneity is a novel approach. WoLA is a model-agnostic method that requires no hyperparameters, no server-side dataset, and no knowledge of the number of Byzantine machines. It adds no computational cost, is easy to implement, and can be combined with other robustness strategies. Its only requirement is a shared training objective.
While many approaches focus on handling update heterogeneity [3, 24, 4, 15, 12], WoLA tackles its root cause—label skewness—by aligning gradients from the start. This enables WoLA to significantly outperform state-of-the-art methods. In the next section, we explain the motivation behind its design.
4 Rationale for Worker Label Alignement Loss
This section first shows that local gradients approximate weighted averages of class-wise gradients. Next, it illustrates that class-wise gradients can differ significantly. Finally, it presents the main conclusion drawn from these observations.
4.1 Gradients as Weighted Averages
The key rationale behind WoLA is that the gradient over a sufficiently large local dataset can be approximated as a weighted average of class-wise gradients, with weights given by the local label distribution.
More precisely, the gradient of worker $i$ for a standard loss is:
$\nabla\mathcal{L}_{i}$ $=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\sum_{c=1}^{C}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}$
$=\sum_{c=1}^{C}\frac{N^{c}_{i}}{N_{i}}\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}_{i}},$
where:
$\frac{N^{c}_{i}}{N_{i}}=\Bigg{(}\text{Proportion of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}\vcentcolon=p^{c}_{i},$
and
$\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}_{i}}$ (1)
$\quad\quad=\Bigg{(}\text{Mean gradient of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}\vcentcolon=\mu_{i}^{c}$
Given a sufficient number of instances of class $c\in[C]$, the following approximation is reasonable (see Appendix Section A.2, Lemma 2 for a formal statement):
(Mean gradient of class $\Bigg{(}\text{Mean gradient of class }c\text{ in }\mathcal{D}_{i}\Bigg{)}$ (2)
$\quad\quad\approx\Bigg{(}\text{Mean gradient of class }c\text{ in}\mathcal{D}\Bigg{)}$
$\quad\quad=\frac{\sum_{(x,y)\in\mathcal{D}}\mathbbm{1}(y=c)\nabla\ell\Big{(}c,\Phi(x)\Big{)}}{N^{c}}\vcentcolon=\mu_{c}$
We call $\mu_{i}^{c}$ the class-$c$ gradient of worker $i$, and $\mu^{c}$ the global class-$c$ gradient. When clear from context, we omit the term global.
Using approximation equation (2), we can express the local gradient as a weighted average of global class gradients:
$\nabla\mathcal{L}_{i}\approx\sum_{c=1}^{C}p_{i}^{c}\mu^{c}$ (3)
Assuming that the vectors $\mu^{c}$ for $c\in[C]$ differ—either in norm or direction—the differences in local gradients arise from the local label distributions, through the weights $p_{i}^{c}$.
In the next subsection, we show that the vectors $\mu^{c}$ for $c\in[C]$ are indeed different.
4.2 Difference in Class Gradients
Figure 1 shows standard (non-FL) training of a shallow neural network on the Iris dataset [17]. The loss decreases and accuracy improves over epochs, indicating successful model convergence. Independently of the optimization, at each step we compute the class gradients for Setosa, Versicolour, and Virginica, and track the cosine similarity between these gradients over training steps.
The figure reveals that the angles between class-wise gradients often exceed $\pi/2$, indicating that the gradients point in widely different directions. In some cases, such as between Versicolour and Virginica, the gradients are nearly opposite.
This highlights that different classes can push the model in distinct—and sometimes conflicting—directions during training. Similar observations have been reported in previous studies [46, 4].
[…]
4.3 Key Takeaways
Local gradients approximate weighted averages of global class gradients, with weights given by the local label distribution. Since class gradients may differ, aligning local gradients requires aligning their weightings—as if local datasets shared the same label distribution. WoLA is designed to achieve this goal.
5 Worker Label Alignement Loss
This section introduces WoLA. It first defines the training objective required for WoLA, then presents its formula. The final subsections provide an intuitive understanding of WoLA’s properties. These properties are rigorously stated in Section 6 and proved in the Appendix A.
5.1 Shared Training Objective
WoLA requires a shared training objective across workers. This objective is formalised by a probability distribution $q=(q^{1},q^{2},\ldots,q^{C})$, where $q^{i}\in[0,1]$ and $\sum_{i\in[C]}q^{i}=1$. A larger $q^{i}$ means that point-wise errors for samples with label $i$ are given more weight during training, and vice versa. To define this distribution, we considered three situations:
Prior knowledge
The global label distribution is known in advance, based on expert knowledge or historical data. For example, in tumor detection on MRI scans or crop type recognition from satellite imagery, national statistics can serve as a reliable proxy for label distributions.
Local label distributions are shared
Each worker shares its local label distribution with the server before training. The server then computes the global distribution. In this case, it is important to note that (i) the process reduces data privacy, and (ii) it introduces a new opportunity for Byzantine workers to attack the system. We carefully evaluate this attack scenario in Section 7.2. Regarding privacy concerns, we assume in this paper that the server is trusted as we consider privacy as an orthogonal problem that is actively being investigated in the literature. Nevertheless, off-the-shelf SecureAggregation protocols such as [6] could be easily used to release the trusted server assumption.
Consensus among workers on the training objective
All workers agree that the global model should favor certain classes during training, regardless of the actual global label distribution. For example, in a task to detect hateful or illegal content on social networks—where hate content is rare—it may be agreed to assign higher importance to these cases. In medical diagnosis, rare diseases may also be given more weight. In contrast, for automated recruitment systems, all classes may be weighted equally to avoid algorithmic bias.
The training objective $q$ is shared by the server with all workers once before training begins, allowing them to adjust their local loss functions accordingly.
5.2 WoLA definition
WoLA locally adjusts the weights of point-wise errors to simulate an identical label distribution across workers. More precisely, WoLA for worker $i$, denoted $\mathcal{WL}_{i}$, is defined as:
$\mathcal{WL}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\ell\Big{(}y,\Phi(x)\Big{)},$
While the formula is undefined for $p^{y}_{i}=0$, it is not problematic, as it indicates that label $y$ is absent from the local dataset $\mathcal{D}_{i}$, meaning the corresponding weight never appears in the sum.
A sample from class $c$ is overweighted when class $c$ is underrepresented locally compared to the training objective, i.e., when $p_{i}^{c}<q^{c}\Leftrightarrow 1<q^{c}/p_{i}^{c}$; and underweighted otherwise.
5.3 Local Updates Alignement
WoLA makes local datasets appear as if they had identical label distributions matching the training objective $q$—that is, as if $p_{i}^{c}=q^{c}$ for all $c\in[C]$ and all $i\in\mathcal{H}$. Specifically, for worker $i$, the gradient under WoLA is:
$\nabla\mathcal{WL}_{i}\vcentcolon=$ $\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\sum_{c=1}^{C}\mathbbm{1}(y=c)\frac{q^{c}}{p^{c}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=$ $\sum_{c=1}^{C}q^{c}\frac{\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell\Big{(}y,\Phi(x)\Big{)}}{N^{c}_{i}},$
using approximation in equation (2), it follows:
$\nabla\mathcal{WL}_{i}\approx\sum_{c=1}^{C}q^{c}\mu^{c}$ (4)
With WoLA, the approximation of the worker $i$’s gradient no longer depends on local label proportions. It becomes a weighted average of class gradients, using the shared training objective $q$ as weights. Since equation (4) holds for all $i\in[C]$, it follows that $\nabla\mathcal{WL}_{i}\approx\nabla\mathcal{WL}_{j}$ for all $j\in[C]$. This suggests that local updates are more similar under WoLA, potentially reducing gradient dissimilarity and improving robustness.
Figure 2 illustrates the impact of WoLA on updates using a fictitious example with three classes: Dog, Flower, and Frog, under the training objective $q=(1/3,1/3,1/3)$. It shows three local datasets with label distributions $(3/6,2/6,1/6)$, $(2/6,2/6,2/6)$, and $(1/6,1/6,4/6)$. Two of them differ from the training objective $q$. WoLA addresses this mismatch by upweighting samples from local minority classes and downweighting those from local majority classes, thereby aligning updates.
(a) Point-wise standard loss gradients on $\mathcal{D}_{1}$
(b) Point-wise standard loss gradients on $\mathcal{D}_{2}$
(c) Point-wise standard loss gradients on $\mathcal{D}_{3}$
(d) Point-wise WoLA gradients on $\mathcal{D}_{1}$
(e) Point-wise WoLA gradients on $\mathcal{D}_{2}$
(f) Point-wise WoLA gradients on $\mathcal{D}_{3}$
Figure 2: Fictitious FL classification task with three classes (Dog, Flower, Frog) under the training objective $q=(1/3,1/3,1/3)$, comparing standard loss with WoLA. Each image represents a data point from a local dataset. Arrows indicate 2D point-wise loss gradients; in black squares, arrows show the resulting updates. Subfigures (a–c) display gradients under the standard loss, while (d–f) show gradients with WoLA. In (d) and (f), dataset sizes are scaled to reflect the effect of WoLA weights, although datasets are actually unchanged. For example, $\mathcal{D}_{3}$ has a local label distribution of $(1/6,1/6,4/6)$, while $q$ is uniform. This results in weights: $1/3\div 1/6=4/2$ for Dog and Flower, and $1/3\div 4/6=1/2$ for Frog. To avoid fractional images, the denominator 2 is omitted, giving 4 duplicates of Dog and Flower gradients, and 1 of Frog. The same logic applies to $\mathcal{D}_{1}$. Subfigures (b) and (e) are identical, as $\mathcal{D}_{2}$ label distribution matches the training objective. Local updates with WoLA are more aligned than with the standard loss.
5.4 Targeted Model
Assuming that the training objective $q$ matches the global label distribution $p$, i.e., $p=q$, the average gradients of the standard loss and the WoLA loss—denoted $\nabla\mathcal{L}$ and $\nabla\mathcal{WL}$ respectively—are similar. Using the approximations from equations (3) and (4), we have:
$\nabla\mathcal{L}$ $=\sum_{i\in\mathcal{H}}\nabla\mathcal{L}i\frac{N_{i}}{N}$
$\approx\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}p_{i}^{c}\mu^{c}\frac{N_{i}}{N}=\sum_{c=1}^{C}p^{c}\mu^{c}=\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}p^{c}\mu^{c}\frac{N_{i}}{N}$
$=\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}q^{c}\mu^{c}\frac{N_{i}}{N}\approx\sum_{i\in\mathcal{H}}\nabla\mathcal{WL}_{i}\frac{N_{i}}{N}=\nabla\mathcal{WL}$
Therefore, when $p=q$, the training targets the same model regardless of whether the standard or WoLA loss is used.
5.5 Key Takeaways
WoLA requires a shared training objective. Previous analyses suggest that: (i) using WoLA reduces gradient dissimilarity, and (ii) when the training objective matches the label distribution, the model targeted with WoLA is the same as the one with the standard loss.
6 Theoretical Arguments
This section uses probabilistic tools to rigorously state the properties of WoLA. We first describe the data modeling framework, then present asymptotic results that offer insights into WoLA’s behavior.
6.1 Data Modeling
Let $(X_{1},Y_{1},Z_{1}),(X_{2},Y_{2},Z_{2}),\ldots,(X_{N},Y_{N},Z_{N})$ be independent and identically distributed (i.i.d.) random triplets drawn from a distribution $\mathfrak{D}$ over the sample space $\mathcal{X}\times[C]\times\mathcal{H}$. Each pair $(X_{i},Y_{i})$ represents a datapoint in the dataset $\mathcal{D}$. For each $k\in[N]$, the random variable $Z_{k}$ takes values in $\mathcal{H}$ and indicates the worker to which the $k$-th datapoint is assigned. We also denote by $(X,Y,Z)$ a generic triplet with the same distribution $\mathfrak{D}$.
For any triplet $(X_{k},Y_{k},Z_{k})$, we assume that $X_{k}$ and $Z_{k}$ are conditionally independent given $Y_{k}$. That is, conditioned on the label, the features provide no information about the assignment, and vice versa. Formally, $P(X_{k},Z_{k}|Y_{k})=P(X_{k}|Y_{k})P(Z_{k}|Y_{k}).$
6.2 Theoretical Results
The results are asymptotic, considering the case where the size of the global dataset tends to infinity, i.e., $N\to\infty$, and are based on the Law of Large Numbers. Full proofs and details are given in Appendix A.
Let $L_{i}$ and $W_{i}$ be the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{WL}_{i}$, respectively. Let $L$ and $W$ denote the average gradients of honest workers for the standard loss and WoLA, respectively.
With the standard loss, local gradients remain different even with an infinite amount of data:
Proposition 1
Under the standard loss, the gradient of worker $i$ converges almost surely as $N\to\infty$:
[…]
Moreover, the average gradient satisfies:
[…]
In contrast, with WoLA, local gradients become asymptotically identical, regardless of the training objective shared by the server:
Proposition 2
Let $q$ be the training objective distribution shared by the server. Under WoLA, the gradient of worker $i$ converges almost surely as $N\to\infty$:
[…]
Moreover, the average gradient also satisfies:
[…]
This directly implies that, under WoLA, the gradient dissimilarity vanishes asymptotically as $N\to\infty$ (see Appendix A.5).
Finally, if the training objective matches the global label distribution, i.e., $q=p$, then the standard loss gradient and the WoLA gradient become asymptotically identical:
Proposition 3
Let $q$ be the training objective and $p$ the global label distribution. If $q=p$, then:
[…]
This equality implies that, in the asymptotic regime, when the training objective matches the label distribution, the model targeted with WoLA is the same as the one with the standard loss.
(a) Results under ALIE attack.
(b) Results under LF attack.
[…]
(a) Boxplot of accuracy gains based on Table 2.
(b) Boxplot of accuracy gains under the strongest attack, based on Table 3.
[…]
Table 2: Global Performance: Test Accuracy (s.d.) %, averaged over training. Results are averaged across aggregators (CWMed, CwTM, GM, MKrum) and attacks (ALIE, FOE, LF, Mimic, SF). Colors rank values within columns: yellow (1st), blue (2nd), red (3rd). See full table in Appendix.
(a) MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 93.6 (0.3) 92.9 (0.4) 86.1 (1.6) 66.3 (4.5) 93.6 (0.3) 92.2 (0.6) 82.3 (3.3) 43.5 (11.7) 91.9 (1.8) 84.6 (4.0) 58.8 (12.1) 17.3 (7.3)
BKT 93.5 (0.3) 92.8 (0.5) 86.1 (1.6) 66.3 (4.5) 93.6 (0.4) 92.5 (0.7) 82.3 (3.3) 43.5 (11.7) 93.3 (0.7) 89.6 (1.9) 58.8 (12.1) 17.3 (7.3)
FoundFL 89.6 (1.8) 74.4 (4.2) 56.8 (4.6) 47.5 (5.0) 85.4 (5.2) 69.6 (7.3) 57.7 (6.7) 50.4 (6.8) 61.6 (15.9) 50.4 (12.8) 43.0 (12.1) 38.0 (11.8)
NNM 93.7 (0.2) 93.3 (0.3) 92.7 (0.4) 75.5 (6.9) 93.9 (0.3) 93.0 (0.6) 90.1 (2.5) 63.2 (9.0) 93.9 (0.6) 91.5 (1.4) 67.5 (14.5) 35.2 (8.4)
WoLA 93.7 (0.3) 93.4 (0.3) 92.7 (0.3) 81.9 (0.6) 93.9 (0.3) 93.5 (0.4) 92.4 (0.4) 80.5 (2.2) 93.6 (0.5) 93.0 (0.7) 88.9 (2.6) 54.5 (8.0)
WoLA+NNM 93.7 (0.3) 93.5 (0.3) 93.2 (0.3) 84.3 (0.6) 94.0 (0.3) 93.6 (0.4) 93.2 (0.3) 83.1 (2.8) 93.7 (0.6) 93.4 (0.7) 90.0 (3.0) 64.6 (8.7)
WoLA 93.7 (0.3) 93.3 (0.3) 92.5 (0.3) 81.3 (0.8) 93.8 (0.3) 93.3 (0.3) 92.1 (0.6) 76.8 (5.4) 93.6 (0.6) 92.9 (0.7) 88.1 (2.8) 47.3 (17.0)
WoLA+NNM 93.7 (0.3) 93.4 (0.3) 93.0 (0.3) 83.8 (0.7) 93.9 (0.4) 93.5 (0.3) 92.8 (0.4) 73.5 (13.0) 93.7 (0.6) 93.3 (0.7) 89.8 (2.8) 48.7 (15.8)
Accuracy Gain Over Top Non-WoLA Methods
WoLA 0.0 0.0 0.0 +6.4 0.0 +0.4 +2.4 +17.3 -0.4 +1.5 +21.4 +16.5
WoLA+NNM 0.0 +0.2 +0.6 +8.8 +0.1 +0.6 +3.1 +19.9 -0.2 +1.8 +22.5 +26.6
WoLA 0.0 0.0 -0.2 +5.9 -0.1 +0.3 +2.0 +13.6 -0.4 +1.4 +20.5 +9.3
WoLA+NNM 0.0 +0.1 +0.3 +8.3 -0.1 +0.5 +2.8 +10.3 -0.2 +1.8 +22.3 +10.7
(b) Fashion MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 77.1 (0.7) 74.0 (1.2) 64.7 (3.1) 45.3 (6.3) 76.6 (1.9) 71.6 (2.9) 59.2 (5.5) 32.3 (7.7) 72.6 (2.6) 60.3 (4.9) 41.2 (7.4) 16.7 (6.9)
BKT 76.8 (0.7) 74.5 (1.1) 64.7 (3.1) 45.3 (6.3) 77.2 (1.7) 73.6 (2.5) 59.2 (5.5) 32.3 (7.7) 75.7 (2.8) 67.4 (4.7) 41.2 (7.4) 16.7 (6.9)
FoundFL 71.2 (3.1) 62.2 (5.8) 52.3 (3.9) 44.7 (7.7) 63.6 (6.9) 54.9 (7.0) 47.7 (8.9) 41.6 (9.5) 47.8 (10.6) 41.7 (9.6) 38.1 (9.1) 31.2 (9.3)
NNM 78.4 (0.6) 76.5 (0.9) 72.3 (1.8) 54.8 (5.9) 79.4 (1.3) 76.7 (1.7) 69.9 (3.3) 48.8 (7.4) 79.4 (1.9) 74.0 (3.5) 54.2 (9.2) 30.5 (6.7)
WoLA 78.6 (0.6) 77.6 (0.6) 75.8 (0.8) 63.9 (1.4) 79.5 (1.3) 78.1 (1.6) 76.0 (1.7) 62.3 (4.0) 78.0 (1.8) 76.3 (2.3) 70.8 (4.8) 42.3 (7.3)
WoLA+NNM 78.6 (0.6) 78.0 (0.6) 77.1 (0.7) 68.5 (1.0) 79.6 (1.3) 78.6 (1.5) 77.5 (1.4) 67.5 (4.0) 78.5 (1.8) 77.4 (2.3) 73.9 (5.0) 48.8 (8.6)
WoLA 78.5 (0.5) 77.3 (0.5) 74.5 (0.9) 59.1 (2.3) 79.3 (1.5) 76.9 (1.6) 72.8 (2.0) 55.0 (6.0) 77.7 (1.7) 73.7 (2.4) 64.9 (4.3) 36.4 (11.2)
WoLA+NNM 78.5 (0.6) 77.6 (0.4) 75.9 (0.8) 64.2 (2.1) 79.6 (1.5) 77.3 (1.6) 74.9 (1.5) 55.7 (9.3) 78.3 (1.6) 75.0 (2.3) 69.3 (4.8) 38.9 (8.3)
Accuracy Gain Over Top Non-WoLA Methods
WoLA +0.2 +1.1 +3.6 +9.1 +0.1 +1.4 +6.1 +13.4 -1.4 +2.2 +16.6 +11.1
WoLA+NNM +0.2 +1.5 +4.8 +13.7 +0.2 +1.9 +7.6 +18.7 -0.9 +3.4 +19.8 +17.6
WoLA +0.1 +0.7 +2.2 +4.3 -0.1 +0.2 +2.9 +6.2 -1.7 -0.4 +10.7 +5.3
WoLA+NNM +0.1 +1.1 +3.6 +9.4 +0.2 +0.6 +5.0 +6.9 -1.1 +0.9 +15.1 +7.7
(c) Purchase100
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 70.4 (0.7) 65.3 (1.1) 50.5 (1.9) 24.8 (2.8) 67.9 (1.0) 58.0 (2.0) 37.1 (3.1) 8.0 (2.1) 62.2 (1.8) 44.4 (3.3) 15.9 (2.3) 2.6 (1.1)
BKT 70.1 (0.9) 64.6 (1.2) 50.5 (1.9) 24.8 (2.8) 68.2 (1.1) 59.3 (1.9) 37.1 (3.1) 8.0 (2.1) 64.5 (1.9) 49.3 (3.5) 15.9 (2.3) 2.6 (1.1)
FoundFL 52.2 (3.6) 36.5 (3.1) 28.4 (2.6) 22.4 (2.4) 38.8 (5.9) 28.9 (4.1) 20.7 (4.5) 13.3 (4.9) 24.6 (8.2) 15.6 (6.6) 10.3 (4.3) 6.4 (2.8)
NNM 71.6 (0.7) 67.9 (0.9) 61.3 (1.8) 38.1 (2.3) 70.6 (0.9) 63.9 (1.6) 52.0 (2.8) 28.3 (4.0) 69.6 (1.2) 58.8 (3.2) 28.9 (5.6) 15.1 (3.2)
WoLA 70.6 (0.6) 69.2 (0.6) 66.7 (0.7) 52.7 (1.0) 69.2 (0.8) 67.0 (0.8) 61.3 (1.2) 43.1 (1.7) 63.2 (1.1) 57.1 (2.1) 43.6 (2.6) 21.8 (2.5)
WoLA+NNM 70.9 (0.6) 69.8 (0.6) 68.3 (0.7) 57.6 (0.9) 69.3 (0.8) 67.8 (0.7) 65.3 (1.1) 53.5 (1.7) 64.0 (1.1) 60.6 (1.7) 52.6 (2.5) 33.0 (3.7)
WoLA 70.5 (0.6) 68.8 (0.5) 65.2 (0.6) 48.2 (1.7) 69.0 (0.8) 66.3 (0.9) 59.3 (1.6) 33.5 (6.1) 62.9 (1.0) 55.9 (1.9) 40.9 (3.4) 11.2 (5.5)
WoLA+NNM 70.7 (0.6) 69.3 (0.5) 67.0 (0.6) 53.3 (2.6) 69.2 (0.7) 67.2 (0.8) 63.1 (1.5) 39.9 (9.9) 63.7 (1.0) 59.3 (1.7) 45.7 (5.2) 19.5 (7.6)
Accuracy Gain Over Top Non-WoLA Methods
WoLA -1.0 +1.3 +5.4 +14.6 -1.4 +3.1 +9.3 +14.8 -6.4 -1.7 +14.6 +6.8
WoLA+NNM -0.8 +1.8 +7.0 +19.5 -1.3 +3.9 +13.3 +25.2 -5.6 +1.8 +23.7 +17.9
WoLA -1.2 +0.8 +4.0 +10.2 -1.6 +2.4 +7.2 +5.2 -6.7 -2.9 +11.9 -3.9
WoLA+NNM -0.9 +1.4 +5.7 +15.3 -1.5 +3.2 +11.1 +11.6 -5.9 +0.5 +16.7 +4.5
(d) CIFAR10
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 69.4 (1.1) 60.4 (2.1) 41.4 (4.2) 21.5 (4.2) 63.9 (1.7) 48.1 (3.3) 32.2 (3.6) 16.1 (3.5) 51.9 (3.5) 36.7 (4.5) 24.5 (4.8) 12.0 (3.5)
BKT 67.9 (1.0) 61.6 (2.0) 41.4 (4.2) 21.5 (4.2) 62.3 (1.8) 51.4 (3.3) 32.2 (3.6) 16.1 (3.5) 52.8 (2.9) 37.5 (3.7) 24.5 (4.8) 12.0 (3.5)
FoundFL 45.1 (12.2) 35.1 (9.9) 32.6 (9.6) 22.5 (8.6) 27.2 (6.8) 20.5 (6.1) 25.2 (5.8) 19.3 (5.7) 20.0 (4.4) 18.2 (4.6) 18.3 (3.9) 15.2 (4.6)
NNM 71.7 (0.8) 66.9 (1.5) 56.7 (3.9) 38.8 (7.0) 68.7 (1.2) 59.1 (3.4) 40.6 (5.2) 29.2 (6.5) 62.7 (2.4) 48.7 (3.7) 32.5 (5.4) 22.5 (3.9)
WoLA 71.6 (0.6) 67.2 (0.8) 58.8 (1.2) 40.0 (2.2) 68.7 (0.7) 63.2 (1.1) 51.9 (2.9) 32.3 (3.9) 62.5 (1.8) 54.0 (3.1) 38.9 (4.7) 20.3 (5.0)
WoLA+NNM 72.7 (0.5) 70.0 (0.8) 64.2 (1.1) 51.6 (2.1) 70.3 (0.6) 66.9 (0.9) 59.1 (2.7) 45.0 (3.8) 65.4 (1.5) 60.1 (3.1) 49.0 (3.9) 34.2 (7.2)
WoLA 71.5 (0.6) 66.4 (0.9) 56.3 (1.6) 34.4 (3.3) 68.4 (0.7) 62.1 (1.5) 46.9 (3.9) 25.6 (6.4) 62.2 (1.9) 52.4 (3.7) 33.8 (5.5) 16.3 (4.8)
WoLA+NNM 72.7 (0.5) 69.4 (0.8) 62.5 (1.2) 45.4 (3.2) 69.9 (0.7) 66.2 (1.1) 55.4 (2.8) 37.5 (8.5) 65.0 (1.6) 58.2 (3.2) 42.7 (4.9) 21.3 (7.5)
Accuracy Gain Over Top Non-WoLA Methods
WoLA -0.1 +0.3 +2.2 +1.2 -0.1 +4.1 +11.3 +3.1 -0.3 +5.4 +6.4 -2.1
WoLA+NNM +1.0 +3.2 +7.5 +12.9 +1.5 +7.9 +18.5 +15.8 +2.6 +11.4 +16.4 +11.8
WoLA -0.2 -0.5 -0.4 -4.4 -0.3 +3.0 +6.3 -3.6 -0.5 +3.7 +1.3 -6.2
WoLA+NNM +0.9 +2.5 +5.9 +6.7 +1.2 +7.1 +14.7 +8.3 +2.2 +9.5 +10.2 -1.2
Table 3: Worst-case Performance: Test Accuracy (s.d.) % under the worst attack, averaged over training. Results are averaged across aggregators (CWMed, CwTM, GM, MKrum). Colors rank values within columns: yellow (1st), blue (2nd), red (3rd). See full table in Appendix.
(a) MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 93.4 (0.3) 91.9 (0.5) 73.6 (5.1) 27.5 (5.8) 93.3 (0.4) 90.1 (1.3) 62.8 (8.1) 12.6 (5.7) 89.3 (3.2) 70.7 (8.6) 29.5 (11.5) 7.1 (3.4)
BKT 93.2 (0.5) 91.5 (0.9) 73.6 (5.1) 27.5 (5.8) 93.2 (0.5) 90.8 (1.5) 62.8 (8.1) 12.6 (5.7) 92.6 (1.1) 84.6 (4.0) 29.5 (11.5) 7.1 (3.4)
FoundFL 80.4 (4.4) 36.5 (8.2) 10.5 (1.2) 10.9 (1.7) 76.4 (15.1) 27.1 (8.7) 10.9 (1.2) 13.5 (3.3) 48.5 (22.4) 20.5 (6.3) 14.7 (4.0) 16.9 (7.6)
NNM 93.6 (0.2) 92.8 (0.3) 91.3 (0.6) 47.0 (27.5) 93.8 (0.3) 91.6 (1.7) 83.9 (8.9) 22.5 (11.7) 93.7 (0.6) 85.6 (3.7) 49.6 (24.0) 18.1 (5.9)
WoLA 93.5 (0.3) 92.9 (0.3) 90.6 (0.4) 66.7 (0.7) 93.8 (0.3) 92.9 (0.3) 89.9 (0.7) 64.3 (3.0) 93.3 (0.5) 92.1 (0.8) 77.6 (8.1) 28.8 (6.2)
WoLA+NNM 93.6 (0.3) 93.2 (0.3) 92.6 (0.2) 70.5 (1.6) 93.9 (0.3) 93.2 (0.3) 92.3 (0.4) 66.7 (4.7) 93.6 (0.5) 92.6 (0.7) 78.9 (11.2) 24.1 (11.4)
WoLA 93.5 (0.3) 92.8 (0.3) 90.6 (0.4) 67.3 (1.0) 93.7 (0.3) 92.8 (0.3) 89.6 (0.8) 63.7 (8.9) 93.3 (0.5) 92.0 (0.8) 75.9 (8.1) 24.5 (7.3)
WoLA+NNM 93.6 (0.3) 93.1 (0.3) 92.3 (0.4) 70.3 (0.7) 93.7 (0.3) 93.1 (0.3) 91.9 (0.5) 51.4 (27.9) 93.6 (0.6) 92.6 (0.8) 80.2 (9.5) 14.9 (6.6)
Accuracy Gain Over Top Non-WoLA Methods
WoLA -0.1 +0.1 -0.8 +19.7 -0.1 +1.3 +6.0 +41.9 -0.4 +6.5 +28.1 +10.7
WoLA+NNM 0.0 +0.3 +1.3 +23.5 0.0 +1.6 +8.3 +44.3 -0.1 +7.0 +29.3 +6.0
WoLA -0.1 0.0 -0.8 +20.3 -0.1 +1.3 +5.7 +41.3 -0.4 +6.4 +26.3 +6.5
WoLA+NNM +0.1 +0.2 +0.9 +23.2 -0.1 +1.5 +7.9 +28.9 -0.1 +7.0 +30.6 -3.2
(b) Fashion MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 75.6 (0.8) 69.6 (1.3) 50.6 (7.7) 18.3 (8.8) 73.7 (2.6) 64.6 (4.0) 40.8 (8.5) 10.1 (4.7) 66.7 (2.6) 45.2 (5.1) 22.5 (6.1) 4.6 (3.3)
BKT 75.9 (0.8) 71.4 (1.5) 50.6 (7.7) 18.3 (8.8) 76.2 (2.0) 70.2 (2.8) 40.8 (8.5) 10.1 (4.7) 74.5 (3.3) 59.2 (6.6) 22.5 (6.1) 4.6 (3.3)
FoundFL 65.5 (4.5) 35.9 (8.8) 11.7 (1.6) 12.2 (3.9) 53.6 (15.5) 23.6 (10.3) 20.6 (11.9) 15.0 (5.2) 31.4 (12.0) 25.7 (10.1) 15.8 (5.9) 14.8 (7.2)
NNM 77.5 (0.5) 73.3 (1.2) 62.9 (4.1) 25.3 (13.4) 77.9 (1.6) 73.4 (2.3) 54.5 (7.2) 18.8 (8.0) 77.2 (2.2) 69.2 (4.7) 36.6 (12.8) 14.1 (2.7)
WoLA 77.9 (0.6) 75.8 (0.5) 71.8 (1.0) 51.5 (1.5) 78.9 (1.3) 76.2 (1.6) 71.6 (2.2) 46.5 (7.0) 77.0 (1.9) 73.4 (2.5) 58.9 (6.2) 24.7 (7.5)
WoLA+NNM 78.0 (0.7) 76.6 (0.6) 74.8 (0.6) 54.8 (1.7) 79.1 (1.4) 76.9 (1.5) 75.0 (1.6) 49.5 (8.3) 77.9 (1.8) 74.9 (2.6) 61.6 (11.0) 18.9 (7.9)
WoLA 77.9 (0.5) 75.4 (0.4) 70.1 (1.0) 48.1 (2.5) 78.7 (1.3) 75.1 (1.6) 67.2 (3.4) 44.5 (5.1) 76.7 (1.5) 70.1 (2.8) 47.7 (5.9) 18.0 (7.4)
WoLA+NNM 77.8 (0.7) 75.9 (0.4) 73.2 (1.1) 50.8 (2.4) 79.1 (1.6) 75.7 (1.8) 71.1 (2.0) 35.4 (18.6) 77.8 (1.6) 71.8 (2.2) 57.9 (8.2) 11.3 (5.1)
Accuracy Gain Over Top Non-WoLA Methods
WoLA +0.5 +2.5 +8.9 +26.1 +1.0 +2.8 +17.1 +27.7 -0.3 +4.2 +22.3 +9.9
WoLA+NNM +0.5 +3.3 +11.9 +29.4 +1.2 +3.5 +20.5 +30.7 +0.7 +5.6 +25.0 +4.1
WoLA +0.4 +2.2 +7.2 +22.8 +0.9 +1.7 +12.7 +25.7 -0.5 +0.9 +11.1 +3.2
WoLA+NNM +0.3 +2.6 +10.3 +25.4 +1.2 +2.3 +16.5 +16.7 +0.6 +2.5 +21.4 -3.5
(c) Purchase100
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 68.2 (0.6) 59.2 (0.8) 29.0 (1.2) 8.7 (2.3) 64.0 (1.5) 48.5 (2.9) 18.5 (3.1) 0.9 (0.3) 54.9 (2.5) 28.5 (3.8) 2.1 (0.8) 0.9 (0.3)
BKT 68.6 (0.8) 58.1 (0.9) 29.0 (1.2) 8.7 (2.3) 65.8 (1.6) 50.3 (3.1) 18.5 (3.1) 0.9 (0.3) 58.2 (3.4) 30.0 (7.1) 2.1 (0.8) 0.9 (0.3)
FoundFL 33.6 (1.9) 4.2 (1.7) 0.6 (0.3) 0.4 (0.1) 21.5 (6.8) 2.6 (1.3) 1.6 (0.9) 1.6 (0.9) 13.2 (4.7) 3.2 (1.1) 2.9 (0.3) 2.5 (1.3)
NNM 70.4 (0.5) 63.3 (1.5) 52.2 (1.4) 6.4 (1.0) 67.1 (1.7) 53.8 (4.3) 33.6 (4.4) 3.2 (1.8) 61.8 (3.1) 36.9 (8.1) 3.2 (2.3) 1.0 (0.4)
WoLA 69.7 (0.6) 66.7 (0.6) 61.4 (0.7) 37.7 (1.0) 68.0 (0.8) 63.8 (0.8) 49.0 (1.4) 19.3 (1.2) 60.3 (1.0) 47.3 (2.3) 18.2 (2.6) 6.1 (2.8)
WoLA+NNM 70.3 (0.6) 68.0 (0.6) 66.3 (0.6) 44.2 (0.8) 68.7 (0.8) 65.8 (0.6) 62.0 (1.5) 37.6 (1.9) 61.7 (1.0) 53.5 (1.5) 39.7 (3.3) 3.0 (1.8)
WoLA 69.5 (0.6) 66.2 (0.5) 59.8 (0.6) 31.9 (1.9) 67.7 (0.8) 62.9 (0.8) 45.7 (2.5) 21.6 (4.6) 59.9 (0.9) 45.8 (2.0) 23.4 (4.5) 4.2 (1.0)
WoLA+NNM 70.2 (0.6) 67.5 (0.5) 64.8 (0.4) 36.3 (6.4) 68.5 (0.7) 65.0 (0.7) 58.4 (1.6) 20.0 (13.5) 61.4 (0.9) 51.6 (1.6) 17.4 (15.2) 1.9 (0.7)
Accuracy Gain Over Top Non-WoLA Methods
WoLA -0.7 +3.4 +9.2 +29.0 +0.8 +10.0 +15.4 +16.1 -1.5 +10.4 +15.1 +3.6
WoLA+NNM -0.1 +4.7 +14.1 +35.5 +1.5 +12.0 +28.4 +34.4 -0.1 +16.6 +36.5 +0.5
WoLA -0.9 +2.8 +7.6 +23.2 +0.5 +9.1 +12.1 +18.4 -1.8 +8.9 +20.3 +1.7
WoLA+NNM -0.2 +4.2 +12.6 +27.7 +1.3 +11.2 +24.8 +16.8 -0.4 +14.8 +14.2 -0.6
(d) CIFAR10
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 63.9 (2.9) 50.1 (4.7) 25.4 (8.0) 14.4 (3.5) 55.3 (5.3) 22.2 (5.0) 16.3 (4.1) 10.9 (3.1) 36.3 (5.4) 17.7 (5.5) 15.4 (5.2) 8.2 (2.7)
BKT 65.2 (1.6) 55.6 (3.7) 25.4 (8.0) 14.4 (3.5) 58.3 (3.8) 34.0 (7.3) 16.3 (4.1) 10.9 (3.1) 42.2 (6.3) 18.0 (5.0) 15.4 (5.2) 8.2 (2.7)
FoundFL 36.7 (13.2) 14.0 (4.2) 10.6 (1.8) 11.0 (2.8) 21.6 (6.2) 12.0 (2.5) 11.0 (1.9) 11.3 (3.3) 16.9 (4.0) 12.6 (3.1) 12.7 (2.8) 9.6 (2.1)
NNM 65.0 (2.3) 53.3 (4.6) 42.0 (8.6) 19.2 (8.3) 58.8 (4.1) 31.3 (8.9) 15.2 (4.2) 14.0 (3.0) 45.7 (5.8) 17.9 (6.4) 15.6 (3.7) 11.8 (3.1)
WoLA 69.7 (0.9) 64.0 (1.1) 49.0 (1.5) 23.5 (3.0) 66.0 (1.1) 59.2 (2.1) 39.7 (5.1) 21.7 (2.1) 57.0 (2.2) 46.9 (4.3) 28.4 (3.8) 11.9 (2.5)
WoLA+NNM 70.5 (1.0) 65.2 (1.6) 57.7 (1.4) 34.6 (3.5) 67.1 (1.6) 60.3 (2.3) 49.2 (3.5) 27.3 (2.9) 59.2 (1.6) 49.7 (6.2) 37.1 (7.3) 22.0 (3.2)
WoLA 69.5 (0.9) 62.9 (1.5) 44.1 (2.5) 18.9 (2.5) 65.6 (1.3) 57.4 (2.1) 33.5 (5.4) 13.3 (2.6) 56.8 (2.2) 45.1 (6.2) 21.5 (5.5) 8.4 (4.7)
WoLA+NNM 70.6 (0.9) 64.3 (1.6) 55.0 (1.2) 34.0 (2.7) 66.3 (2.0) 59.5 (2.4) 44.2 (4.1) 26.5 (10.4) 58.7 (1.8) 45.6 (6.0) 28.8 (7.2) 13.8 (5.7)
Accuracy Gain Over Top Non-WoLA Methods
WoLA +4.5 +8.4 +6.9 +4.3 +7.2 +25.2 +23.4 +7.7 +11.3 +28.8 +12.8 +0.1
WoLA+NNM +5.4 +9.5 +15.7 +15.5 +8.3 +26.3 +32.9 +13.4 +13.5 +31.7 +21.5 +10.2
WoLA +4.4 +7.3 +2.1 -0.3 +6.8 +23.4 +17.2 -0.7 +11.1 +27.0 +5.8 -3.4
WoLA+NNM +5.4 +8.6 +13.0 +14.8 +7.5 +25.4 +28.0 +12.5 +13.0 +27.5 +13.1 +2.0
Table 4: Gradient Dissimilarity (s.d.), averaged over training. Results are averaged across aggregators (CWMed, CwTM, GM, MKrum) and attacks (ALIE, FOE, LF, Mimic, SF). Colors rank values within columns: yellow (1st), blue (2nd), red (3rd).
(a) MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 0.1 (0.03) 0.33 (0.09) 1.09 (0.3) 2.07 (0.35) 0.69 (0.15) 1.8 (0.36) 3.27 (0.85) 4.28 (1.18) 4.46 (0.94) 7.04 (1.35) 8.91 (1.55) 4.56 (0.98)
BKT 0.06 (0.01) 0.12 (0.03) 1.09 (0.3) 2.07 (0.35) 0.13 (0.02) 0.36 (0.14) 3.27 (0.85) 4.28 (1.18) 0.55 (0.18) 2.79 (0.67) 8.91 (1.55) 4.56 (0.98)
FoundFL 0.37 (0.13) 0.84 (0.32) 0.59 (0.19) 0.52 (0.27) 0.84 (0.45) 1.28 (0.58) 0.88 (0.46) 0.89 (0.47) 4.54 (1.83) 3.86 (1.83) 3.61 (1.78) 3.45 (1.86)
NNM 0.06 (0.01) 0.08 (0.02) 0.13 (0.03) 0.83 (0.19) 0.12 (0.02) 0.36 (0.24) 0.71 (0.31) 1.92 (0.81) 0.38 (0.08) 1.22 (0.35) 4.88 (1.38) 4.74 (1.02)
WoLA 0.03 (0.0) 0.03 (0.0) 0.06 (0.02) 0.11 (0.02) 0.04 (0.0) 0.06 (0.01) 0.11 (0.03) 0.33 (0.18) 0.07 (0.01) 0.16 (0.05) 0.69 (0.35) 1.68 (0.42)
WoLA+NNM 0.03 (0.0) 0.03 (0.0) 0.04 (0.01) 0.05 (0.01) 0.04 (0.0) 0.05 (0.01) 0.06 (0.01) 0.19 (0.12) 0.05 (0.01) 0.07 (0.01) 0.46 (0.33) 0.81 (0.18)
WoLA 0.03 (0.0) 0.04 (0.0) 0.06 (0.01) 0.1 (0.03) 0.04 (0.01) 0.05 (0.01) 0.13 (0.06) 0.68 (0.68) 0.07 (0.01) 0.15 (0.03) 0.55 (0.26) 2.24 (0.8)
WoLA+NNM 0.03 (0.0) 0.03 (0.0) 0.03 (0.01) 0.06 (0.01) 0.04 (0.0) 0.05 (0.01) 0.06 (0.02) 0.36 (0.29) 0.05 (0.01) 0.07 (0.01) 0.37 (0.25) 1.5 (0.87)
(b) Fashion MNIST
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 3.84 (0.56) 4.2 (0.83) 4.22 (0.76) 2.84 (0.62) 8.15 (1.76) 7.67 (1.56) 7.32 (1.5) 4.15 (1.39) 14.01 (1.24) 13.47 (1.37) 11.15 (1.69) 4.55 (1.5)
BKT 1.26 (0.11) 1.82 (0.23) 4.22 (0.76) 2.84 (0.62) 2.45 (0.56) 3.8 (1.17) 7.32 (1.5) 4.15 (1.39) 5.5 (0.84) 8.82 (1.23) 11.15 (1.69) 4.55 (1.5)
FoundFL 1.28 (0.69) 1.31 (0.68) 1.08 (0.45) 0.76 (0.32) 3.01 (1.32) 2.18 (1.18) 2.24 (1.28) 1.92 (1.01) 6.18 (2.86) 5.93 (2.26) 5.7 (2.17) 4.02 (1.96)
NNM 2.54 (0.26) 2.54 (0.34) 2.92 (0.48) 1.92 (0.67) 5.11 (1.32) 5.0 (1.02) 5.26 (1.52) 3.23 (1.12) 9.94 (1.62) 10.1 (1.87) 9.48 (2.09) 6.0 (1.29)
WoLA 0.18 (0.02) 0.17 (0.02) 0.17 (0.03) 0.14 (0.02) 0.41 (0.17) 0.38 (0.15) 0.43 (0.17) 0.29 (0.12) 0.89 (0.15) 1.16 (0.24) 1.57 (0.48) 1.51 (0.7)
WoLA+NNM 0.17 (0.02) 0.15 (0.02) 0.13 (0.02) 0.11 (0.02) 0.38 (0.16) 0.29 (0.11) 0.28 (0.11) 0.28 (0.14) 0.78 (0.12) 0.72 (0.12) 0.97 (0.38) 0.99 (0.33)
WoLA 0.19 (0.02) 0.19 (0.03) 0.18 (0.05) 0.15 (0.04) 0.51 (0.2) 0.51 (0.21) 0.51 (0.25) 0.73 (0.7) 0.95 (0.17) 1.35 (0.31) 1.53 (0.27) 2.17 (0.88)
WoLA+NNM 0.18 (0.02) 0.15 (0.02) 0.14 (0.04) 0.12 (0.03) 0.45 (0.18) 0.4 (0.16) 0.35 (0.16) 0.4 (0.29) 0.82 (0.1) 0.79 (0.13) 0.91 (0.23) 1.46 (0.75)
(c) Purchase100
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 0.14 (0.01) 0.2 (0.02) 0.27 (0.02) 0.42 (0.05) 0.55 (0.08) 0.67 (0.1) 0.99 (0.11) 1.55 (0.23) 2.14 (0.29) 2.84 (0.33) 3.59 (0.43) 3.74 (0.42)
BKT 0.13 (0.01) 0.17 (0.01) 0.27 (0.02) 0.42 (0.05) 0.39 (0.04) 0.53 (0.05) 0.99 (0.11) 1.55 (0.23) 1.29 (0.16) 2.12 (0.28) 3.59 (0.43) 3.74 (0.42)
FoundFL 0.24 (0.02) 0.32 (0.03) 0.28 (0.02) 0.32 (0.03) 0.98 (0.17) 1.01 (0.12) 0.98 (0.15) 1.04 (0.17) 4.29 (0.73) 3.79 (0.68) 3.4 (0.46) 2.94 (0.54)
NNM 0.12 (0.01) 0.15 (0.01) 0.16 (0.02) 0.24 (0.02) 0.34 (0.04) 0.45 (0.05) 0.6 (0.1) 0.99 (0.19) 1.03 (0.14) 1.76 (0.36) 3.03 (0.52) 3.45 (0.49)
WoLA 0.02 (0.0) 0.02 (0.0) 0.02 (0.0) 0.02 (0.0) 0.03 (0.0) 0.04 (0.0) 0.04 (0.0) 0.05 (0.01) 0.09 (0.01) 0.12 (0.01) 0.15 (0.02) 0.24 (0.03)
WoLA+NNM 0.02 (0.0) 0.01 (0.0) 0.02 (0.0) 0.01 (0.0) 0.03 (0.0) 0.03 (0.0) 0.04 (0.0) 0.03 (0.0) 0.08 (0.0) 0.09 (0.01) 0.11 (0.01) 0.16 (0.02)
WoLA 0.02 (0.0) 0.02 (0.0) 0.02 (0.0) 0.05 (0.01) 0.03 (0.0) 0.04 (0.0) 0.06 (0.01) 0.13 (0.05) 0.09 (0.0) 0.14 (0.03) 0.26 (0.09) 0.44 (0.19)
WoLA+NNM 0.02 (0.0) 0.02 (0.0) 0.02 (0.0) 0.03 (0.01) 0.03 (0.0) 0.03 (0.0) 0.04 (0.01) 0.11 (0.06) 0.08 (0.0) 0.11 (0.02) 0.23 (0.09) 0.41 (0.19)
(d) CIFAR10
Strategy $\alpha=3\quad f=2,4,6,8$ $\alpha=1\quad f=2,4,6,8$ $\alpha=0.3\quad f=2,4,6,8$
Accuracy
- 1.81 (0.27) 2.49 (0.26) 2.45 (0.33) 1.67 (0.48) 5.06 (0.8) 5.73 (0.88) 4.68 (0.83) 2.12 (0.96) 8.56 (0.7) 9.14 (0.86) 7.11 (1.04) 2.93 (1.25)
BKT 0.8 (0.06) 1.09 (0.08) 2.45 (0.33) 1.67 (0.48) 1.57 (0.2) 2.26 (0.49) 4.68 (0.83) 2.12 (0.96) 2.84 (0.26) 4.12 (0.64) 7.11 (1.04) 2.93 (1.25)
FoundFL 1.36 (0.56) 1.24 (0.48) 1.46 (0.56) 1.08 (0.63) 2.45 (1.0) 1.91 (1.22) 3.39 (1.24) 3.08 (1.34) 3.84 (1.68) 3.6 (1.37) 4.11 (2.21) 2.93 (1.91)
NNM 1.13 (0.11) 1.49 (0.23) 1.36 (0.34) 1.28 (0.4) 2.24 (0.29) 3.11 (0.74) 2.29 (0.76) 1.96 (0.78) 3.39 (0.24) 5.01 (0.92) 3.43 (0.73) 3.07 (0.94)
WoLA 0.47 (0.03) 0.44 (0.03) 0.38 (0.04) 0.36 (0.11) 0.72 (0.05) 0.78 (0.1) 0.68 (0.11) 0.54 (0.14) 1.01 (0.08) 1.23 (0.24) 1.16 (0.38) 0.87 (0.35)
WoLA+NNM 0.44 (0.02) 0.39 (0.02) 0.37 (0.03) 0.43 (0.11) 0.63 (0.04) 0.63 (0.06) 0.65 (0.1) 0.61 (0.13) 0.84 (0.06) 0.99 (0.14) 1.01 (0.23) 0.94 (0.28)
WoLA 0.47 (0.03) 0.44 (0.03) 0.38 (0.04) 0.32 (0.05) 0.72 (0.05) 0.76 (0.1) 0.69 (0.15) 0.54 (0.25) 0.99 (0.09) 1.18 (0.19) 1.18 (0.25) 1.09 (0.47)
WoLA+NNM 0.44 (0.02) 0.39 (0.02) 0.37 (0.04) 0.35 (0.05) 0.63 (0.04) 0.62 (0.05) 0.6 (0.08) 0.53 (0.2) 0.82 (0.07) 0.96 (0.15) 0.92 (0.22) 1.07 (0.47)
7 Experimental Setup
This section describes our experimental setup, including the baseline methods, the WoLA approach under training objective attack, and the datasets used. We detail the configuration of the heterogeneity regimes, workers, and the types of Byzantine attacks we consider. We also present the robust aggregation methods, evaluation metrics, optimization algorithms, models, and hyperparameters we used. Finally, we explain how we ensure the reproducibility of our experiments.
7.1 Baseline Approaches
We compare WoLA against three baselines: Bucketing (BKT) [24], NNM [3], and FoundationFL (FoundFL) [15]. BKT requires setting a bucket size $s$. Following the authors’ recommendation, we use $s=\left\lfloor{n}/({2f})\right\rfloor$. When $f>n/4$, the bucket size drops to $1$, making BKT ineffective. For instance, with $17$ workers, BKT has no effect when $f=6$ or $8$. FoundFL requires a hyperparameter $m$. As suggested in the original paper, we set $m=n/2$, i.e., $m=8$ in all our experiments. We also include a baseline using only a robust aggregator without any additional defense, denoted as -.
We evaluate different variants of WoLA. In particular, to assess robustness against training objective attacks, we implement the worst-case scenario described in Subsection 7.2, denoting the resulting loss as WoLA. Additionally, we examine the combination of WoLA and WoLA with NNM.
We do not compare WoLA with methods that (i) require a server-side dataset [4, 28, 11, 10, 30], (ii) involve high computational costs [4, 12, 32, 47, 39], or (iii) rely on hyperparameter tuning [46, 35].
7.2 WoLA under Training Objective Attack
To define the training objective, we consider a scenario where workers share their local label distributions with the server. In this case, Byzantine machines may try to manipulate the objective by sending malicious distributions. We analyze the impact of such attacks.
Just as Byzantine workers can send arbitrary updates with full knowledge of honest gradients, we assume they can also send arbitrary distributions with full knowledge of local label distributions. Rather than listing all possible attacks, we focus on one that causes the maximum distortion: Byzantine machines send a zero vector except for a single entry set to $1$, corresponding to the minority class in the global dataset (proof are provided in Appendix A.7).
To defend against attacks, we propose using the geometric median as the aggregator. This approach offers two key advantages: (i) Unlike methods like Trimmed Mean or Krum, which exclude outliers, the geometric median treats all updates equally. This is important because outlier distributions are not necessarily malicious. Excluding them may discard honest distribution and degrade performance. (ii) When aggregating probability distributions, the geometric median produces a probability distribution—unlike the coordinate-wise median, for instance—ensuring the training objective retains its expected form.
Our goal is to show that a strong attack, even when countered by a simple defense, has limited impact on model performance. As the chosen attack represents a worst-case scenario, it provides a meaningful benchmark for evaluating training objective attacks.
We refer to our approach in this scenario as WoLA.
7.3 Datasets, Heterogeneity & Workers
We conduct the experiments on four datasets: MNIST [27], Fashion-MNIST [42], CIFAR10 [26], and Purchase100 [37]. Purchase100 contains shopping records for thousands of customers. We use a simplified version [37], where each customer is represented by $600$ binary features indicating item purchases. The data is clustered into $100$ classes, and the task is to predict each customer’s class.
To simulate data heterogeneity across workers, we sample from the global dataset using a Dirichlet distribution with concentration parameter $\alpha$, following the method in [3, 23]. We consider three heterogeneity levels—low, medium, and high—corresponding to $\alpha\in\{3,1,0.3\}$.
We use $n=17$ workers in all experiments. The number of Byzantine workers varies as $f\in\{2,4,6,8\}$. Each worker holds a local dataset of fixed size, equal to the total dataset size divided by the number of honest workers.
7.4 Byzantine Attacks & Robust Aggregators
Byzantine workers carry out one of five state-of-the-art gradient attacks: Fall of Empires (FOE) [43], A Little is Enough (ALIE) [5], Sign Flipping (SF) [2], Label Flipping (LF) [2], and Mimic [24].
In the Mimic attack, all Byzantine workers replicate a selected honest update and send it to the server. Inspired by the ALIE attack—which places Byzantine values between the honest mean and outliers to overcome robustness defenses—we instead select the honest outlier that is most surrounded by other honest updates (see Appendix for details). With this strategy, the “extreme” honest outliers hide Byzantine updates, which makes the attack harder to detect while remaining effective in heterogeneous settings. For instance, based on the extensive results in Appendix B.4, Tables 14, 15, and 16, Mimic often appears as the worst attack for the strategies –, BKT, and NNM.
All state-of-the-art strategies are evaluated using four robust aggregators: CWMed, CwTM, GM, and MKrum.
7.5 Metrics, Tables & Boxplots
We use two metrics: (i) Test Accuracy, averaged over training with a test set that matches the global training label distribution, and (ii) Gradient Dissimilarity, also averaged over training.
We ran experiments on four datasets, three levels of heterogeneity, four aggregation methods, and five attacks, averaging results over three random seeds for each scenario. Full results are in Appendix B.4. For clarity, we summarize key findings in three tables. Table 2 shows Test Accuracy, averaged over all aggregation methods (CWMed, CwTM, GM, MKrum) and all attacks (ALIE, FOE, LF, Mimic, SF), to reflect overall performance. Table 4 presents the same summary for Gradient Dissimilarity. Table 3 shows Test Accuracy under the strongest attack (i.e., the one causing the lowest performance), averaged over all aggregators, to highlight worst-case performance.
In Tables 2 and 3, we report the Accuracy Gain Over Top Non-WoLA Methods—the difference between each WoLA variant and the best-performing non-WoLA method (among –, BKT, FoundFL, and NNM). These values highlight performance gains over state-of-the-art strategies.
We also include boxplots of the gains from Tables 2 and 3 in Figures 4(a) and 4(b), to provide an overall view. In the following, when describing a given situation, we refer to the typical performance as the interquartile range of the corresponding boxplot. The average gain is also taken from these boxplots.
7.6 Optimization Algorithm
WoLA relies on approximating the global class gradients using local datasets (see Subsection 5.3). Approximation accuracy improves with increased data per class at each optimization step. As a result, the best performance is likely achieved with non-stochastic gradient descent, where honest workers compute gradients over the entire dataset. In practice, however, stochastic methods are more common—workers typically compute gradients on random mini-batches. To match this setting, we follow [3] and adopt the distributed stochastic heavy ball method (see Algorithm in Appendix B.1). This method has demonstrated strong robustness against Byzantine behavior in both heterogeneous [3] and homogeneous settings [13, 25, 16].
We use momentum $\beta=0.9$ for MNIST and Fashion-MNIST and $0.99$ for CIFAR10 and Purchase100.
7.7 Models & Hyperparameters
On MNIST and Fashion-MNIST, we train a lightweight CNN for $T=800$ steps. On CIFAR10, we use a deeper CNN; on Purchase100, a fully connected neural network. Both models are trained for $T=2000$ steps. The model architectures are provided in Appendix B.2. Unless otherwise specified, the loss is cross-entropy. The batch size is $b=128$, learning rates decay over time, gradient clipping is fixed at $5$, and $\ell_{2}$ regularization is set to $10^{-4}$.
7.8 Reproducibility
All experiments are run using the three seeds $\{1,2,3\}$ to ensure reproducibility. The code will be made publicly available for reuse at https://anonymous.4open.science/r/WoLA-E4CB.
8 Empirical Results
In this section, we present experiments to answer the following Research Questions (RQ). We use WoLA methods to refer collectively to WoLA, WoLA, and their combinations with NNM.
  • RQ1
    Does WoLA guide training toward a better model?
  • RQ2
    Are WoLA and WoLA more robust overall than state-of-the-art strategies?
  • RQ3
    Are WoLA and WoLA more robust in worst-case scenarios than state-of-the-art strategies?
  • RQ4
    Does combining WoLA or WoLA with NNM improve robustness?
  • RQ5
    What explains the robustness of WoLA methods?
  • RQ6
    How do batch size and momentum affect WoLA performance?
[…]
8.1 Targeted Model with WoLA (RQ1)
As expected from Proposition 3 (Subsection 6.2), WoLA does not target a better model during training.
Figure 5 shows Test Accuracy over a training without Byzantine workers, using simple mean aggregation. The model trained with WoLA performs similarly to the unweighted loss, though slightly worse on Purchase100 under high heterogeneity.
8.2 Robustness of WoLA and WoLA (RQ2)
WoLA and WoLA significantly improve robustness. They perform slightly below the best state-of-the-art methods when $f=2$, offer marginal gains at $f=4$, and outperform all non-WoLA methods as the number of Byzantine workers increases.
When $f=2$: WoLA and WoLA show similar performance, slightly below the best non-WoLA strategies. Typical negative gains between $-1.1$ and 00 points (Figure 4(a)). Notable exceptions exist, such as Purchase100 with high heterogeneity (Table 2, $\alpha=0.3$), where NNM outperforms WoLA and WoLA by more than $6$ points.
When $f=4$: WoLA and WoLA still perform similarly and yield slightly better results than non-WoLA methods, with typical gains of a few points (Figure 4(a)). According to Table 2, gains range from $+5.4$ (CIFAR10, $\alpha=0.3$) to $-2.9$ (Purchase100, $\alpha=0.3$).
When $f=6$: WoLA slightly surpasses WoLA, and both outperform other methods. WoLA shows typical gains between $+3.3$ and $+12.1$ points, while WoLA reaches between $+1.8$ and $+8.1$ (Figure 4(a)). Table 2 reports maximum gains of $+21.4$ (WoLA) and $+20.5$ (WoLA) on MNIST with $\alpha=0.3$.
When $f=8$: WoLA is better than WoLA, and both methods clearly outperform all others (Figure 4(a)). WoLA typically achieves gains between $+5.6$ and $+14.7$ points (median around $+10.1$, Figure 4(a)), while WoLA ranges from $-3.7$ to $+7$ (median around $+5.2$, Figure 4(a)). The best results appear on MNIST with medium heterogeneity ($\alpha=1$): $+17.3$ for WoLA and $+13.6$ for WoLA (Table 2).
Importantly, when $f>2$, Table 2 shows that WoLA almost always ranks in the top three regardless of the scenario (exceptions: Purchase100 $\alpha=0.3$, $f=4$ and CIFAR10 $\alpha=0.3$, $f=8$). While WoLA does not frequently appear among the top values, it outperforms non-WoLA strategies most of the time on MNIST and Fashion MNIST, often on Purchase100, and occasionally on CIFAR10.
8.3 Worst-Case Robustness of WoLA and WoLA (RQ3)
Considering the strongest attack for each approach, WoLA and WoLA consistently enhance robustness, almost always outperforming other state-of-the-art methods, with gains reaching up to $+41.9$ points (Table 3, MNIST, $\alpha=1$, $f=8$, WoLA).
For $f=2$: WoLA and WoLA perform similarly, slightly surpass non-WoLA methods with typical gains of a few points (Figure 4(b)), and up to $+11.3$ points on CIFAR10, $\alpha=0.3$ (Table 3).
For $f=4$: WoLA and WoLA continue to perform similarly, outperforming all baselines with median gains around $+6.3$ and $4.6$ points respectively (Figure 4(b)). According to Table 3, WoLA’s gains range from $+0.1$ (MNIST, $\alpha=3$) to $+28.8$ (CIFAR10, $\alpha=0.3$), while WoLA ranges from 00 (MNIST, $\alpha=3$) to $+27$ (CIFAR10, $\alpha=0.3$).
For $f=6$ or $8$: WoLA outperforms WoLA, and both clearly surpass all other methods (Figure 4(b)). WoLA achieves median gains around $+13.4$ points, WoLA around $+12.4$ (Figure 4(b)). Table 3 shows that WoLA’s gains range from $-0.8$ (MNIST, $\alpha=3$, $f=6$) to $+41.9$ (MNIST, $\alpha=1$, $f=8$); WoLA ranges from $-3.4$ (CIFAR10, $\alpha=0.3$, $f=8$) to $+41.3$ (MNIST, $\alpha=1$, $f=8$).
8.4 Robustness Gains from Combining WoLA and WoLA with NNM (RQ4)
Combining WoLA or WoLA with NNM significantly improves robustness, consistently outperforming WoLA, WoLA or NNM alone.
WoLA+NNM consistently achieves the top or top-three performance in Table 2 as soon as $f>2$, reflecting strong robustness across settings. Gains over the best non-WoLA methods can reach around $20$ points (Table 2). For example, on MNIST with $\alpha=0.3$ and $f=8$, the gain is $+26.6$; on Fashion MNIST with $\alpha=0.3$ and $f=6$, $+19.8$; on Purchase100 with $\alpha=1$ and $f=8$, $+25.2$; and on CIFAR10 with $\alpha=1$ and $f=6$, $+18.5$. The median gain is $+10.4$ when $f=6$ and $+17.8$ when $f=8$ (Figure 4(a)). Considering worst attacks, improvements are even larger, with median gains around $+20$ for both $f=6$ and $f=8$ (Figure 4(b)).
WoLA+NNM also regularly reaches top-three values in Tables 2 and 3, confirming its robustness. It performs similarly to WoLA+NNM when $f\leq 4$, and slightly worse when $f>4$, but still outperforms WoLA alone and all non-WoLA strategies (Figure 4(a), 4(b)). Importantly, it turns nearly all negative gains of WoLA alone into positive ones (Table 2). For instance, on CIFAR10 with $\alpha=3$ and $f=8$, the score improves from $-4.4$ to $+6.7$; with $\alpha=1$ and $f=8$, from $-3.6$ to $+8.3$; and with $\alpha=0.3$ and $f=2$, from $-0.5$ to $+2.2$. Overall, median gains over non-WoLA methods are around $+8$ when $f=6$ or $8$ (Figure 4(a)). Considering worst attacks, the median gain is around $+13$ for both $f=6$ and $f=8$ (Figure 4(b)).
8.5 Sources of Robustness in WoLA Methods (RQ5)
Gradient dissimilarity is a key indicator of system robustness [24, 3]. As expected from Proposition 2 (Subsection 6.2), WoLA methods significantly reduce gradient dissimilarity, thereby enhancing robustness against Byzantine attacks.
Table 4 shows that WoLA and its variants consistently achieve much lower gradient dissimilarity across all scenarios. Interestingly, combining NNM with WoLA further reduces gradient dissimilarity, which may explain its greater robustness over WoLA alone.
Figure 3 shows the evolution of Test Accuracy and Gradient Dissimilarity during training on Fashion-MNIST with high heterogeneity ($\alpha=0.3$) and six Byzantine workers ($f=6$) under two attacks (ALIE and LF). We also observe that WoLA’s methods maintain lower gradient dissimilarity than other state-of-the-art approaches.
(a) Fashion MNIST
(b) CIFAR10
Figure 6: Heatmaps of mean Test Accuracy over training (averaged over three seeds). Each row corresponds to a different level of data heterogeneity: low ($\alpha=3$), medium ($\alpha=1$), and high ($\alpha=0.3$). Each column shows: WoLA accuracy, NNM accuracy, and WoLA’s accuracy gain over NNM. In each heatmap, the vertical axis represents batch size, and the horizontal axis represents momentum. Both scenarios involve $17$ workers ($n=17$), including $5$ Byzantine workers ($f=5$), under the ALIE attack mitigated by the CwTM aggregator. Accuracy is reported in %, with a shared color scale for WoLA and NNM accuracy within each row.
8.6 Impact of Batch Size & Momentum on WoLA (RQ6)
Figure 6 reports the mean Test Accuracy on Fashion MNIST and CIFAR-10 achieved by WoLA under three heterogeneity levels ($\alpha=3$, $1$, and $0.3$), with $17$ workers including $5$ Byzantine nodes ($f=5$), using the ALIE attack and the CwTM aggregator. The figure also shows the performance of NNM and the accuracy gain of WoLA over NNM.
For Fashion MNIST, regardless of the heterogeneity level or momentum value, increasing the batch size consistently improves WoLA’s performance. Similarly, increasing the momentum coefficient improves performance up to $0.9$.
For CIFAR-10, under high heterogeneity ($\alpha=0.3$), increasing the batch size (up to $256$) or momentum coefficient improves performance. Under low and medium heterogeneity ($\alpha=3$ and $1$), the effects are less consistent. However, with high momentum ($0.9$ or $0.99$), increasing the batch size still tends to help. The best performance is observed with a high momentum coefficient ($0.99$).
Comparing WoLA and NNM, we find that the best results are obtained with moderately large batch sizes and high momentum. These conclusions are specific to the given setting and are not a substitute for a full-scale study, but they provide useful insights into parameter sensitivity.
9 Conclusion and Future Work
We show that label skew heterogeneity can be partly addressed with WoLA, a weighted loss where weights are set according to each worker’s local label distribution and a shared training objective. WoLA is a simple, model-agnostic method that requires no hyperparameter tuning, no server-side dataset, no knowledge of the number of Byzantine workers, and adds no computational overhead. It is easy to implement and can be combined with other robustness strategies, such as pre-aggregators and robust aggregators.
We provide theoretical insights into WoLA’s behavior, supported by empirical evaluation. Extensive experiments demonstrate WoLA’s efficiency and its superiority over state-of-the-art strategies—both on average and in worst-case scenarios (i.e., under the strongest attack for each method)—especially in adversarial settings involving both Byzantine behavior and heterogeneity. Even better results are achieved when WoLA is combined with NNM. We also show that WoLA significantly reduces gradient dissimilarity among honest workers, which is key to its robustness.
For future work, WoLA’s robustness can be improved and its scope extended—for instance, by refining the weighting scheme, such as introducing thresholds or adapting it dynamically at the mini-batch scale. The impact of training objective attacks can also be explored more thoroughly, along with the design of stronger defenses against them. Finally, developing a decentralized version of WoLA also seems to be a relevant direction.
Appendix A Theoretical Results
In this section, we define the random counterparts of the local updates and provide all proofs of the theoretical results given in Section 6.2.
A.1 Gradients Modeling
In this subsection, we define the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{WL}_{i}$. Let the global dataset be $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$ with $(x_{i},y_{i})\in\mathcal{X}\times[C]$, and let $z_{1},z_{2},\ldots,z_{N}$ be integers in $\mathcal{H}$ such that $z_{k}=i$ if $(x_{k},y_{k})\in\mathcal{D}_{i}$.
Since $N_{i}=\sum_{l=1}^{N}\mathbbm{1}(z_{l}=i)$, the following holds for $\nabla\mathcal{L}_{i}$:
$\nabla\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\ell\left(y,\Phi\left(x\right)\right)$
$=$ $\frac{N}{N_{i}}\frac{\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\ell\left(y_{k},\Phi\left(x_{k}\right)\right)}{N}$
$=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}$
Let $L_{i}$ be the random counterpart of $\nabla\mathcal{L}_{i}$. Replacing realizations with random variables (as defined in Subsection 6.1), $L_{i}$ is given by:
$L_{i}=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (5)
$\quad\quad\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
We now turn to $\nabla\mathcal{WL}_{i}$:
$\nabla\mathcal{WL}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\frac{q^{y_{k}}}{p^{y_{k}}_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\sum_{c=1}^{C}\mathbbm{1}(y_{k}=c)\frac{q^{c}}{N_{i}^{c}/N_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
$=$ $\sum_{c=1}^{C}q^{c}\frac{N}{N_{i}^{c}}\frac{\sum_{k=1}^{N}\mathbbm{1}(y_{k}=c)\mathbbm{1}(z_{k}=i)\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}}{N}$
Since $N^{c}_{i}=\sum_{l=1}^{N}\mathbbm{1}(y_{l}=c)\mathbbm{1}(z_{l}=i)$, we have:
$\nabla\mathcal{WL}_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(y_{l}=c)\mathbbm{1}(z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(y_{k}=c)\mathbbm{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}\Bigg{)}$
Let $W_{i}$ be the random counterpart of $\nabla\mathcal{WL}_{i}$. Replacing realizations with random variables (as defined in Subsection 6.1), $W_{i}$ is given by:
$W_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (6)
$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
A.2 Lemmas
In this subsection, we give two lemmas that will be used in the subsequent theoretical proofs.
Lemma 2
Let $X$, $Y$, and $Z$ be random variables such that $X\mid Y$ is independent of $Z\mid Y$. Let $f$ be a function defined on the space of realizations of $X$. Then, for all $y$ and $z$ such that $\mathbb{P}(Y=y)>0$ and $\mathbb{P}(Z=z)>0$, we have:
$\mathbb{E}\Big{[}f(X)\Big{|}Y=y,Z=z\Big{]}=\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}$
Proof.
$\mathbb{E}\Big{[}f(X)\Big{|}Y=y,Z=z\Big{]}=\frac{\mathbb{E}\Big{[}\mathbbm{1}(Y=y)\mathbbm{1}(Z=z)f(X)\Big{]}}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}\mathbbm{1}(Z=z)f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}\mathbbm{1}(Z=z)\Big{|}Y=y\Big{]}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\frac{\mathbb{E}\Big{[}\mathbbm{1}(Z=z)\mathbbm{1}(Y=y)\Big{]}}{\mathbb{P}(Y=y)}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}$
Lemma 3
The following expression converges almost surely (a.s.) as $N\to\infty$:
$\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
[…]
Proof.
By the continuity of the involved functions and the fact that the product of limits equals the limit of the product, it holds:
$\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$=\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
By the Law of Large Numbers, as $N\to\infty$, the expressions converge almost surely to:
$\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\Big{]}^{-1}$
$\quad\quad\quad\quad\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}$
$=$ $\frac{\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}}{\mathbb{P}(Y=c,Z=i)}$
$=$ $\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c,Z=i\Big{]}$
According to Lemma 2:
$\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c,Z=i\Big{]}=\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
A.3 Proof of Proposition 1
We begin by calculating the limit of $L_{i}$. By definition, we have:
$L_{i}=\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
By the Law of Large Numbers, it holds:
[…]
$\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
[…]
By the continuity of the involved functions and the fact that the product of limits equals the limit of the product, it holds:
$\lim_{N\to\infty}L_{i}=\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
$=$ $\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
[…]
Similarly, for $L$:
$\lim_{N\to\infty}L=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}L_{i}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}L_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
[…]
By the Law of total expectation, it holds:
[…]
A.4 Proof of Proposition 2
We begin by calculating the limit of $W_{i}$. By the continuity of the involved functions, it holds:
$\lim_{N\to\infty}W_{i}$
$=$ $\lim_{N\to\infty}\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$=$ $\sum_{c=1}^{C}q^{c}\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
Using Lemma 3, as $N\to\infty$, the expression converge almost surely to:
[…]
Similarly, for $W$:
$\lim_{N\to\infty}W=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}W_{i}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}W_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
By the Law of Large Numbers, it holds:
[…]
Using above results, it concludes the lemma:
[…]
$=$ $\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
A.5 Proof of Gradient Dissimilarity Vanishing
According to Proposition 2, by the continuity of the gradient dissimilarity formula, we obtain:
$\lim_{N\to\infty}\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\left\|W_{i}-W\right\|^{2}$
[…]
A.6 Proof of Proposition 3
According to Proposition 2, it holds:
[…]
By assumption, it holds $\tilde{p}_{c}=\mathbb{P}(Y=c)$ for all $c\in[C]$. Then, using the Law of total expectation, it follows:
[…]
$=\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{]}$
A.7 Worst Training Objective Attack
We show that the training objective $q$ is maximally deviated by a Byzantine attack that concentrates all its mass on the class with minimal weight among honest workers, under the $L_{1}$ norm denoted by $\|\cdot\|_{1}$.
Let $u$ and $v$ be two probability distributions over $C$ classes, with $u=(u^{1},u^{2},\ldots,u^{C})$ such that $u^{i}\in[0,1]$ and $\sum_{i=1}^{C}u^{i}=1$, and similarly for $v$. Let $u^{\min}$ and $u^{\max}$ denote the minimum and maximum components of $u$, and $v^{\max}$ the maximum of $v$.
We first recall the identity:
$\|u-v\|_{1}=\sum_{i=1}^{C}|u^{i}-v^{i}|$ $=\sum_{i=1}^{C}u^{i}+v^{i}-2\min(u^{i},v^{i})$
$=2-2\sum_{i=1}^{C}\min(u^{i},v^{i}).$
Observe that $v^{\max}\geq u^{\min}$. Otherwise, we would have $v^{i}<u^{i}$ for all $i$, which implies $\sum_{i}v^{i}<\sum_{i}u^{i}=1$, contradicting the assumption that $v$ is a probability distribution. It follows that:
$u^{\min}\leq\min(u^{\min},v^{\max})$ $\leq\min(u^{\max},v^{\max})$
$\leq\sum_{i=1}^{C}\min(u^{i},v^{i}),$
which gives the upper bound:
$\|u-v\|_{1}\leq 2-2u^{\min}.$
Now, consider the training objective $q=\sum_{i=1}^{n}q_{i}$, where each $q_{i}$ is a probability distribution sent by worker $i$. Let $u=\frac{1}{n-f}\sum_{i\in\mathcal{H}}q_{i}$ be the average contribution from honest workers, and $v=\frac{1}{f}\sum_{i\notin\mathcal{H}}q_{i}$ the average contribution from Byzantine workers. Then:
$\|q-u\|_{1}=\frac{f}{n}\|v-u\|_{1}\leq\frac{f}{n}(2-2u^{\min}).$
This bound is tight: it is achieved by setting each $q_{i}$ for $i\notin\mathcal{H}$ to put all mass on the class where $u$ is minimal, completing the proof.
Appendix B Experiment
B.1 Robust Distributed Stochastic Heavy Ball Algorithm
We present here the robust distributed stochastic Heavy Ball algorithm. More details can be found in [3]. We use superscript in brackets to denote the iteration index, i.e., (t), and subscripts to indicate the worker index.
Input: Initial model $\theta^{(0)}$, initial momentum $m_{i}^{(0)}=0$ for honest workers, robust aggregation $F$, learning rate $\gamma$, momentum coefficient $\beta$, and number of steps $T$.
for $t=1,\ldots,T$ do
       Server broadcasts $\theta^{(t-1)}$ to all workers;
       for every honest worker $i$ do
             Compute a stochastic gradient $g^{(t)}_{i}$;
             Update local momentum $m_{i}^{(t)}\leftarrow\beta m_{i}^{(t-1)}+(1-\beta)g^{(t)}_{i}$;
             Send $m_{i}^{(t)}$ to the server;
       end for
      Server aggregates the momentums:
           $R^{(t)}=F(m_{1}^{(t)},\ldots,m_{n}^{(t)})$;
       Server updates the model: $\theta^{(t)}=\theta^{(t-1)}-\gamma R^{(t)}$;
end for
Algorithm 1 Robust Distributed Stochastic Heavy Ball
B.2 Model Architectures
We adopt the following abreviations : L(#outputs) denotes a fully connected linear layer; R is a ReLU activation; T is a Tanh activation; C(#channels) represents a 2D convolutional layer with kernel size $5$, padding 00, and stride $1$; M denotes 2D max-pooling with kernel size $2$; B stands for batch normalization; and D represents dropout with a fixed probability of $0.25$.
The model architectures are as follows:
  • MNIST and Fashion MNIST (CNN from [3]): C(20)-R-M–C(20)-R-M–L(500)-R–L(10)
  • CIFAR10 (CNN from [3]):(3,32×32)-C(64)-R-B-C(64)-R-B-M-D-C(128)-R-B-C(128)-R-B-M-D-L(128)-R-D-L(10)
  • Purchase100 (fully connected neural networks from [31]): L(1024)-L(512)-T-L(256)-T-L(100)
The learning rates are defined as:
  • MNIST and Fashion MNIST (from [3]): $\gamma_{t}=0.75/(1+\left\lfloor\frac{t}{50}\right\rfloor)$
  • CIFAR10 (from [3]) and Purchase100: $\gamma_{t}=0.25$ if $t\leq 1500$, $0.025$ otherwise.
B.3 Mimic with Most Surrounded Outlier Selection
In the Mimic attack, we choose the honest outlier that is most surrounded by other honest updates. The selection procedure is detailed below. We first introduce the operators $\operatorname{dist}$ and $\operatorname{rank}$ used in this process.
Let $\mathbf{v}_{1},\ldots,\mathbf{v}_{h}$ be the honest update vectors. For each vector $\mathbf{v}_{i}$, define $\sigma(i,\cdot)$ as the permutation that orders the vectors in decreasing order of their Euclidean distance from $\mathbf{v}_{i}$:
$\operatorname{dist}(\mathbf{v}_{i},\mathbf{v}_{\sigma(i,1)})\geq\cdots\geq\operatorname{dist}(\mathbf{v}_{i},\mathbf{v}_{\sigma(i,h)}),$
where $\operatorname{dist}$ denotes the Euclidean distance. For any vector $\mathbf{v}_{j}$, there exists an index $r$ such that $\sigma(i,r)=j$, meaning that $\mathbf{v}_{j}$ is the $r$‑th farthest vector from $\mathbf{v}_{i}$. Note that $\sigma(i,h)=i$. We define the operator $\operatorname{rank}$ as:
$\operatorname{rank}(\mathbf{v}_{i},\mathbf{v}_{j})=r,$
where $j=\sigma(i,r)$.
The operators $\operatorname{dist}$ and $\operatorname{rank}$ allow us to compute a score for each update based on how surrounded a vector is and how far it is from the mean of the honest updates. The update to mimic is the one with the highest score, see Algorithm 2.
Input: Honest updates $\mathbf{v}_{1},\ldots,\mathbf{v}_{h}\in\mathbb{R}^{d}$, number of Byzantine workers $f$
Output: Update $\mathbf{v}_{\text{argmax}}$ with highest score
for $i=1,\ldots,h$ do
       $s_{i}\leftarrow 0$ ;
end for
$\textit{center}\leftarrow\frac{1}{h}\sum_{i=1}^{h}\mathbf{v}_{i}$ ;
for $i=1,\ldots,h$ do
       for $j=1,\ldots,h$ do
             $r\leftarrow\operatorname{rank}(\mathbf{v}_{i},\mathbf{v}_{j})$ ;
             $s_{j}\leftarrow s_{j}+\min(f,r)\cdot\operatorname{dist}(\mathbf{v}_{j},\textit{center})$ ;
       end for
end for
$k\leftarrow\arg\max_{j}s_{j}$ return $\mathbf{v}_{k}$
Algorithm 2 Mimic with most surrounded outlier selection
B.4 Comprehensive Results
We report the full experimental results. For MNIST, see Tables 5, 6, and 7; for Fashion MNIST, see Tables 8, 9, and 10; for Purchase100, see Tables 11, 12, and 13; and for CIFAR10, see Tables 14, 15, and 16. Each set of three tables corresponds to low, medium, and high heterogeneity levels ($\alpha=3$, $1$, and $0.3$).
Table 5: Accuracy (s.d.) % on MNIST with $\alpha=3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 93.7
93.2
92.4
84.8
93.3
92.5
91.4
85.5
93.6
92.8
88.9
52.7
93.6
93.1
92.6
91.2
93.2
92.5
91.6
80.5
93.5
92.8
91.4
78.9
93.2
92.5
88.9
52.7
BKT 93.7
93.6
92.4
84.8
93.6
93.3
91.4
85.5
93.3
92.1
88.9
52.7
93.7
93.4
92.6
91.2
93.5
92.8
91.6
80.5
93.6
93.0
91.4
78.9
93.3
92.1
88.9
52.7
FoundFL 93.0
93.3
93.1
92.5
91.0
85.8
12.1
12.3
90.1
73.2
50.7
42.9
92.8
92.5
93.3
93.3
90.7
82.5
10.2
11.0
91.5
85.5
51.9
50.4
90.1
73.2
10.2
11.0
NNM 93.8
93.5
93.2
93.7
93.7
93.4
92.7
77.4
93.8
93.5
93.6
93.3
93.6
93.3
92.3
90.6
93.6
92.7
91.7
61.0
93.7
93.3
92.7
83.2
93.6
92.7
91.7
61.0
WoLA 93.7
93.5
93.1
91.5
93.5
93.0
92.1
90.8
93.8
93.6
93.3
91.0
93.7
93.6
93.3
93.5
93.5
92.9
92.0
90.8
93.6
93.3
92.8
91.5
93.5
92.9
92.0
90.8
WoLA+NNM 93.8
93.6
93.4
93.8
93.7
93.4
93.0
90.9
93.8
93.7
93.8
93.8
93.8
93.6
93.4
93.4
93.6
93.2
92.6
92.2
93.7
93.5
93.2
92.8
93.6
93.2
92.6
90.9
WoLA 93.7
93.4
92.8
90.9
93.5
93.0
91.9
90.5
93.7
93.5
92.8
90.5
93.7
93.5
93.3
92.7
93.4
92.8
91.8
90.7
93.6
93.2
92.5
91.1
93.4
92.8
91.8
90.5
WoLA+NNM 93.8
93.4
93.1
93.2
93.7
93.3
92.9
90.5
93.8
93.6
93.4
92.9
93.7
93.5
93.3
92.7
93.6
93.0
92.3
91.0
93.7
93.4
93.0
92.1
93.6
93.0
92.3
90.5
CwTM - 93.7
93.5
92.7
84.8
93.6
93.2
91.3
85.5
93.6
92.6
88.6
54.1
93.5
93.2
92.5
90.2
93.6
92.6
91.4
80.7
93.6
93.0
91.3
79.0
93.5
92.6
88.6
54.1
BKT 93.7
93.6
92.7
84.8
93.6
93.3
91.3
85.5
93.3
92.1
88.6
54.1
93.6
93.3
92.5
90.2
93.5
92.8
91.4
80.7
93.6
93.0
91.3
79.0
93.3
92.1
88.6
54.1
FoundFL 92.6
92.9
92.7
92.3
92.7
91.6
90.4
19.6
72.7
48.9
26.0
16.6
93.0
92.8
93.1
92.9
90.9
24.5
10.8
11.2
88.4
70.1
62.6
46.5
72.7
24.5
10.8
11.2
NNM 93.8
93.6
93.1
93.7
93.7
93.4
92.7
75.8
93.8
93.5
93.7
93.3
93.6
93.3
92.8
91.0
93.6
92.8
91.7
59.6
93.7
93.3
92.8
82.7
93.6
92.8
91.7
59.6
WoLA 93.8
93.5
93.2
91.4
93.7
93.3
92.5
90.7
93.7
93.5
93.2
91.0
93.7
93.6
93.5
93.4
93.5
92.9
91.9
90.9
93.7
93.4
92.9
91.5
93.5
92.9
91.9
90.7
WoLA+NNM 93.8
93.6
93.3
93.8
93.7
93.4
93.0
91.0
93.8
93.7
93.8
93.8
93.8
93.6
93.5
93.4
93.6
93.2
92.6
92.2
93.7
93.5
93.2
92.8
93.6
93.2
92.6
91.0
WoLA 93.7
93.5
92.9
90.9
93.6
93.2
92.3
90.5
93.7
93.4
92.7
90.3
93.7
93.5
93.4
92.7
93.5
92.9
91.7
90.4
93.7
93.3
92.6
90.9
93.5
92.9
91.7
90.3
WoLA+NNM 93.8
93.4
93.1
93.3
93.7
93.3
92.9
90.4
93.8
93.6
93.4
92.9
93.7
93.5
93.4
92.7
93.6
93.1
92.3
91.4
93.7
93.4
93.0
92.1
93.6
93.1
92.3
90.4
MKrum - 93.8
93.6
93.1
78.3
93.7
93.4
92.6
57.0
93.8
93.4
43.3
1.7
93.6
93.1
92.6
90.6
93.5
92.3
30.6
10.1
93.7
93.1
70.4
47.5
93.5
92.3
30.6
1.7
BKT 93.7
93.3
93.1
78.3
93.4
92.9
92.6
57.0
92.4
89.0
43.3
1.7
93.5
92.9
92.6
90.6
93.4
91.5
30.6
10.1
93.3
91.9
70.4
47.5
92.4
89.0
30.6
1.7
FoundFL 92.5
92.7
91.1
91.0
92.4
92.2
85.5
27.1
69.9
52.2
33.9
14.5
92.9
92.3
91.8
92.0
86.3
10.7
10.1
10.6
86.8
68.0
62.5
47.0
69.9
10.7
10.1
10.6
NNM 93.7
93.6
93.1
83.1
93.7
93.5
92.7
71.6
93.8
93.5
93.6
8.0
93.7
93.2
92.2
90.7
93.6
92.7
90.6
10.1
93.7
93.3
92.5
52.7
93.6
92.7
90.6
8.0
WoLA 93.8
93.6
93.3
91.0
93.7
93.3
92.7
76.8
93.8
93.6
93.8
1.7
93.7
93.5
93.4
93.4
93.6
92.9
87.5
10.1
93.7
93.4
92.2
54.6
93.6
92.9
87.5
1.7
WoLA+NNM 93.8
93.6
93.3
91.8
93.7
93.4
93.0
89.1
93.8
93.7
93.8
9.3
93.8
93.6
93.4
93.4
93.6
93.2
92.6
10.4
93.7
93.5
93.2
58.8
93.6
93.2
92.6
9.3
WoLA 93.8
93.4
93.1
90.3
93.7
93.3
92.6
70.0
93.8
93.5
93.4
4.6
93.7
93.4
93.4
92.7
93.6
92.8
87.8
10.1
93.7
93.3
92.1
53.6
93.6
92.8
87.8
4.6
WoLA+NNM 93.7
93.4
93.1
91.0
93.7
93.3
92.9
88.2
93.8
93.6
93.4
13.7
93.7
93.5
93.4
92.8
93.6
93.1
92.2
10.1
93.7
93.4
93.0
59.1
93.6
93.1
92.2
10.1
GM - 93.8
93.6
93.0
87.1
93.5
90.3
86.1
36.4
93.8
93.5
93.6
1.4
93.8
93.2
92.3
91.5
93.3
92.2
90.5
81.5
93.6
92.5
91.1
59.6
93.3
90.3
86.1
1.4
BKT 93.8
93.6
93.0
87.1
93.6
93.4
86.1
36.4
93.6
93.2
93.6
1.4
93.6
93.4
92.3
91.5
93.6
92.9
90.5
81.5
93.6
93.3
91.1
59.6
93.6
92.9
86.1
1.4
FoundFL 92.9
92.9
92.6
92.8
91.0
72.7
12.7
11.0
92.3
37.8
42.8
22.8
92.6
92.7
92.7
92.9
88.8
73.7
10.7
11.7
91.5
74.0
50.3
46.3
88.8
37.8
10.7
11.0
NNM 93.8
93.6
93.2
93.3
93.7
93.4
92.9
79.6
93.8
93.5
93.7
93.2
93.7
93.4
92.6
91.0
93.6
92.9
91.4
59.5
93.7
93.4
92.7
83.3
93.6
92.9
91.4
59.5
WoLA 93.8
93.6
93.5
92.2
93.7
93.3
92.9
87.6
93.8
93.6
93.6
93.5
93.8
93.6
93.5
93.6
93.5
92.7
90.9
83.6
93.7
93.4
92.9
90.1
93.5
92.7
90.9
83.6
WoLA+NNM 93.8
93.6
93.4
93.4
93.7
93.4
93.1
91.0
93.8
93.7
93.7
93.6
93.8
93.5
93.5
93.5
93.7
93.1
92.6
91.2
93.7
93.5
93.2
92.5
93.7
93.1
92.6
91.0
WoLA 93.8
93.5
93.1
91.7
93.6
93.2
92.5
87.9
93.7
93.6
93.2
92.7
93.7
93.5
93.2
92.8
93.5
92.7
90.9
83.9
93.6
93.3
92.6
89.8
93.5
92.7
90.9
83.9
WoLA+NNM 93.8
93.5
93.1
92.9
93.7
93.3
92.8
89.9
93.8
93.5
93.3
92.9
93.7
93.5
93.4
92.7
93.6
93.0
92.2
90.6
93.7
93.4
93.0
91.8
93.6
93.0
92.2
89.9
Table 6: Accuracy (s.d.) % on MNIST with $\alpha=1$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 93.9
93.1
90.6
55.4
93.2
91.7
86.7
62.4
93.2
91.0
68.4
21.5
93.2
92.6
89.8
79.9
92.6
90.9
86.1
35.1
93.2
91.8
84.3
50.9
92.6
90.9
68.4
21.5
BKT 93.9
93.6
90.6
55.4
93.8
93.3
86.7
62.4
93.2
91.5
68.4
21.5
93.8
92.8
89.8
79.9
93.6
92.6
86.1
35.1
93.7
92.8
84.3
50.9
93.2
91.5
68.4
21.5
FoundFL 86.7
91.4
91.4
89.5
84.2
64.8
23.9
10.0
80.7
78.5
78.9
50.7
82.7
87.8
92.5
92.4
81.4
30.8
11.3
13.6
83.1
70.6
59.6
51.2
80.7
30.8
11.3
10.0
NNM 94.0
93.6
92.6
93.1
93.9
93.5
92.2
58.3
93.9
93.7
93.7
92.4
93.9
93.0
87.0
73.7
93.8
90.8
87.2
32.1
93.9
92.9
90.5
69.9
93.8
90.8
87.0
32.1
WoLA 94.0
93.6
93.1
89.8
93.7
93.1
91.8
88.0
93.8
93.7
93.0
88.8
94.0
93.7
93.5
93.3
93.7
92.9
92.0
87.3
93.8
93.4
92.7
89.4
93.7
92.9
91.8
87.3
WoLA+NNM 94.1
93.7
93.2
93.9
93.9
93.4
92.9
84.7
94.1
93.9
93.9
93.9
94.1
93.7
93.5
93.1
93.9
93.2
92.4
91.3
94.0
93.6
93.2
91.4
93.9
93.2
92.4
84.7
WoLA 93.9
93.5
92.8
85.9
93.6
93.0
91.8
89.2
93.8
93.4
92.1
86.7
93.8
93.6
92.8
85.9
93.6
92.9
92.2
88.7
93.7
93.3
92.3
87.3
93.6
92.9
91.8
85.9
WoLA+NNM 93.9
93.6
93.2
91.7
93.8
93.5
92.7
72.3
94.0
93.8
93.4
92.4
93.9
93.5
92.9
86.6
93.7
93.1
91.8
65.1
93.9
93.5
92.8
81.6
93.7
93.1
91.8
65.1
CwTM - 94.1
93.5
91.4
56.4
93.9
93.1
88.1
63.5
93.6
91.9
74.9
20.5
93.7
92.3
90.4
78.3
93.8
91.5
86.5
39.1
93.8
92.5
86.3
51.6
93.6
91.5
74.9
20.5
BKT 93.9
93.6
91.4
56.4
93.8
93.3
88.1
63.5
93.2
91.5
74.9
20.5
93.8
92.9
90.4
78.3
93.6
92.6
86.5
39.1
93.7
92.8
86.3
51.6
93.2
91.5
74.9
20.5
FoundFL 91.9
91.1
90.0
89.7
91.6
87.8
80.0
20.6
79.6
48.5
55.4
21.5
90.1
90.7
91.3
91.5
88.1
19.3
10.6
14.3
88.3
67.5
65.5
47.5
79.6
19.3
10.6
14.3
NNM 94.0
93.7
92.5
93.1
94.0
93.5
92.4
59.0
94.0
93.7
93.8
92.3
93.9
92.5
90.9
76.5
93.9
92.4
87.0
25.0
93.9
93.2
91.3
69.2
93.9
92.4
87.0
25.0
WoLA 94.1
93.7
93.1
89.6
93.9
93.4
92.4
88.2
93.8
93.6
92.8
89.2
94.0
93.6
93.4
93.1
93.8
93.0
91.6
86.6
93.9
93.5
92.7
89.3
93.8
93.0
91.6
86.6
WoLA+NNM 94.1
93.8
93.3
93.9
93.9
93.4
92.9
85.5
94.1
93.9
93.9
93.9
94.0
93.6
93.5
93.2
93.9
93.2
92.2
91.8
94.0
93.6
93.2
91.7
93.9
93.2
92.2
85.5
WoLA 94.0
93.6
92.9
85.5
93.8
93.3
92.2
89.2
93.8
93.3
92.3
86.2
93.9
93.6
92.7
84.6
93.7
92.9
91.7
88.5
93.8
93.4
92.3
86.8
93.7
92.9
91.7
84.6
WoLA+NNM 93.9
93.7
93.2
91.7
93.8
93.5
92.8
71.9
94.0
93.8
93.4
92.4
93.9
93.5
92.9
83.3
93.8
93.1
92.0
65.0
93.9
93.5
92.9
80.8
93.8
93.1
92.0
65.0
MKrum - 93.9
93.6
92.4
32.4
93.8
93.5
92.1
56.4
93.8
93.5
49.4
11.4
93.8
92.6
90.7
73.7
93.7
90.6
30.6
6.5
93.8
92.7
71.0
36.1
93.7
90.6
30.6
6.5
BKT 93.7
93.3
92.4
32.4
93.4
92.8
92.1
56.4
92.7
87.4
49.4
11.4
93.4
92.1
90.7
73.7
93.1
90.1
30.6
6.5
93.3
91.1
71.0
36.1
92.7
87.4
30.6
6.5
FoundFL 91.6
91.2
88.7
77.3
91.7
90.0
55.2
29.9
85.8
72.8
53.8
29.3
91.8
90.9
90.2
86.8
82.3
12.8
11.0
14.2
88.6
71.6
59.8
47.5
82.3
12.8
11.0
14.2
NNM 94.1
93.6
92.4
39.9
94.0
93.5
92.2
70.7
93.9
93.7
93.5
26.1
93.9
92.7
74.6
59.9
93.9
92.0
81.4
6.5
93.9
93.1
86.8
40.6
93.9
92.0
74.6
6.5
WoLA 94.1
93.8
93.3
89.0
93.9
93.4
92.7
76.2
94.1
93.9
93.7
12.5
94.0
93.7
93.4
93.1
93.8
92.9
85.6
6.5
94.0
93.5
91.7
55.5
93.8
92.9
85.6
6.5
WoLA+NNM 94.1
93.8
93.2
90.5
93.9
93.5
92.9
71.4
94.1
93.9
93.9
23.9
94.1
93.7
93.6
93.1
93.9
93.2
92.2
6.7
94.0
93.6
93.2
57.1
93.9
93.2
92.2
6.7
WoLA 93.9
93.6
93.1
83.2
93.7
93.4
92.6
49.2
94.0
93.8
93.3
8.6
93.9
93.5
92.8
86.6
93.7
92.7
84.5
12.3
93.8
93.4
91.3
48.0
93.7
92.7
84.5
8.6
WoLA+NNM 93.9
93.7
93.1
85.9
93.8
93.4
92.8
58.1
94.0
93.8
93.4
17.7
93.8
93.5
92.8
86.3
93.8
93.1
91.7
12.3
93.9
93.5
92.8
52.0
93.8
93.1
91.7
12.3
GM - 94.0
93.6
92.4
51.0
93.1
87.4
77.2
24.6
93.9
93.7
93.6
2.1
93.9
92.9
84.6
31.2
93.2
91.9
89.5
68.4
93.6
91.9
87.5
35.5
93.1
87.4
77.2
2.1
BKT 94.0
93.7
92.4
51.0
93.9
93.5
77.2
24.6
93.7
93.1
93.6
2.1
93.9
93.3
84.6
31.2
93.7
92.8
89.5
68.4
93.8
93.3
87.5
35.5
93.7
92.8
77.2
2.1
FoundFL 91.0
92.7
91.7
90.6
62.7
49.6
10.9
17.2
86.1
69.7
24.0
61.7
87.3
85.5
91.2
91.9
81.2
45.7
11.2
15.5
81.7
68.6
45.8
55.4
62.7
45.7
10.9
15.5
NNM 94.1
93.7
92.7
92.7
94.0
93.6
92.8
72.7
94.0
93.7
93.8
90.9
93.8
92.7
91.5
82.9
93.9
91.1
87.2
26.2
93.9
93.0
91.6
73.1
93.8
91.1
87.2
26.2
WoLA 94.1
93.8
93.4
91.0
93.9
93.4
92.3
77.0
93.9
93.8
93.5
93.2
94.0
93.6
93.5
93.4
93.7
92.8
90.5
83.5
93.9
93.5
92.7
87.6
93.7
92.8
90.5
77.0
WoLA+NNM 94.1
93.8
93.3
93.6
93.9
93.5
93.1
90.1
93.9
93.9
93.6
93.5
94.0
93.6
93.5
93.3
93.8
93.2
92.3
90.7
94.0
93.6
93.2
92.2
93.8
93.2
92.3
90.1
WoLA 94.0
93.6
93.0
87.0
93.8
93.3
92.0
75.7
93.8
93.5
93.4
92.2
93.9
93.5
92.7
91.5
93.7
92.7
90.6
79.8
93.8
93.3
92.3
85.3
93.7
92.7
90.6
75.7
WoLA+NNM 93.9
93.7
93.2
91.0
93.8
93.4
92.9
64.5
93.8
93.5
93.5
92.7
93.9
93.6
92.9
87.0
93.8
93.1
91.9
63.2
93.9
93.5
92.9
79.7
93.8
93.1
91.9
63.2
Table 7: Accuracy (s.d.) % on MNIST with $\alpha=0.3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 93.1
91.1
65.2
15.0
90.0
85.8
70.4
16.0
82.9
68.9
40.1
8.4
89.3
76.8
55.5
12.6
80.5
60.5
50.3
16.2
87.2
76.6
56.3
13.6
80.5
60.5
40.1
8.4
BKT 93.8
93.0
65.2
15.0
93.6
92.4
70.4
16.0
92.8
87.6
40.1
8.4
93.6
90.7
55.5
12.6
93.4
86.9
50.3
16.2
93.4
90.1
56.3
13.6
92.8
86.9
40.1
8.4
FoundFL 66.5
78.7
66.3
69.5
30.1
23.7
12.0
18.3
55.0
43.5
43.8
21.4
50.4
39.5
50.8
67.4
36.1
31.5
16.5
15.5
47.6
43.4
37.9
38.4
30.1
23.7
12.0
15.5
NNM 94.0
93.7
79.4
92.8
93.9
93.4
84.7
29.9
94.1
93.8
65.8
25.4
93.7
90.4
79.2
35.7
93.7
80.0
31.0
26.9
93.9
90.3
68.0
42.1
93.7
80.0
31.0
25.4
WoLA 93.7
93.4
92.0
68.9
93.2
92.4
87.0
49.1
93.4
92.9
88.9
41.3
93.5
93.3
92.8
91.4
93.0
92.0
88.1
42.5
93.4
92.8
89.8
58.6
93.0
92.0
87.0
41.3
WoLA+NNM 93.8
93.7
92.6
91.6
93.7
93.4
91.2
22.6
93.7
94.0
94.0
93.6
93.6
93.1
92.7
91.4
93.6
92.5
75.0
46.9
93.7
93.3
89.1
69.2
93.6
92.5
75.0
22.6
WoLA 93.7
93.2
91.7
57.1
93.2
92.2
86.9
65.5
93.4
92.7
87.7
36.4
93.5
92.9
91.3
56.2
93.1
92.1
85.7
61.7
93.4
92.6
88.7
55.4
93.1
92.1
85.7
36.4
WoLA+NNM 93.8
93.6
92.5
87.1
93.7
93.3
91.3
27.7
93.7
93.9
93.0
71.5
93.6
93.1
91.8
57.5
93.5
92.4
73.4
15.8
93.7
93.3
88.4
51.9
93.5
92.4
73.4
15.8
CwTM - 93.9
92.4
75.5
14.5
93.5
90.2
62.3
16.7
93.4
80.9
35.6
8.7
93.4
90.1
60.4
29.1
93.3
69.6
42.1
17.8
93.5
84.6
55.2
17.4
93.3
69.6
35.6
8.7
BKT 93.8
93.0
75.5
14.5
93.6
92.4
62.3
16.7
92.8
87.8
35.6
8.7
93.5
91.1
60.4
29.1
93.4
87.2
42.1
17.8
93.4
90.3
55.2
17.4
92.8
87.2
35.6
8.7
FoundFL 79.0
73.4
78.9
69.4
81.0
63.3
27.9
20.0
53.5
37.6
21.6
22.7
56.9
81.6
76.5
68.9
56.5
18.0
13.9
17.4
65.4
54.8
43.8
39.7
53.5
18.0
13.9
17.4
NNM 94.1
93.6
75.1
92.8
93.9
93.5
85.2
22.4
94.2
93.9
58.5
25.3
93.8
91.1
56.5
27.7
93.7
87.7
55.5
35.8
93.9
92.0
66.1
40.8
93.7
87.7
55.5
22.4
WoLA 93.8
93.7
92.4
69.3
93.6
93.3
91.3
49.2
93.6
93.2
91.1
44.9
93.6
93.3
92.8
89.7
93.5
92.5
90.4
43.8
93.6
93.2
91.6
59.4
93.5
92.5
90.4
43.8
WoLA+NNM 93.8
93.7
92.7
92.2
93.7
93.4
91.9
26.2
93.7
93.9
94.0
93.6
93.6
93.2
92.3
91.1
93.6
92.6
83.6
50.5
93.7
93.4
90.9
70.7
93.6
92.6
83.6
26.2
WoLA 93.8
93.6
92.1
57.3
93.7
93.2
90.8
65.2
93.6
92.7
91.0
36.4
93.7
93.3
91.1
60.7
93.5
92.4
89.8
62.0
93.6
93.0
91.0
56.3
93.5
92.4
89.8
36.4
WoLA+NNM 93.8
93.7
92.6
86.3
93.7
93.3
91.4
28.9
93.7
93.9
93.0
70.4
93.6
93.3
91.2
57.5
93.6
92.7
88.8
23.9
93.7
93.4
91.4
53.4
93.6
92.7
88.8
23.9
MKrum - 94.0
93.6
77.7
5.3
93.9
93.6
85.0
31.7
94.1
73.7
55.6
31.6
93.4
91.6
79.2
35.7
93.6
86.5
15.7
6.0
93.8
87.8
62.6
22.0
93.4
73.7
15.7
5.3
BKT 93.3
92.9
77.7
5.3
93.0
91.9
85.0
31.7
91.2
78.0
55.6
31.6
92.9
86.7
79.2
35.7
92.6
81.3
15.7
6.0
92.6
86.2
62.6
22.0
91.2
78.0
15.7
5.3
FoundFL 82.3
67.4
46.1
30.2
76.3
73.3
64.5
21.8
87.0
52.5
34.0
37.1
85.1
74.1
75.4
49.8
74.5
16.3
16.5
17.4
81.0
56.7
47.3
31.3
74.5
16.3
16.5
17.4
NNM 94.1
93.6
78.9
6.9
93.9
93.3
85.1
24.0
94.1
93.8
57.4
24.3
93.9
91.5
55.1
30.3
93.7
85.7
53.7
6.0
93.9
91.6
66.0
18.3
93.7
85.7
53.7
6.0
WoLA 93.9
93.6
92.6
67.9
93.7
93.3
91.4
50.8
93.7
94.0
93.6
18.3
93.6
93.2
92.5
91.4
93.5
92.2
48.4
5.1
93.7
93.2
83.7
46.7
93.5
92.2
48.4
5.1
WoLA+NNM 93.8
93.7
92.6
73.0
93.7
93.3
91.6
34.6
93.7
93.9
94.0
24.2
93.6
93.1
92.4
91.3
93.6
92.5
85.7
5.1
93.7
93.3
91.3
45.6
93.6
92.5
85.7
5.1
WoLA 93.8
93.6
92.3
39.3
93.7
93.3
91.3
33.5
93.7
93.9
93.1
15.4
93.5
93.0
91.3
57.5
93.5
92.1
42.4
7.5
93.7
93.2
82.1
30.7
93.5
92.1
42.4
7.5
WoLA+NNM 93.8
93.6
92.4
55.7
93.7
93.3
91.0
17.7
93.7
93.9
93.0
21.0
93.7
93.2
91.6
57.7
93.6
92.6
84.1
7.6
93.7
93.3
90.4
31.9
93.6
92.6
84.1
7.6
GM - 94.0
93.4
86.9
7.3
90.1
79.0
48.2
6.2
94.0
93.5
66.1
9.6
94.0
89.6
26.4
47.8
92.8
91.4
76.8
10.0
93.0
89.4
60.9
16.2
90.1
79.0
26.4
6.2
BKT 94.0
93.8
86.9
7.3
93.8
93.4
48.2
6.2
93.6
86.2
66.1
9.6
93.8
92.2
26.4
47.8
93.6
92.6
76.8
10.0
93.8
91.6
60.9
16.2
93.6
86.2
26.4
6.2
FoundFL 64.3
78.2
78.9
59.7
52.1
24.1
24.0
23.7
35.9
43.0
34.7
31.2
40.3
56.9
60.7
81.2
69.6
31.2
16.5
17.4
52.4
46.7
42.9
42.6
35.9
24.1
16.5
17.4
NNM 94.1
93.7
83.2
88.7
94.0
93.5
86.1
38.0
94.1
93.9
58.9
29.6
93.8
91.0
63.1
18.6
93.7
89.0
58.2
22.3
93.9
92.2
69.9
39.4
93.7
89.0
58.2
18.6
WoLA 93.8
93.8
92.9
76.7
93.5
92.1
84.8
37.1
93.7
93.8
93.3
24.9
93.6
93.3
92.2
92.0
93.3
91.6
89.4
36.6
93.6
92.9
90.5
53.4
93.3
91.6
84.8
24.9
WoLA+NNM 93.7
93.7
93.0
93.1
93.7
93.5
92.8
42.3
93.7
93.8
93.5
93.3
93.6
93.3
92.9
92.1
93.6
92.7
71.5
43.8
93.7
93.4
88.7
72.9
93.6
92.7
71.5
42.3
WoLA 93.7
93.6
92.4
71.2
93.5
92.1
85.8
27.2
93.7
93.6
92.9
17.8
93.7
93.4
91.6
86.4
93.3
91.4
89.9
31.1
93.6
92.8
90.5
46.7
93.3
91.4
85.8
17.8
WoLA+NNM 93.8
93.7
92.7
86.9
93.7
93.4
92.2
31.1
93.7
93.7
93.0
68.1
93.6
93.2
92.3
89.7
93.6
92.7
74.4
12.2
93.7
93.4
88.9
57.6
93.6
92.7
74.4
12.2
Table 8: Accuracy (s.d.) % on Fashion MNIST with $\alpha=3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 78.0
75.1
70.5
61.4
74.2
70.6
65.3
55.3
75.5
71.3
61.3
32.2
75.9
73.8
71.2
70.7
74.1
71.2
65.2
45.2
75.5
72.4
66.7
53.0
74.1
70.6
61.3
32.2
BKT 77.7
76.6
70.5
61.4
77.0
75.5
65.3
55.3
76.0
72.4
61.3
32.2
76.9
75.9
71.2
70.7
76.5
72.8
65.2
45.2
76.8
74.7
66.7
53.0
76.0
72.4
61.3
32.2
FoundFL 72.1
73.2
73.1
69.6
70.0
58.2
11.4
11.7
70.0
60.3
58.4
44.5
71.5
72.4
73.4
74.8
66.1
52.9
11.7
14.1
69.9
63.4
45.6
42.9
66.1
52.9
11.4
11.7
NNM 79.0
77.4
74.1
76.4
78.3
76.8
73.1
30.7
79.3
79.2
79.1
79.3
77.6
75.2
72.0
70.8
77.8
72.6
62.1
38.7
78.4
76.2
72.1
59.2
77.6
72.6
62.1
30.7
WoLA 78.8
78.3
77.1
73.1
77.9
76.6
74.1
71.9
78.9
78.2
77.1
74.1
78.7
77.9
78.0
77.5
78.1
76.6
74.3
71.9
78.5
77.5
76.1
73.7
77.9
76.6
74.1
71.9
WoLA+NNM 79.0
78.6
78.0
78.7
78.3
77.4
75.8
71.5
78.9
79.2
79.2
79.1
78.7
77.9
77.7
77.1
78.0
76.6
75.1
74.8
78.6
77.9
77.2
76.2
78.0
76.6
75.1
71.5
WoLA 78.9
77.8
76.0
68.1
77.9
76.0
72.4
67.5
78.9
77.9
75.5
68.2
78.3
77.9
76.8
72.4
77.9
76.1
72.8
67.1
78.4
77.1
74.7
68.7
77.9
76.0
72.4
67.1
WoLA+NNM 78.7
78.6
76.5
74.9
78.6
77.4
74.6
63.7
78.9
78.6
77.9
75.6
78.5
77.4
77.0
73.1
77.7
75.9
73.1
67.8
78.5
77.6
75.8
71.0
77.7
75.9
73.1
63.7
CwTM - 78.8
77.0
72.4
58.8
78.0
75.0
66.4
55.2
77.6
72.1
63.7
29.8
77.4
75.4
72.7
70.9
77.2
70.9
67.3
46.2
77.8
74.1
68.5
52.2
77.2
70.9
63.7
29.8
BKT 77.6
76.6
72.4
58.8
77.0
75.5
66.4
55.2
75.8
72.4
63.7
29.8
76.6
75.9
72.7
70.9
76.4
72.9
67.3
46.2
76.7
74.7
68.5
52.2
75.8
72.4
63.7
29.8
FoundFL 75.0
73.7
73.2
69.7
74.0
72.2
66.6
18.8
68.6
59.9
54.1
36.3
75.0
74.2
74.3
74.9
68.9
21.5
13.2
13.2
72.3
60.3
56.3
42.6
68.6
21.5
13.2
13.2
NNM 79.1
77.6
73.9
76.3
78.6
77.1
73.1
30.7
79.3
79.4
79.1
79.3
77.7
75.6
72.2
71.8
78.1
73.7
63.8
38.7
78.6
76.7
72.4
59.4
77.7
73.7
63.8
30.7
WoLA 79.2
78.7
77.4
72.5
78.8
77.4
74.4
71.9
79.0
78.2
75.9
74.4
79.0
78.2
77.6
77.7
78.2
76.3
73.9
71.7
78.8
77.7
75.9
73.6
78.2
76.3
73.9
71.7
WoLA+NNM 79.1
78.8
78.1
78.7
78.5
77.3
75.7
71.7
78.8
79.3
79.2
79.0
78.5
77.9
77.9
77.5
78.1
76.7
74.7
74.3
78.6
78.0
77.1
76.3
78.1
76.7
74.7
71.7
WoLA 79.0
77.9
76.2
68.5
78.7
77.0
73.2
67.3
78.9
77.8
75.3
69.0
78.8
77.8
76.6
71.4
78.2
75.9
72.1
67.3
78.7
77.3
74.7
68.7
78.2
75.9
72.1
67.3
WoLA+NNM 79.1
78.6
76.4
74.7
78.9
77.5
74.6
64.8
78.9
78.6
77.9
75.6
78.2
77.5
76.9
72.6
77.8
75.8
73.5
68.0
78.6
77.6
75.9
71.1
77.8
75.8
73.5
64.8
MKrum - 78.5
77.0
73.5
52.8
78.3
76.4
72.6
21.4
79.4
78.8
37.8
1.7
77.1
75.3
73.1
70.8
77.7
72.0
21.4
8.5
78.2
75.9
55.7
31.0
77.1
72.0
21.4
1.7
BKT 77.4
75.5
73.5
52.8
76.2
73.9
72.6
21.4
75.7
73.0
37.8
1.7
76.9
73.7
73.1
70.8
75.2
67.5
21.4
8.5
76.3
72.7
55.7
31.0
75.2
67.5
21.4
1.7
FoundFL 74.4
74.2
71.6
67.3
74.4
73.4
66.8
25.6
68.7
61.0
75.6
10.6
74.3
74.7
72.5
73.6
64.6
22.4
11.0
41.9
71.3
61.1
59.5
43.8
64.6
22.4
11.0
10.6
NNM 79.0
77.7
73.7
59.0
78.5
76.9
72.9
44.2
79.4
79.3
78.7
4.8
77.3
75.7
72.1
69.8
77.9
73.5
61.8
8.5
78.4
76.6
71.8
37.3
77.3
73.5
61.8
4.8
WoLA 78.8
78.5
77.8
72.5
78.4
77.6
74.6
23.7
79.0
79.2
79.1
1.8
78.5
77.9
77.6
77.1
78.1
75.6
68.5
9.1
78.6
77.8
75.5
36.8
78.1
75.6
68.5
1.8
WoLA+NNM 79.0
79.0
77.9
74.2
78.3
77.7
75.5
65.0
79.0
79.3
79.2
5.1
79.0
77.8
77.6
77.5
78.1
76.7
74.9
8.5
78.7
78.1
77.0
46.0
78.1
76.7
74.9
5.1
WoLA 78.5
78.6
76.4
68.5
77.9
77.2
73.3
8.5
78.9
78.9
78.4
3.5
78.5
77.6
76.6
73.1
77.6
75.4
66.2
9.2
78.3
77.5
74.2
32.5
77.6
75.4
66.2
3.5
WoLA+NNM 78.8
78.3
76.4
69.8
78.3
77.4
74.9
57.0
78.9
78.6
77.9
9.8
78.6
77.6
76.5
72.6
77.9
75.7
73.1
9.2
78.5
77.5
75.8
43.7
77.9
75.7
73.1
9.2
GM - 78.6
77.2
74.3
64.5
73.9
64.9
59.0
34.1
79.0
78.9
78.1
9.4
77.8
74.7
72.0
72.7
76.0
71.4
56.1
45.3
77.1
73.4
67.9
45.2
73.9
64.9
56.1
9.4
BKT 78.0
77.5
74.3
64.5
77.5
75.9
59.0
34.1
77.2
77.2
78.1
9.4
77.0
76.0
72.0
72.7
76.5
73.2
56.1
45.3
77.3
75.9
67.9
45.2
76.5
73.2
56.1
9.4
FoundFL 75.5
76.8
74.3
71.4
72.1
49.1
13.8
22.9
72.3
71.7
63.8
65.1
74.9
74.6
75.2
75.0
62.8
46.9
11.1
13.0
71.5
63.8
47.6
49.5
62.8
46.9
11.1
13.0
NNM 79.0
77.9
74.2
76.0
78.4
77.1
73.4
54.0
79.4
79.3
78.9
79.2
77.4
75.1
72.7
72.1
77.7
73.5
63.9
35.2
78.4
76.6
72.6
63.3
77.4
73.5
63.9
35.2
WoLA 78.9
78.9
77.7
74.4
78.1
76.3
73.8
66.5
79.0
79.1
78.4
78.3
79.0
77.7
78.0
78.1
77.6
74.8
70.8
60.5
78.5
77.4
75.7
71.5
77.6
74.8
70.8
60.5
WoLA+NNM 78.9
78.9
78.0
78.7
78.5
77.8
76.1
71.3
79.1
78.9
78.7
78.8
78.9
77.9
78.0
77.8
77.9
76.4
74.5
70.8
78.7
78.0
77.1
75.5
77.9
76.4
74.5
70.8
WoLA 79.3
78.5
76.7
69.4
78.0
76.0
71.8
60.1
79.1
78.6
77.5
73.4
78.7
77.6
76.5
74.2
77.7
74.5
69.7
54.5
78.6
77.0
74.4
66.3
77.7
74.5
69.7
54.5
WoLA+NNM 78.9
78.5
77.0
74.3
78.3
77.6
74.9
65.9
79.1
78.5
78.0
75.4
78.4
77.9
76.6
73.5
78.0
76.1
73.3
65.4
78.6
77.7
75.9
70.9
78.0
76.1
73.3
65.4
Table 9: Accuracy (s.d.) % on Fashion MNIST with $\alpha=1$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 77.5
73.8
63.0
42.8
72.4
67.5
55.5
29.1
73.0
60.2
41.3
15.0
73.1
70.9
64.8
63.8
70.2
63.4
53.9
25.8
73.2
67.2
55.7
35.3
70.2
60.2
41.3
15.0
BKT 78.0
76.1
63.0
42.8
77.2
74.3
55.5
29.1
76.4
72.0
41.3
15.0
77.0
73.9
64.8
63.8
76.2
70.7
53.9
25.8
77.0
73.4
55.7
35.3
76.2
70.7
41.3
15.0
FoundFL 64.8
69.8
66.8
66.6
56.3
44.8
12.2
26.5
62.7
54.3
49.8
27.4
41.3
63.9
69.1
73.0
44.6
21.2
12.0
11.2
53.9
50.8
42.0
40.9
41.3
21.2
12.0
11.2
NNM 80.2
78.0
72.4
75.4
79.4
77.4
73.0
21.7
80.1
80.2
79.6
75.0
78.1
73.3
68.8
62.9
78.5
72.8
57.4
31.6
79.3
76.4
70.2
53.3
78.1
72.8
57.4
21.7
WoLA 80.0
79.0
76.9
71.6
79.0
76.8
74.2
70.8
79.9
78.5
77.5
72.3
79.7
78.8
78.4
78.0
78.8
76.7
74.6
69.8
79.5
78.0
76.3
72.5
78.8
76.7
74.2
69.8
WoLA+NNM 80.2
79.5
78.2
78.6
79.5
78.3
76.2
60.1
79.9
79.6
79.9
80.2
79.5
78.5
78.2
77.9
79.0
76.8
75.4
72.5
79.6
78.6
77.6
73.8
79.0
76.8
75.4
60.1
WoLA 79.8
77.3
73.5
60.7
78.8
75.8
71.2
65.4
79.3
77.7
74.0
62.9
79.5
77.5
76.1
69.7
78.7
76.0
71.8
62.5
79.2
76.9
73.3
64.2
78.7
75.8
71.2
60.7
WoLA+NNM 79.9
77.9
76.0
68.7
79.5
77.1
73.4
49.1
80.2
78.6
78.0
75.0
79.3
77.5
75.1
69.0
79.1
75.4
71.1
46.1
79.6
77.3
74.7
61.5
79.1
75.4
71.1
46.1
CwTM - 79.4
76.4
68.0
43.4
77.3
72.6
61.6
27.0
76.6
69.0
48.0
14.1
77.3
73.1
68.5
61.6
76.6
64.5
57.3
21.5
77.5
71.1
60.7
33.5
76.6
64.5
48.0
14.1
BKT 78.1
76.1
68.0
43.4
77.2
74.3
61.6
27.0
76.4
72.4
48.0
14.1
77.5
74.4
68.5
61.6
76.2
70.6
57.3
21.5
77.1
73.6
60.7
33.5
76.2
70.6
48.0
14.1
FoundFL 70.1
68.7
62.0
58.4
68.9
63.4
59.0
36.7
69.2
51.0
45.2
26.7
70.3
65.8
68.1
71.6
65.3
16.7
26.4
15.1
68.7
53.2
52.1
41.7
65.3
16.7
26.4
15.1
NNM 80.5
78.4
72.7
75.2
80.0
77.9
73.2
22.5
80.5
80.3
79.6
71.9
77.5
74.2
70.9
62.6
79.2
74.1
60.6
33.5
79.5
77.0
71.4
53.2
77.5
74.1
60.6
22.5
WoLA 80.0
79.6
77.7
71.6
79.7
78.0
74.7
70.9
79.5
78.3
76.7
72.3
79.8
78.8
78.7
78.3
79.1
76.6
74.3
69.2
79.6
78.3
76.4
72.5
79.1
76.6
74.3
69.2
WoLA+NNM 80.0
79.5
78.1
78.8
79.6
78.3
76.2
62.0
79.9
79.6
79.9
80.2
79.3
78.6
77.8
78.0
79.2
76.9
75.2
72.5
79.6
78.6
77.4
74.3
79.2
76.9
75.2
62.0
WoLA 79.9
78.1
74.1
59.9
79.4
77.0
71.5
65.6
79.2
77.4
70.7
64.4
79.1
77.5
76.1
69.1
79.1
75.3
71.0
62.1
79.3
77.1
72.7
64.2
79.1
75.3
70.7
59.9
WoLA+NNM 79.7
78.4
76.1
69.3
79.5
77.2
73.5
50.0
80.2
78.6
77.9
75.0
79.5
77.0
75.7
68.7
79.1
75.5
71.4
45.4
79.6
77.3
74.9
61.7
79.1
75.5
71.4
45.4
MKrum - 79.9
77.6
71.3
19.0
78.9
77.8
72.8
32.6
80.3
78.0
40.6
11.0
77.9
74.0
70.3
62.9
78.4
73.0
26.3
6.8
79.1
76.1
56.3
26.5
77.9
73.0
26.3
6.8
BKT 77.8
75.5
71.3
19.0
76.4
74.1
72.8
32.6
75.9
68.9
40.6
11.0
77.1
73.4
70.3
62.9
75.1
65.9
26.3
6.8
76.5
71.6
56.3
26.5
75.1
65.9
26.3
6.8
FoundFL 71.4
66.8
64.4
52.0
69.6
65.7
61.2
41.8
75.0
71.1
19.8
19.4
68.2
67.0
69.9
69.1
58.5
25.8
24.1
26.7
68.5
59.3
47.9
41.8
58.5
25.8
19.8
19.4
NNM 80.5
78.4
72.6
29.3
79.9
77.8
73.1
29.8
80.5
80.2
79.0
19.2
77.9
73.4
70.5
62.3
78.8
73.1
42.9
7.1
79.5
76.6
67.6
29.5
77.9
73.1
42.9
7.1
WoLA 80.1
79.2
77.8
72.1
79.3
77.9
75.2
24.1
79.9
79.6
79.9
8.1
79.5
78.4
78.2
77.9
79.0
75.9
67.4
5.4
79.6
78.2
75.7
37.5
79.0
75.9
67.4
5.4
WoLA+NNM 79.9
79.5
78.2
72.7
79.7
78.3
76.0
59.4
79.9
79.6
79.9
15.9
79.5
78.5
78.3
77.9
79.2
76.9
74.9
5.4
79.6
78.6
77.5
46.3
79.2
76.9
74.9
5.4
WoLA 79.8
77.6
75.6
61.0
79.2
76.9
71.9
20.5
80.1
78.6
77.8
5.8
79.2
77.2
75.7
69.0
78.9
74.8
59.7
8.4
79.4
77.0
72.1
32.9
78.9
74.8
59.7
5.8
WoLA+NNM 79.7
78.5
75.9
61.2
79.5
77.1
73.6
41.7
80.1
78.6
78.0
11.1
79.6
77.0
75.7
69.6
79.1
75.8
70.9
8.4
79.6
77.4
74.8
38.4
79.1
75.8
70.9
8.4
GM - 80.1
77.1
72.0
31.5
70.2
60.7
47.5
21.8
80.2
79.7
72.5
4.6
78.4
74.3
67.2
67.5
75.1
69.2
61.6
44.7
76.8
72.2
64.2
34.0
70.2
60.7
47.5
4.6
BKT 79.0
77.5
72.0
31.5
78.3
76.2
47.5
21.8
78.4
75.9
72.5
4.6
78.0
75.9
67.2
67.5
77.4
73.5
61.6
44.7
78.2
75.8
64.2
34.0
77.4
73.5
47.5
4.6
FoundFL 73.6
72.7
70.3
63.3
65.6
43.2
24.4
20.2
70.1
67.8
52.7
39.3
49.5
66.9
71.6
73.5
57.2
30.6
24.7
14.1
63.2
56.2
48.7
42.1
49.5
30.6
24.4
14.1
NNM 80.2
78.4
72.5
75.7
79.7
77.8
74.1
51.6
80.5
80.5
79.7
77.9
78.0
74.9
68.4
67.7
78.8
73.6
57.2
23.7
79.4
77.0
70.4
59.3
78.0
73.6
57.2
23.7
WoLA 79.9
79.4
77.9
73.0
78.6
76.4
70.5
61.1
80.1
79.6
79.7
78.9
79.6
78.7
77.7
78.6
78.5
75.8
71.5
41.5
79.4
77.9
75.5
66.6
78.5
75.8
70.5
41.5
WoLA+NNM 79.9
79.7
78.3
78.3
79.7
78.4
76.3
70.7
80.2
79.7
79.6
80.0
79.6
78.7
78.2
78.8
79.1
76.9
74.6
70.4
79.7
78.7
77.4
75.6
79.1
76.9
74.6
70.4
WoLA 79.8
78.0
75.5
62.5
78.6
74.7
67.3
51.6
79.9
78.7
77.8
56.7
79.6
77.1
75.5
70.3
78.3
74.6
68.8
52.6
79.3
76.6
73.0
58.7
78.3
74.6
67.3
51.6
WoLA+NNM 79.8
78.0
76.4
68.9
79.5
77.0
73.5
52.3
79.9
78.6
77.7
74.3
79.4
77.2
76.5
69.5
79.2
76.1
70.9
41.9
79.6
77.4
75.0
61.3
79.2
76.1
70.9
41.9
Table 10: Accuracy (s.d.) % on Fashion MNIST with $\alpha=0.3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 72.6
65.3
39.6
19.5
65.1
54.5
39.4
16.0
58.2
41.6
23.6
6.0
65.3
47.2
50.0
33.1
50.3
38.2
29.0
16.3
62.3
49.4
36.3
18.2
50.3
38.2
23.6
6.0
BKT 76.1
69.8
39.6
19.5
75.0
68.8
39.4
16.0
74.6
63.7
23.6
6.0
76.2
68.3
50.0
33.1
73.4
55.2
29.0
16.3
75.1
65.1
36.3
18.2
73.4
55.2
23.6
6.0
FoundFL 30.6
47.7
41.2
39.6
43.8
23.7
11.6
13.0
34.5
42.6
27.2
15.5
42.9
40.5
43.9
37.2
24.1
22.5
16.5
9.7
35.2
35.4
28.1
23.0
24.1
22.5
11.6
9.7
NNM 80.4
76.0
52.2
68.5
78.9
76.0
65.6
27.1
79.9
77.4
58.4
15.8
77.3
69.4
58.9
35.1
78.2
69.1
29.4
16.6
79.0
73.6
52.9
32.7
77.3
69.1
29.4
15.8
WoLA 78.7
76.7
72.6
49.3
77.2
75.1
65.8
40.1
77.9
76.3
68.7
32.1
78.2
76.2
75.9
71.9
76.7
74.5
64.4
32.7
77.7
75.8
69.5
45.2
76.7
74.5
64.4
32.1
WoLA+NNM 78.9
78.2
76.2
69.4
78.4
77.5
75.0
19.6
79.0
79.7
80.5
79.3
78.5
76.3
77.0
71.4
77.6
75.0
60.2
25.6
78.5
77.4
73.8
53.1
77.6
75.0
60.2
19.6
WoLA 78.6
74.4
68.1
40.9
76.8
71.7
61.1
47.0
77.8
73.3
65.6
22.5
77.9
73.7
71.0
48.4
76.1
71.7
58.7
39.9
77.5
72.9
64.9
39.8
76.1
71.7
58.7
22.5
WoLA+NNM 78.6
76.3
71.1
53.7
78.3
75.0
69.0
17.9
78.7
77.7
76.8
70.7
77.7
73.8
70.0
63.3
77.8
71.9
55.7
12.2
78.2
74.9
68.5
43.6
77.7
71.9
55.7
12.2
CwTM - 77.0
70.0
43.3
19.1
74.6
65.2
41.1
16.0
74.2
54.8
23.0
5.9
73.1
62.7
54.9
26.0
72.1
38.1
24.3
17.2
74.2
58.2
37.3
16.8
72.1
38.1
23.0
5.9
BKT 76.1
69.6
43.3
19.1
75.0
68.8
41.1
16.0
74.6
63.7
23.0
5.9
75.1
68.6
54.9
26.0
73.4
55.3
24.3
17.2
74.9
65.2
37.3
16.8
73.4
55.3
23.0
5.9
FoundFL 59.3
52.0
65.4
49.3
65.8
41.5
31.6
23.4
57.3
44.7
35.9
14.7
65.5
29.6
55.0
62.3
28.1
14.1
18.8
17.4
55.2
36.4
41.3
33.4
28.1
14.1
18.8
14.7
NNM 80.9
76.5
51.9
69.6
80.1
77.3
66.6
26.7
80.6
76.0
61.2
15.6
77.6
71.1
56.5
35.4
79.7
72.8
40.9
16.3
79.8
74.7
55.4
32.7
77.6
71.1
40.9
15.6
WoLA 79.1
78.5
74.7
49.8
78.4
77.1
72.5
40.3
78.2
77.1
71.9
32.6
78.5
77.0
76.1
72.7
77.8
74.8
69.9
32.9
78.4
76.9
73.0
45.7
77.8
74.8
69.9
32.6
WoLA+NNM 79.1
78.4
76.1
66.8
78.2
77.4
75.5
21.0
79.0
79.7
80.4
79.2
78.4
76.7
76.7
70.6
77.9
74.9
65.3
23.6
78.5
77.4
74.8
52.2
77.9
74.9
65.3
21.0
WoLA 78.8
76.0
70.4
35.7
78.0
75.0
67.3
47.0
77.9
74.4
69.5
19.1
77.9
74.2
71.5
62.1
77.7
71.3
62.3
40.1
78.1
74.2
68.2
40.8
77.7
71.3
62.3
19.1
WoLA+NNM 78.8
76.2
72.0
50.8
78.2
75.2
69.8
19.3
78.8
77.7
76.8
70.7
77.9
73.9
71.3
48.8
78.0
71.6
60.7
15.5
78.4
74.9
70.1
41.0
77.9
71.6
60.7
15.5
MKrum - 79.4
75.3
47.5
5.1
79.5
75.9
62.6
22.0
80.1
60.8
45.2
18.0
77.1
64.5
47.7
35.1
78.4
66.0
13.8
6.1
78.9
68.5
43.4
17.3
77.1
60.8
13.8
5.1
BKT 76.9
72.2
47.5
5.1
75.5
71.5
62.6
22.0
74.1
60.4
45.2
18.0
75.6
68.0
47.7
35.1
74.0
56.2
13.8
6.1
75.2
65.7
43.4
17.3
74.0
56.2
13.8
5.1
FoundFL 70.5
57.5
53.6
28.5
68.9
59.4
48.8
26.4
36.0
45.1
29.4
17.8
70.1
61.9
52.0
54.4
54.3
41.3
16.4
17.4
60.0
53.0
40.0
28.9
36.0
41.3
16.4
17.4
NNM 80.6
75.9
55.6
6.2
80.2
77.1
66.6
23.7
80.4
75.7
50.1
16.7
76.9
66.6
63.2
42.1
79.5
69.7
33.6
6.1
79.5
73.0
53.8
19.0
76.9
66.6
33.6
6.1
WoLA 78.8
78.0
75.9
51.5
77.8
76.8
74.7
28.6
79.0
79.7
80.4
4.8
77.9
76.3
74.3
71.4
77.3
73.8
38.6
5.2
78.2
76.9
68.8
32.3
77.3
73.8
38.6
4.8
WoLA+NNM 79.1
78.3
75.8
52.4
78.3
77.5
75.0
26.9
79.0
79.7
80.5
10.3
78.4
76.3
76.0
71.9
78.0
74.5
64.1
5.1
78.5
77.3
74.3
33.3
78.0
74.5
64.1
5.1
WoLA 78.4
75.8
71.7
40.0
77.8
74.6
69.5
23.9
78.8
77.8
76.7
3.0
77.6
73.7
69.8
63.3
77.4
70.7
15.8
6.4
78.0
74.5
60.7
27.3
77.4
70.7
15.8
3.0
WoLA+NNM 78.8
76.2
71.0
40.7
78.2
74.9
69.2
15.9
78.8
77.7
76.8
7.6
77.9
74.1
70.0
61.7
78.0
71.8
58.5
6.4
78.3
74.9
69.1
26.5
77.9
71.8
58.5
6.4
GM - 79.4
75.2
58.5
6.5
67.2
43.6
29.5
12.6
80.4
76.3
50.7
1.4
77.2
68.2
48.1
44.8
71.1
62.3
52.7
8.1
75.1
65.1
47.9
14.7
67.2
43.6
29.5
1.4
BKT 78.6
76.3
58.5
6.5
78.0
75.9
29.5
12.6
77.0
70.1
50.7
1.4
77.1
74.5
48.1
44.8
77.3
71.3
52.7
8.1
77.6
73.6
47.9
14.7
77.0
70.1
29.5
1.4
FoundFL 42.5
62.5
61.6
52.3
42.0
35.9
26.7
41.2
37.3
48.5
48.2
29.1
42.7
37.1
62.8
56.4
40.1
25.0
16.4
17.4
40.9
41.8
43.1
39.3
37.3
25.0
16.4
17.4
NNM 80.5
76.6
52.9
68.0
79.8
77.5
65.9
29.8
80.6
78.5
51.1
19.1
77.1
70.2
60.4
51.1
79.3
71.0
42.4
19.3
79.5
74.8
54.5
37.5
77.1
70.2
42.4
19.1
WoLA 78.9
78.1
76.0
54.6
76.0
70.7
62.6
36.6
79.1
79.4
78.7
35.2
78.5
76.8
75.9
74.5
76.1
72.5
65.6
29.3
77.7
75.5
71.8
46.0
76.0
70.7
62.6
29.3
WoLA+NNM 79.0
78.5
76.5
69.7
78.2
77.5
75.4
29.8
79.0
79.5
79.7
78.5
78.4
76.9
75.8
73.1
78.1
75.2
56.9
31.2
78.6
77.5
72.8
56.4
78.1
75.2
56.9
29.8
WoLA 78.7
76.0
71.2
47.6
75.7
66.8
54.0
28.1
78.8
77.5
74.8
27.7
77.8
75.0
70.4
58.5
75.8
70.0
58.2
27.4
77.3
73.1
65.7
37.8
75.7
66.8
54.0
27.4
WoLA+NNM 78.8
76.4
72.1
54.6
78.2
75.4
70.1
24.8
78.9
77.7
76.6
71.1
77.7
74.3
71.7
61.0
78.1
71.8
56.9
11.2
78.3
75.1
69.5
44.5
77.7
71.8
56.9
11.2
Table 11: Accuracy (s.d.) % on Purchase100 with $\alpha=3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 72.2
69.3
64.2
40.2
67.2
60.7
50.2
21.7
70.3
68.0
61.2
29.9
68.5
61.6
50.4
34.9
66.8
60.6
49.8
16.8
69.0
64.0
55.1
28.7
66.8
60.6
49.8
16.8
BKT 72.4
71.7
64.2
40.2
70.5
67.0
50.2
21.7
69.4
66.4
61.2
29.9
70.5
65.0
50.4
34.9
69.1
62.9
49.8
16.8
70.4
66.6
55.1
28.7
69.1
62.9
49.8
16.8
FoundFL 64.2
61.3
55.8
49.0
59.7
30.1
2.7
8.8
52.0
24.9
0.2
0.3
61.2
60.4
58.3
62.7
43.7
13.8
0.9
0.6
56.2
38.1
23.6
24.3
43.7
13.8
0.2
0.3
NNM 72.9
72.3
67.4
72.7
71.1
67.2
58.3
3.8
73.0
72.9
72.7
70.9
70.7
63.3
55.3
34.9
70.4
63.2
52.7
44.9
71.6
67.7
61.3
45.5
70.4
63.2
52.7
3.8
WoLA 71.5
71.2
69.9
66.5
69.9
67.8
64.9
58.5
71.2
70.7
69.4
64.8
71.0
69.4
67.9
63.7
69.8
67.3
64.2
55.8
70.7
69.3
67.3
61.8
69.8
67.3
64.2
55.8
WoLA+NNM 71.2
71.1
70.1
70.9
70.6
69.3
67.3
62.4
71.3
71.0
71.0
71.0
70.9
69.6
67.4
63.7
70.4
68.4
67.3
66.7
70.9
69.9
68.6
66.9
70.4
68.4
67.3
62.4
WoLA 71.3
70.7
68.5
61.3
69.8
67.3
63.3
53.8
71.1
70.3
68.3
55.5
70.7
69.3
65.9
61.8
69.6
66.9
62.8
53.7
70.5
68.9
65.8
57.2
69.6
66.9
62.8
53.7
WoLA+NNM 71.1
70.7
69.1
68.2
70.5
68.8
65.7
47.0
71.2
70.7
69.9
68.8
70.7
68.8
66.1
61.8
70.2
67.8
65.9
60.5
70.7
69.4
67.3
61.3
70.2
67.8
65.7
47.0
CwTM - 72.4
72.0
65.9
40.2
70.4
65.7
54.3
21.7
71.1
68.4
61.8
29.9
70.1
63.8
52.0
34.9
69.1
61.3
49.6
16.5
70.7
66.2
56.7
28.6
69.1
61.3
49.6
16.5
BKT 72.4
71.7
65.9
40.2
70.5
67.0
54.3
21.7
69.4
66.4
61.8
29.9
70.5
65.0
52.0
34.9
69.1
62.9
49.6
16.5
70.4
66.6
56.7
28.6
69.1
62.9
49.6
16.5
FoundFL 62.3
57.3
53.0
49.5
61.2
55.0
54.5
7.8
45.1
27.4
10.5
0.5
58.9
55.1
57.0
61.1
37.3
1.7
0.9
0.6
53.0
39.3
35.2
23.9
37.3
1.7
0.9
0.5
NNM 72.9
72.6
68.7
72.7
71.1
67.3
58.3
3.8
73.0
72.9
72.7
70.9
70.9
63.3
55.6
34.9
70.4
63.8
53.3
44.9
71.7
68.0
61.7
45.5
70.4
63.3
53.3
3.8
WoLA 71.4
71.4
70.4
66.5
70.2
68.1
65.0
58.5
71.1
70.4
69.1
64.8
71.0
69.5
67.9
63.7
69.8
67.1
63.5
55.8
70.7
69.3
67.2
61.9
69.8
67.1
63.5
55.8
WoLA+NNM 71.3
71.2
70.3
70.9
70.7
69.4
67.3
62.4
71.3
71.0
71.0
71.0
70.8
69.6
67.4
63.7
70.4
68.4
67.2
66.7
70.9
69.9
68.7
66.9
70.4
68.4
67.2
62.4
WoLA 71.2
70.9
69.2
61.3
70.0
67.7
63.4
53.8
71.0
70.0
68.0
55.2
70.8
69.2
66.4
61.8
69.6
66.6
62.1
53.7
70.5
68.9
65.8
57.2
69.6
66.6
62.1
53.7
WoLA+NNM 71.1
70.7
69.1
68.2
70.5
68.9
65.8
47.0
71.2
70.7
69.9
68.8
70.7
69.0
66.2
61.8
70.2
67.9
65.9
60.5
70.7
69.4
67.4
61.3
70.2
67.9
65.8
47.0
MKrum - 72.8
72.2
67.1
38.5
71.0
66.9
57.9
1.9
72.8
70.6
24.4
0.8
70.8
64.5
52.7
34.9
69.8
57.7
1.6
0.7
71.4
66.4
40.7
15.4
69.8
57.7
1.6
0.7
BKT 71.8
70.1
67.1
38.5
69.3
63.1
57.9
1.9
66.8
58.3
24.4
0.8
69.0
54.9
52.7
34.9
67.2
44.7
1.6
0.7
68.8
58.2
40.7
15.4
66.8
44.7
1.6
0.7
FoundFL 61.0
58.2
55.8
41.5
57.4
54.6
46.0
1.4
42.1
20.9
5.5
0.5
58.7
55.7
49.8
49.5
39.9
1.0
0.7
0.7
51.8
38.1
31.6
18.7
39.9
1.0
0.7
0.5
NNM 72.9
72.4
67.9
44.9
71.1
67.2
58.2
4.0
73.0
72.9
72.7
3.0
70.8
64.0
54.6
34.9
70.4
63.2
50.6
1.0
71.6
67.9
60.8
17.6
70.4
63.2
50.6
1.0
WoLA 71.1
71.0
70.1
66.8
70.5
68.8
61.6
3.4
71.3
71.0
71.0
0.8
70.8
69.2
67.0
63.7
70.2
66.8
58.3
0.7
70.8
69.4
65.6
27.1
70.2
66.8
58.3
0.7
WoLA+NNM 71.2
71.1
70.2
67.4
70.6
69.3
67.2
32.2
71.3
71.0
71.0
3.2
70.8
69.4
67.3
63.7
70.4
68.3
66.9
0.7
70.9
69.8
68.5
33.4
70.4
68.3
66.9
0.7
WoLA 70.9
70.5
68.6
61.8
70.3
68.3
59.1
1.9
71.2
70.7
69.8
0.9
70.6
68.7
65.3
61.8
70.0
66.3
56.3
0.5
70.6
68.9
63.8
25.4
70.0
66.3
56.3
0.5
WoLA+NNM 71.1
70.7
68.9
63.2
70.5
68.9
65.7
22.3
71.2
70.7
69.9
2.9
70.7
68.8
65.7
61.8
70.2
67.7
65.4
0.6
70.7
69.4
67.1
30.2
70.2
67.7
65.4
0.6
GM - 73.2
71.7
67.0
50.6
69.3
63.3
53.9
39.0
72.2
69.8
64.4
0.8
70.8
61.1
45.6
41.7
66.9
57.1
15.1
1.1
70.5
64.6
49.2
26.6
66.9
57.1
15.1
0.8
BKT 72.9
72.6
67.0
50.6
71.1
67.9
53.9
39.0
69.7
66.3
64.4
0.8
71.3
66.5
45.6
41.7
69.4
62.1
15.1
1.1
70.9
67.1
49.2
26.6
69.4
62.1
15.1
0.8
FoundFL 58.4
55.6
55.6
52.5
54.6
37.2
3.6
2.0
55.9
0.4
0.3
0.4
57.0
53.9
56.2
58.5
13.5
4.6
1.0
0.7
47.9
30.4
23.3
22.8
13.5
0.4
0.3
0.4
NNM 72.9
72.7
68.2
69.8
71.1
67.4
59.3
22.6
72.9
72.4
71.3
68.6
70.8
63.7
55.7
40.7
70.3
64.1
52.3
17.0
71.6
68.1
61.3
43.7
70.3
63.7
52.3
17.0
WoLA 71.3
71.3
70.9
68.3
69.8
67.6
65.0
58.4
71.0
70.2
69.0
65.7
70.9
70.0
69.2
68.4
69.0
65.5
59.7
38.5
70.4
68.9
66.8
59.9
69.0
65.5
59.7
38.5
WoLA+NNM 71.3
71.2
70.5
70.6
70.5
68.7
65.5
59.3
71.2
70.5
69.5
68.2
70.9
69.4
68.1
65.4
70.1
67.1
63.7
51.4
70.8
69.4
67.4
63.0
70.1
67.1
63.7
51.4
WoLA 71.2
70.9
69.6
65.3
69.6
67.1
63.6
53.2
70.9
69.9
68.1
61.7
70.8
69.3
68.0
66.3
68.8
64.9
58.2
19.7
70.3
68.4
65.5
53.2
68.8
64.9
58.2
19.7
WoLA+NNM 71.1
70.8
69.3
66.6
70.4
68.2
64.1
55.4
71.1
70.2
68.6
66.6
70.7
68.9
66.5
63.2
70.0
66.6
62.3
50.8
70.7
68.9
66.2
60.5
70.0
66.6
62.3
50.8
Table 12: Accuracy (s.d.) % on Purchase100 with $\alpha=1$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 69.9
63.2
50.0
5.9
61.6
49.0
29.3
12.4
66.6
61.3
41.4
11.1
61.6
50.3
29.6
12.4
61.1
49.3
27.9
1.5
64.2
54.6
35.6
8.7
61.1
49.0
27.9
1.5
BKT 71.5
68.9
50.0
5.9
69.2
63.9
29.3
12.4
68.0
62.7
41.4
11.1
65.8
53.7
29.6
12.4
67.1
57.3
27.9
1.5
68.3
61.3
35.6
8.7
65.8
53.7
27.9
1.5
FoundFL 58.5
50.6
17.5
4.7
50.8
22.5
1.9
2.2
24.3
5.9
5.2
10.0
54.6
55.7
56.2
56.8
23.7
10.3
1.5
1.5
42.4
29.0
16.5
15.0
23.7
5.9
1.5
1.5
NNM 72.9
71.2
59.6
73.1
71.1
66.6
57.0
1.2
72.6
71.6
71.7
69.4
67.3
53.5
29.5
12.6
69.0
55.7
41.0
24.1
70.6
63.7
51.7
36.1
67.3
53.5
29.5
1.2
WoLA 70.7
69.8
67.2
60.2
68.6
65.5
60.0
39.8
70.5
69.4
66.5
53.3
69.9
67.7
63.1
57.9
68.4
64.9
59.3
37.7
69.6
67.5
63.2
49.8
68.4
64.9
59.3
37.7
WoLA+NNM 70.0
69.7
68.0
69.8
68.8
67.0
63.6
51.1
70.1
69.9
69.7
69.8
69.0
66.7
62.4
57.9
68.7
66.0
63.2
62.3
69.3
67.9
65.4
62.1
68.7
66.0
62.4
51.1
WoLA 70.5
69.3
65.2
36.6
68.4
64.6
58.1
36.4
70.2
68.9
64.6
37.7
69.7
66.5
61.1
42.0
68.1
64.1
57.8
36.9
69.4
66.7
61.4
37.9
68.1
64.1
57.8
36.4
WoLA+NNM 69.7
69.2
67.1
59.0
68.7
66.3
58.3
24.6
69.9
69.4
68.8
67.2
69.0
66.1
60.4
41.8
68.5
65.3
60.9
34.3
69.1
67.2
63.1
45.4
68.5
65.3
58.3
24.6
CwTM - 71.6
67.5
52.1
6.0
69.0
61.4
37.7
12.4
69.7
64.0
47.2
10.5
64.4
50.2
31.5
11.7
66.9
53.1
31.9
1.5
68.3
59.2
40.1
8.4
64.4
50.2
31.5
1.5
BKT 71.5
69.1
52.1
6.0
69.2
63.9
37.7
12.4
68.0
62.7
47.2
10.5
65.8
53.7
31.5
11.7
67.1
57.3
31.9
1.5
68.3
61.3
40.1
8.4
65.8
53.7
31.5
1.5
FoundFL 45.0
47.5
17.0
3.3
48.9
46.9
52.8
4.1
35.9
23.2
9.1
11.4
44.9
40.2
40.3
52.6
24.1
1.6
2.0
1.9
39.7
31.9
24.2
14.6
24.1
1.6
2.0
1.9
NNM 73.2
71.9
59.9
73.1
71.2
66.6
56.5
1.2
72.6
71.7
71.6
69.4
67.0
52.4
34.5
14.4
69.1
56.2
40.7
24.1
70.6
63.7
52.6
36.4
67.0
52.4
34.5
1.2
WoLA 70.4
70.4
67.8
59.7
68.9
66.1
60.9
39.8
69.7
68.7
66.2
53.4
69.7
67.4
63.9
57.9
68.2
64.3
58.3
37.7
69.4
67.4
63.4
49.7
68.2
64.3
58.3
37.7
WoLA+NNM 70.0
69.8
68.4
69.8
68.9
67.0
63.6
51.1
70.1
69.9
69.7
69.8
69.1
67.0
62.7
57.9
68.7
66.1
63.6
62.3
69.4
68.0
65.6
62.1
68.7
66.1
62.7
51.1
WoLA 70.2
69.9
66.7
37.0
68.7
65.4
58.8
36.4
69.4
68.1
64.7
37.5
69.6
66.4
61.4
41.7
67.9
63.4
56.5
36.9
69.2
66.6
61.6
37.9
67.9
63.4
56.5
36.4
WoLA+NNM 69.8
69.1
67.1
59.1
68.7
66.3
58.9
24.6
69.9
69.4
68.8
67.2
69.1
66.2
61.0
41.9
68.5
65.3
60.9
34.3
69.2
67.2
63.3
45.4
68.5
65.3
58.9
24.6
MKrum - 72.9
69.8
58.4
6.3
71.0
66.7
56.3
1.3
72.4
63.4
29.0
3.1
67.3
55.2
36.0
12.6
69.0
55.4
1.3
0.3
70.5
62.1
36.2
4.7
67.3
55.2
1.3
0.3
BKT 70.6
67.1
58.4
6.3
67.7
60.9
56.3
1.3
65.1
51.3
29.0
3.1
63.4
39.0
36.0
12.6
64.9
35.0
1.3
0.3
66.3
50.7
36.2
4.7
63.4
35.0
1.3
0.3
FoundFL 49.7
41.5
40.5
11.1
44.3
37.2
43.0
1.9
25.0
26.0
12.3
2.1
43.9
37.9
20.5
30.9
42.8
2.0
1.7
1.9
41.2
28.9
23.6
9.6
25.0
2.0
1.7
1.9
NNM 73.1
71.3
58.5
13.1
71.1
66.7
56.2
1.2
72.6
71.6
71.7
7.1
67.5
54.1
33.9
14.8
69.1
55.5
33.7
0.3
70.7
63.8
50.8
7.3
67.5
54.1
33.7
0.3
WoLA 69.7
69.0
67.6
60.4
68.5
64.5
55.3
3.6
70.1
69.9
69.7
2.1
69.0
67.0
61.7
57.9
68.4
64.1
25.8
0.3
69.2
66.9
56.0
24.8
68.4
64.1
25.8
0.3
WoLA+NNM 70.0
69.8
68.2
62.3
68.9
67.0
63.4
26.3
70.1
69.9
69.7
6.5
69.1
66.7
62.1
57.9
68.7
66.0
63.2
0.3
69.4
67.9
65.3
30.6
68.7
66.0
62.1
0.3
WoLA 69.5
68.5
66.6
40.7
68.3
63.6
53.4
1.4
69.9
69.4
68.8
1.7
68.8
65.9
60.1
41.8
68.1
63.2
19.5
0.5
68.9
66.1
53.7
17.2
68.1
63.2
19.5
0.5
WoLA+NNM 69.8
69.2
67.2
44.4
68.7
66.2
58.0
12.0
69.9
69.4
68.8
5.4
69.0
65.8
59.3
41.6
68.5
65.2
60.8
0.5
69.2
67.1
62.8
20.8
68.5
65.2
58.0
0.5
GM - 73.4
68.6
56.2
15.7
66.6
56.1
40.2
18.8
71.7
68.3
53.9
0.4
67.4
48.2
19.8
14.9
63.4
39.8
13.3
1.9
68.5
56.2
36.7
10.3
63.4
39.8
13.3
0.4
BKT 72.9
71.2
56.2
15.7
70.9
67.1
40.2
18.8
68.9
63.2
53.9
0.4
68.3
58.8
19.8
14.9
68.9
59.6
13.3
1.9
70.0
64.0
36.7
10.3
68.3
58.8
13.3
0.4
FoundFL 44.9
50.3
45.7
19.2
43.9
22.9
2.0
1.8
13.2
1.1
1.0
1.1
42.3
42.3
42.6
46.5
16.4
12.3
1.7
1.9
32.1
25.8
18.6
14.1
13.2
1.1
1.0
1.1
NNM 73.2
71.1
59.9
60.3
71.2
67.0
57.8
14.1
72.5
71.6
71.0
67.5
66.8
55.0
36.9
10.2
69.2
57.6
38.8
14.7
70.6
64.5
52.9
33.4
66.8
55.0
36.9
10.2
WoLA 70.1
70.0
68.4
63.6
68.0
65.1
60.7
52.1
69.8
69.2
68.0
59.8
69.3
66.3
63.9
62.6
66.8
61.8
52.6
1.6
68.8
66.5
62.7
47.9
66.8
61.8
52.6
1.6
WoLA+NNM 70.0
69.9
68.6
66.5
68.8
66.7
62.6
53.9
70.0
69.5
68.9
68.0
69.2
67.1
63.3
59.3
68.5
64.9
60.9
48.0
69.3
67.6
64.9
59.2
68.5
64.9
60.9
48.0
WoLA 69.9
69.5
67.0
50.8
67.8
64.3
57.7
42.6
69.6
68.7
66.8
53.0
69.2
65.3
61.5
44.4
66.6
60.8
49.1
13.2
68.6
65.7
60.4
40.8
66.6
60.8
49.1
13.2
WoLA+NNM 69.8
69.4
68.0
56.4
68.6
65.9
59.3
42.9
69.8
69.1
67.9
65.2
69.0
66.8
62.0
44.9
68.3
64.1
58.6
30.1
69.1
67.1
63.1
47.9
68.3
64.1
58.6
30.1
Table 13: Accuracy (s.d.) % on Purchase100 with $\alpha=0.3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 61.5
42.1
8.4
2.5
47.8
30.9
14.8
3.4
54.0
38.4
11.5
2.1
48.4
30.3
5.1
2.1
44.1
21.8
1.8
1.3
51.2
32.7
8.3
2.3
44.1
21.8
1.8
1.3
BKT 69.2
62.2
8.4
2.5
65.4
55.0
14.8
3.4
64.0
52.3
11.5
2.1
57.9
35.2
5.1
2.1
62.2
40.9
1.8
1.3
63.7
49.1
8.3
2.3
57.9
35.2
1.8
1.3
FoundFL 17.2
6.0
2.9
4.9
22.3
16.2
2.9
3.1
20.0
8.8
5.4
2.4
23.8
17.0
27.0
24.2
3.1
4.9
2.9
2.4
17.3
10.6
8.2
7.4
3.1
4.9
2.9
2.4
NNM 73.3
66.1
30.2
74.4
71.3
65.8
53.1
1.1
72.7
71.4
48.4
14.3
60.4
39.5
4.7
2.1
68.8
50.6
3.1
1.8
69.3
58.7
27.9
18.7
60.4
39.5
3.1
1.1
WoLA 66.1
62.9
54.0
28.5
59.4
47.5
29.8
13.8
63.3
57.1
43.3
18.8
62.8
55.0
45.6
36.6
58.5
45.8
29.1
11.0
62.0
53.6
40.4
21.7
58.5
45.8
29.1
11.0
WoLA+NNM 65.6
65.5
60.2
66.1
63.2
59.3
48.8
2.4
65.7
65.9
65.7
66.1
63.7
57.7
49.5
36.6
61.6
53.4
39.2
21.4
64.0
60.4
52.7
38.5
61.6
53.4
39.2
2.4
WoLA 65.8
61.7
41.1
7.4
59.1
46.2
28.7
15.1
63.0
55.4
38.0
10.9
62.8
55.2
41.5
15.8
58.3
44.9
29.5
11.9
61.8
52.7
35.7
12.2
58.3
44.9
28.7
7.4
WoLA+NNM 65.2
64.3
55.3
27.6
62.9
57.9
47.1
2.3
65.3
64.7
63.4
61.7
63.7
57.5
44.2
15.7
61.3
51.4
21.9
12.8
63.7
59.2
46.4
24.0
61.3
51.4
21.9
2.3
CwTM - 69.4
59.0
15.5
2.4
65.1
50.2
20.4
3.4
65.6
50.4
15.9
2.1
57.8
38.8
4.7
2.0
61.4
31.9
2.2
1.1
63.9
46.1
11.8
2.2
57.8
31.9
2.2
1.1
BKT 69.2
62.4
15.5
2.4
65.4
55.0
20.4
3.4
64.0
52.3
15.9
2.1
57.9
35.2
4.7
2.0
62.2
40.9
2.2
1.1
63.7
49.1
11.8
2.2
57.9
35.2
2.2
1.1
FoundFL 29.9
10.1
3.9
4.3
31.9
26.2
31.1
3.6
19.4
17.8
17.9
1.7
32.8
17.9
15.7
15.0
4.8
1.6
2.9
3.0
23.8
14.7
14.3
5.5
4.8
1.6
2.9
1.7
NNM 73.5
68.8
36.9
74.4
71.2
65.8
53.1
1.1
72.8
71.6
48.2
14.3
62.3
32.9
4.7
2.2
68.7
52.0
3.6
1.7
69.7
58.2
29.3
18.7
62.3
32.9
3.6
1.1
WoLA 66.1
66.0
57.6
28.8
63.6
58.9
41.5
13.8
64.2
60.7
49.3
18.7
64.1
58.6
47.8
36.6
62.2
53.0
32.6
10.9
64.0
59.4
45.8
21.8
62.2
53.0
32.6
10.9
WoLA+NNM 65.6
65.6
61.1
66.1
63.2
59.4
49.3
2.4
65.7
65.9
65.7
66.1
63.6
59.0
46.4
36.6
61.7
53.5
40.4
21.4
64.0
60.7
52.6
38.5
61.7
53.5
40.4
2.4
WoLA 65.7
64.5
51.5
7.4
63.3
57.4
38.4
15.1
63.8
59.1
44.1
11.1
63.8
56.6
41.0
15.5
61.9
51.4
33.0
11.9
63.7
57.8
41.6
12.2
61.9
51.4
33.0
7.4
WoLA+NNM 65.3
64.4
55.7
27.4
62.9
58.0
47.5
2.3
65.3
64.7
63.4
61.7
63.5
57.2
41.7
15.7
61.4
51.6
13.7
12.8
63.7
59.2
44.4
24.0
61.4
51.6
13.7
2.3
MKrum - 73.3
68.0
26.7
2.4
71.2
66.1
53.1
1.0
71.2
53.6
35.5
10.5
60.9
33.8
6.3
2.1
69.0
50.1
1.6
0.4
69.1
54.3
24.6
3.3
60.9
33.8
1.6
0.4
BKT 68.3
60.8
26.7
2.4
65.4
58.2
53.1
1.0
61.5
43.5
35.5
10.5
54.7
10.2
6.3
2.1
61.5
31.1
1.6
0.4
62.3
40.8
24.6
3.3
54.7
10.2
1.6
0.4
FoundFL 33.7
39.7
5.0
5.4
36.4
42.7
18.5
3.0
22.4
20.8
10.6
5.8
30.8
14.7
6.4
10.7
23.5
3.2
2.9
3.0
29.4
24.2
8.7
5.6
22.4
3.2
2.9
3.0
NNM 73.5
67.6
36.4
2.6
71.2
66.0
53.0
1.1
72.7
71.7
47.2
8.9
62.3
38.1
5.7
2.1
68.7
51.9
3.1
0.4
69.7
59.1
29.1
3.0
62.3
38.1
3.1
0.4
WoLA 65.4
64.6
58.8
29.0
62.9
58.6
48.2
4.1
65.7
65.9
65.7
2.6
63.5
57.8
47.8
36.6
61.1
48.2
2.6
0.3
63.7
59.0
44.6
14.5
61.1
48.2
2.6
0.3
WoLA+NNM 65.6
65.4
60.4
36.7
63.2
59.2
48.4
2.7
65.7
65.9
65.7
5.9
63.8
58.7
46.6
36.6
61.7
53.5
40.4
0.4
64.0
60.6
52.3
16.5
61.7
53.5
40.4
0.4
WoLA 65.0
63.3
54.7
8.5
62.6
57.8
46.9
1.7
65.3
64.7
63.4
1.9
63.1
57.7
40.5
15.7
60.8
46.4
2.1
0.3
63.4
58.0
41.5
5.6
60.8
46.4
2.1
0.3
WoLA+NNM 65.3
64.3
55.2
9.8
62.9
58.0
47.1
2.1
65.3
64.7
63.4
4.8
63.6
57.4
44.3
15.5
61.4
51.4
12.6
0.2
63.7
59.1
44.5
6.5
61.4
51.4
12.6
0.2
GM - 72.1
61.5
30.7
3.2
61.3
41.2
17.2
2.7
71.8
65.4
28.6
0.6
62.0
26.4
2.7
5.1
56.7
29.2
14.8
1.1
64.8
44.7
18.8
2.5
56.7
26.4
2.7
0.6
BKT 72.8
69.1
30.7
3.2
70.7
66.0
17.2
2.7
67.6
60.2
28.6
0.6
62.3
39.3
2.7
5.1
68.5
55.6
14.8
1.1
68.4
58.0
18.8
2.5
62.3
39.3
2.7
0.6
FoundFL 27.4
17.7
5.3
8.4
36.4
9.8
3.0
3.0
22.7
2.9
16.1
3.1
30.2
29.1
22.5
17.0
23.0
4.2
2.9
3.0
28.0
12.7
10.0
6.9
22.7
2.9
2.9
3.0
NNM 73.6
67.8
37.1
45.2
71.4
66.6
55.0
1.6
72.7
71.7
48.7
47.6
62.1
37.0
3.8
2.8
68.9
53.3
2.9
1.5
69.7
59.3
29.5
19.8
62.1
37.0
2.9
1.5
WoLA 65.9
64.8
59.5
42.5
61.2
55.0
46.0
35.3
65.1
64.4
61.3
24.4
63.8
55.9
42.4
42.0
59.2
42.4
8.7
2.2
63.0
56.5
43.6
29.3
59.2
42.4
8.7
2.2
WoLA+NNM 65.7
65.9
61.5
57.3
63.3
59.5
50.0
19.7
65.5
65.5
64.8
65.1
63.7
59.6
49.6
42.6
61.7
53.6
38.6
6.7
64.0
60.8
52.9
38.3
61.7
53.6
38.6
6.7
WoLA 65.5
64.1
54.8
15.4
60.9
53.2
40.2
20.3
64.7
63.2
58.2
8.6
63.5
55.8
40.2
27.4
58.8
40.5
29.9
1.8
62.7
55.4
44.7
14.7
58.8
40.5
29.9
1.8
WoLA+NNM 65.3
64.6
57.7
24.2
63.0
58.1
47.6
13.5
65.2
64.3
62.4
60.4
63.6
58.7
47.8
17.6
61.4
52.2
21.4
2.7
63.7
59.6
47.4
23.7
61.4
52.2
21.4
2.7
Table 14: Accuracy (s.d.) % on CIFAR10 with $\alpha=3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 69.6
55.9
35.8
19.7
68.0
59.6
51.0
28.8
66.7
56.3
32.9
18.9
62.6
45.0
18.6
18.5
63.3
51.9
35.9
17.8
66.0
53.7
34.8
20.7
62.6
45.0
18.6
17.8
BKT 70.6
64.7
35.8
19.7
69.8
68.4
51.0
28.8
65.5
58.6
32.9
18.9
66.2
57.7
18.6
18.5
69.0
63.6
35.9
17.8
68.2
62.6
34.8
20.7
65.5
57.7
18.6
17.8
FoundFL 47.7
61.9
54.1
27.6
47.2
27.0
9.6
12.5
42.0
25.3
20.0
18.9
42.8
39.3
44.5
45.7
37.5
16.8
11.0
11.1
43.5
34.1
27.8
23.1
37.5
16.8
9.6
11.1
NNM 73.3
67.5
42.9
67.2
73.1
71.0
66.5
38.7
73.6
73.3
72.4
52.8
64.2
52.5
45.2
24.5
72.6
67.7
54.8
39.9
71.4
66.4
56.4
44.6
64.2
52.5
42.9
24.5
WoLA 71.7
63.4
47.2
28.4
70.6
65.2
59.9
51.7
71.0
68.1
60.7
42.4
69.4
64.1
58.7
52.3
68.6
62.2
56.5
42.4
70.2
64.6
56.6
43.5
68.6
62.2
47.2
28.4
WoLA+NNM 73.4
70.1
57.2
69.6
73.1
71.6
67.7
37.9
73.6
73.4
72.5
72.1
70.2
65.2
60.0
53.9
72.8
69.3
62.2
58.2
72.6
69.9
63.9
58.3
70.2
65.2
57.2
37.9
WoLA 71.3
62.3
44.1
19.9
70.5
64.5
58.5
48.0
70.9
67.3
58.0
33.1
69.5
63.1
56.4
45.6
68.4
61.7
54.3
37.4
70.1
63.8
54.3
36.8
68.4
61.7
44.1
19.9
WoLA+NNM 73.3
69.5
54.3
53.5
72.9
70.9
66.2
41.5
73.5
73.0
71.7
68.9
70.6
64.7
58.0
47.0
72.7
68.7
61.1
50.6
72.6
69.3
62.3
52.3
70.6
64.7
54.3
41.5
CwTM - 73.2
64.9
43.5
20.5
72.7
69.1
56.8
29.4
71.3
61.7
39.6
18.9
63.7
52.1
25.0
18.9
72.2
59.4
41.8
19.0
70.6
61.4
41.3
21.3
63.7
52.1
25.0
18.9
BKT 70.6
64.6
43.5
20.5
69.9
68.4
56.8
29.4
65.5
58.8
39.6
18.9
65.3
58.1
25.0
18.9
69.0
63.6
41.8
19.0
68.1
62.7
41.3
21.3
65.3
58.1
25.0
18.9
FoundFL 42.6
44.3
44.2
31.1
53.0
50.8
58.0
13.7
32.5
34.1
30.1
21.7
57.0
45.7
54.6
42.7
46.0
12.3
10.9
11.0
46.2
37.4
39.6
24.0
32.5
12.3
10.9
11.0
NNM 73.5
67.9
43.5
66.9
73.4
71.6
67.0
39.2
73.8
73.4
72.3
53.4
65.8
54.6
40.8
17.9
73.1
68.8
60.5
41.1
71.9
67.3
56.8
43.7
65.8
54.6
40.8
17.9
WoLA 73.2
68.7
51.6
27.9
72.8
70.7
62.1
51.5
72.3
68.3
60.5
42.7
70.5
64.3
59.6
51.9
72.4
65.2
56.6
41.2
72.3
67.4
58.1
43.0
70.5
64.3
51.6
27.9
WoLA+NNM 73.3
70.3
57.8
69.8
73.2
71.7
68.0
38.3
73.6
73.4
72.6
72.1
70.6
65.6
60.8
52.9
73.0
69.7
62.8
58.8
72.7
70.2
64.4
58.4
70.6
65.6
57.8
38.3
WoLA 73.3
68.0
48.7
19.8
72.8
70.1
60.0
47.8
72.1
67.7
58.0
33.1
70.3
63.3
58.2
44.8
72.4
64.4
55.2
36.5
72.2
66.7
56.0
36.4
70.3
63.3
48.7
19.8
WoLA+NNM 73.3
69.6
54.9
55.4
73.0
71.2
66.6
41.5
73.7
72.9
71.7
69.0
70.6
64.4
57.3
47.1
72.9
69.1
62.7
48.6
72.7
69.4
62.6
52.3
70.6
64.4
54.9
41.5
MKrum - 73.2
66.6
39.1
17.9
73.1
70.2
66.0
28.8
73.2
65.1
33.1
9.4
64.0
50.5
40.0
24.5
72.6
64.3
31.6
9.5
71.2
63.4
42.0
18.0
64.0
50.5
31.6
9.4
BKT 67.9
60.6
39.1
17.9
66.2
63.2
66.0
28.8
64.0
51.9
33.1
9.4
63.1
49.5
40.0
24.5
65.2
52.7
31.6
9.5
65.3
55.6
42.0
18.0
63.1
49.5
31.6
9.4
FoundFL 40.4
42.5
41.4
19.2
42.1
37.1
44.3
11.0
49.9
32.4
30.3
10.7
51.8
52.1
47.2
35.5
46.3
10.7
11.0
11.0
46.1
34.9
34.8
17.5
40.4
10.7
11.0
10.7
NNM 73.5
67.7
42.8
19.1
73.5
71.4
66.4
36.9
73.8
73.5
72.7
13.6
66.3
54.2
41.4
19.4
73.1
68.4
57.8
9.5
72.0
67.0
56.2
19.7
66.3
54.2
41.4
9.5
WoLA 73.0
69.9
56.5
29.0
72.5
70.8
66.7
34.0
73.6
72.9
67.7
10.8
70.3
66.2
60.4
53.9
72.3
67.1
41.7
9.1
72.4
69.4
58.6
27.4
70.3
66.2
41.7
9.1
WoLA+NNM 73.3
70.4
57.0
33.3
73.0
71.5
67.5
36.9
73.7
73.4
72.6
14.9
70.3
64.3
59.5
51.9
72.9
69.4
61.2
7.5
72.6
69.8
63.6
28.9
70.3
64.3
57.0
7.5
WoLA 72.9
69.2
53.5
21.0
72.6
70.2
65.4
30.2
73.5
72.5
66.1
11.5
70.4
64.2
58.3
47.0
72.2
66.4
30.9
10.5
72.3
68.5
54.8
24.0
70.4
64.2
30.9
10.5
WoLA+NNM 73.3
69.6
54.7
22.3
73.1
70.8
66.3
35.8
73.6
72.9
71.8
14.9
70.7
64.0
58.3
44.9
72.8
68.8
60.2
10.3
72.7
69.2
62.2
25.6
70.7
64.0
54.7
10.3
GM - 71.4
62.7
48.1
25.0
70.0
63.3
53.1
32.3
73.4
71.0
50.8
11.5
65.5
53.0
26.5
23.2
68.3
64.5
58.6
37.9
69.7
62.9
47.4
26.0
65.5
53.0
26.5
11.5
BKT 71.8
67.5
48.1
25.0
71.3
70.7
53.1
32.3
69.1
64.7
50.8
11.5
66.7
57.3
26.5
23.2
70.3
66.8
58.6
37.9
69.8
65.4
47.4
26.0
66.7
57.3
26.5
11.5
FoundFL 42.3
41.3
44.0
37.0
54.4
36.4
11.0
11.8
36.6
16.4
23.0
18.0
40.0
43.6
52.4
48.3
48.8
31.0
10.9
11.0
44.4
33.8
28.3
25.2
36.6
16.4
10.9
11.0
NNM 73.6
68.0
44.5
63.8
73.5
71.8
67.9
45.6
73.9
73.2
71.8
51.2
63.8
51.7
43.0
24.9
73.0
69.0
59.1
49.6
71.6
66.7
57.3
47.0
63.8
51.7
43.0
24.9
WoLA 72.5
67.3
55.3
35.7
71.6
68.6
64.9
59.4
73.3
72.6
70.2
49.7
70.8
63.5
58.5
57.3
69.2
64.6
61.2
28.5
71.5
67.3
62.0
46.1
69.2
63.5
55.3
28.5
WoLA+NNM 73.4
70.6
58.7
64.0
73.1
71.9
68.8
57.9
73.6
73.3
72.1
71.2
71.0
65.6
61.7
56.2
72.9
69.8
62.6
54.9
72.8
70.2
64.8
60.8
71.0
65.6
58.7
54.9
WoLA 72.4
66.6
52.6
25.3
71.7
67.9
63.0
48.9
73.2
72.1
68.9
34.4
70.7
62.5
56.8
52.7
69.0
63.7
59.4
40.3
71.4
66.5
60.1
40.3
69.0
62.5
52.6
25.3
WoLA+NNM 73.4
69.9
56.1
42.8
73.2
71.3
67.7
49.7
73.6
72.9
71.2
67.5
70.4
64.0
60.4
50.3
72.9
69.2
60.0
46.7
72.7
69.4
63.1
51.4
70.4
64.0
56.1
42.8
Table 15: Accuracy (s.d.) % on CIFAR10 with $\alpha=1$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 63.2
46.6
25.8
12.9
60.8
51.1
37.4
18.2
53.9
33.9
22.1
15.9
46.9
15.4
16.4
15.2
50.0
34.7
23.3
13.9
55.0
36.3
25.0
15.2
46.9
15.4
16.4
12.9
BKT 65.5
55.6
25.8
12.9
64.7
62.3
37.4
18.2
60.9
52.2
22.1
15.9
58.7
37.1
16.4
15.2
63.0
53.4
23.3
13.9
62.6
52.1
25.0
15.2
58.7
37.1
16.4
12.9
FoundFL 48.4
39.2
40.3
29.0
43.6
20.5
12.7
14.1
28.6
30.5
37.1
20.4
26.1
16.7
33.2
32.8
28.0
15.6
11.3
11.3
34.9
24.5
26.9
21.5
26.1
15.6
11.3
11.3
NNM 71.1
61.3
27.1
50.4
71.4
69.3
64.6
25.6
70.7
65.2
47.9
37.4
60.7
25.1
13.0
16.3
70.9
65.0
37.6
31.1
68.9
57.2
38.0
32.2
60.7
25.1
13.0
16.3
WoLA 68.3
55.6
38.1
25.1
67.1
61.2
54.6
41.5
67.6
62.9
53.1
30.7
66.0
59.6
50.9
41.8
63.8
55.8
47.0
26.9
66.5
59.0
48.8
33.2
63.8
55.6
38.1
25.1
WoLA+NNM 71.1
66.4
48.7
64.4
71.0
69.3
65.0
28.4
71.5
71.4
70.7
70.1
66.9
59.6
53.1
43.1
70.5
65.7
56.6
48.2
70.2
66.5
58.8
50.9
66.9
59.6
48.7
28.4
WoLA 68.1
54.4
31.2
12.8
66.8
60.1
51.6
41.5
67.3
61.9
48.1
24.4
65.3
57.5
45.9
31.4
63.1
54.9
45.3
24.9
66.1
57.7
44.4
27.0
63.1
54.4
31.2
12.8
WoLA+NNM 70.9
65.8
44.0
36.9
70.6
68.5
61.6
41.8
71.3
70.8
69.2
66.1
66.5
59.6
46.0
31.7
70.4
64.9
54.0
37.5
69.9
65.9
55.0
42.8
66.5
59.6
44.0
31.7
CwTM - 70.3
56.8
34.0
12.2
69.7
64.2
47.6
18.7
67.6
51.2
25.3
16.7
54.9
20.0
15.7
15.1
68.4
47.5
26.4
14.4
66.2
47.9
29.8
15.4
54.9
20.0
15.7
12.2
BKT 65.3
55.7
34.0
12.2
64.5
62.4
47.6
18.7
60.7
52.2
25.3
16.7
59.5
38.9
15.7
15.1
63.1
53.4
26.4
14.4
62.6
52.5
29.8
15.4
59.5
38.9
15.7
12.2
FoundFL 19.1
26.3
39.8
29.0
19.5
28.6
45.7
12.5
18.6
25.9
31.1
23.0
19.3
24.3
23.1
26.7
22.8
8.7
11.3
11.3
19.8
22.8
30.2
20.5
18.6
8.7
11.3
11.3
NNM 71.4
63.1
27.6
51.1
71.6
69.9
65.3
25.8
71.2
66.6
56.7
39.1
55.1
36.8
13.1
15.4
71.1
66.5
50.4
30.9
68.1
60.6
42.6
32.5
55.1
36.8
13.1
15.4
WoLA 70.8
64.7
43.9
24.7
70.5
67.9
58.2
41.2
69.7
64.3
54.0
31.5
66.4
60.6
53.2
43.7
69.9
60.6
49.0
26.6
69.5
63.6
51.6
33.5
66.4
60.6
43.9
24.7
WoLA+NNM 71.1
67.0
48.8
64.1
71.0
69.5
65.1
31.9
71.4
71.4
70.8
70.1
66.9
59.6
52.3
44.5
70.7
67.1
59.2
48.8
70.2
66.9
59.2
51.9
66.9
59.6
48.8
31.9
WoLA 70.8
64.1
36.6
13.6
70.4
67.2
54.1
42.0
69.6
63.5
49.8
24.5
66.4
59.7
45.7
29.6
69.7
59.0
46.4
27.9
69.4
62.7
46.5
27.5
66.4
59.0
36.6
13.6
WoLA+NNM 71.0
66.4
45.4
37.3
70.8
68.6
62.5
37.3
71.4
70.9
69.2
66.2
65.8
59.5
44.6
30.7
70.5
66.2
56.1
39.2
69.9
66.4
55.6
42.1
65.8
59.5
44.6
30.7
MKrum - 70.8
59.7
25.4
10.1
71.0
69.2
64.7
24.3
70.4
57.3
40.4
15.5
61.7
26.0
19.8
16.3
69.5
64.2
19.3
8.5
68.7
55.3
33.9
14.9
61.7
26.0
19.3
8.5
BKT 61.8
51.0
25.4
10.1
59.8
57.4
64.7
24.3
58.1
41.5
40.4
15.5
53.8
18.6
19.8
16.3
58.0
43.3
19.3
8.5
58.3
42.4
33.9
14.9
53.8
18.6
19.3
8.5
FoundFL 27.7
18.1
33.3
20.8
21.5
16.3
35.0
13.9
23.4
24.6
16.3
18.3
26.9
19.0
25.8
17.6
32.9
11.6
10.0
11.3
26.5
17.9
24.1
16.4
21.5
11.6
10.0
11.3
NNM 71.2
62.5
27.8
11.1
71.5
69.3
63.8
32.1
71.3
66.1
44.1
17.9
61.2
29.8
18.4
14.5
71.0
65.3
48.0
8.4
69.2
58.6
40.4
16.8
61.2
29.8
18.4
8.4
WoLA 70.8
66.0
48.8
25.2
70.4
68.4
63.0
30.2
71.6
71.1
61.8
14.8
67.0
60.4
52.6
43.1
69.9
63.8
29.8
8.4
69.9
66.0
51.2
24.3
67.0
60.4
29.8
8.4
WoLA+NNM 71.2
66.8
49.0
27.8
70.9
69.5
64.4
25.7
71.5
71.4
70.7
18.0
67.1
60.9
51.4
40.4
70.7
66.5
56.8
8.4
70.3
67.0
58.5
24.0
67.1
60.9
49.0
8.4
WoLA 70.5
65.6
42.8
12.4
70.3
67.8
60.9
25.7
71.3
70.4
50.2
12.8
66.5
58.3
46.2
31.7
69.6
63.4
25.4
10.1
69.6
65.1
45.1
18.5
66.5
58.3
25.4
10.1
WoLA+NNM 71.0
66.2
42.8
14.7
70.7
68.6
61.9
36.0
71.3
70.8
69.3
15.9
66.6
58.9
46.5
33.0
70.3
65.6
53.9
10.2
70.0
66.0
54.9
22.0
66.6
58.9
42.8
10.2
GM - 67.4
54.1
38.9
13.1
66.1
56.7
45.9
25.3
71.2
64.6
52.9
10.0
57.8
27.2
13.7
15.0
66.1
61.4
49.5
31.0
65.7
52.8
40.2
18.9
57.8
27.2
13.7
10.0
BKT 68.3
61.1
38.9
13.1
67.9
68.1
45.9
25.3
65.3
59.6
52.9
10.0
61.2
41.6
13.7
15.0
66.6
62.2
49.5
31.0
65.8
58.5
40.2
18.9
61.2
41.6
13.7
10.0
FoundFL 20.3
16.8
31.1
27.7
25.8
17.0
11.8
14.2
20.2
22.6
18.1
18.4
23.7
16.1
25.7
21.2
47.6
12.1
11.3
11.3
27.5
16.9
19.6
18.6
20.2
12.1
11.3
11.3
NNM 71.4
63.1
29.5
54.5
71.6
70.1
65.6
41.6
71.2
66.9
47.5
28.5
58.2
33.6
16.2
15.7
71.2
66.0
47.8
35.9
68.7
59.9
41.3
35.2
58.2
33.6
16.2
15.7
WoLA 69.5
62.3
47.0
28.5
68.9
65.5
61.2
44.9
71.4
70.1
66.4
38.0
67.2
60.2
48.2
48.3
66.8
63.1
57.0
30.8
68.7
64.2
56.0
38.1
66.8
60.2
47.0
28.5
WoLA+NNM 71.3
67.2
50.2
60.7
71.0
69.7
66.5
40.5
71.6
71.3
70.3
69.3
67.4
61.3
55.6
47.9
70.8
67.5
56.7
46.8
70.4
67.4
59.9
53.0
67.4
61.3
50.2
40.5
WoLA 69.5
61.4
42.1
16.7
68.8
64.2
57.4
36.9
71.2
69.1
63.8
27.9
66.9
58.1
40.8
33.2
66.3
61.0
53.5
31.7
68.5
62.8
51.5
29.3
66.3
58.1
40.8
16.7
WoLA+NNM 71.1
66.4
45.5
33.3
70.9
69.1
63.7
43.2
71.4
70.6
68.7
65.1
66.2
59.8
46.8
36.2
70.5
66.5
55.2
37.6
70.0
66.5
56.0
43.1
66.2
59.8
45.5
33.3
Table 16: Accuracy (s.d.) % on CIFAR10 with $\alpha=0.3$ and $f=2,4,6,8$. Colors rank values within sub-columns: yellow (1st), blue (2nd), and red (3rd).
Agg. Other Strategies ALIE FOE LF Mimic SF Average of Attacks Worst Case
CWMed - 44.0
32.9
15.8
10.9
43.5
40.1
21.9
11.7
25.4
21.8
22.7
11.7
15.7
12.3
16.0
13.6
29.3
20.8
16.4
11.2
31.6
25.6
18.6
11.8
15.7
12.3
15.8
10.9
BKT 57.1
37.2
15.8
10.9
56.4
47.3
21.9
11.7
51.9
35.9
22.7
11.7
45.5
19.5
16.0
13.6
54.2
28.8
16.4
11.2
53.0
33.7
18.6
11.8
45.5
19.5
15.8
10.9
FoundFL 26.6
29.7
25.9
15.8
23.6
17.2
14.0
13.5
24.6
26.4
19.5
19.2
17.8
21.3
15.2
15.5
19.2
16.7
10.3
9.2
22.4
22.3
17.0
14.6
17.8
16.7
10.3
9.2
NNM 67.1
43.8
16.9
49.3
66.6
62.5
57.6
27.0
64.8
54.7
37.0
15.4
45.0
17.7
20.7
13.6
66.3
56.6
28.3
14.8
62.0
47.1
32.1
24.0
45.0
17.7
16.9
13.6
WoLA 61.3
43.1
31.3
18.8
61.3
53.9
42.7
23.3
60.0
51.2
31.6
19.1
54.0
43.3
33.2
28.3
52.9
45.1
30.4
16.6
57.9
47.3
33.8
21.2
52.9
43.1
30.4
16.6
WoLA+NNM 66.4
58.3
38.7
46.0
66.4
64.6
56.6
23.9
67.9
67.5
66.8
57.1
58.4
50.1
42.6
29.2
65.8
57.4
37.3
36.3
65.0
59.6
48.4
38.5
58.4
50.1
37.3
23.9
WoLA 61.4
41.3
19.8
9.0
60.9
51.7
44.3
29.2
60.0
49.1
27.2
15.6
54.0
39.6
29.6
19.8
53.5
42.2
35.3
19.6
58.0
44.8
31.2
18.6
53.5
39.6
19.8
9.0
WoLA+NNM 66.0
57.3
29.3
17.0
66.2
62.5
52.9
25.0
67.6
66.3
63.5
31.0
58.3
42.9
27.8
24.9
65.6
56.5
30.7
20.3
64.7
57.1
40.9
23.7
58.3
42.9
27.8
17.0
CwTM - 63.8
40.0
20.9
10.8
62.8
48.8
33.7
11.4
59.5
32.4
22.3
11.8
37.9
18.2
13.9
14.2
60.0
21.6
14.2
11.2
56.8
32.2
21.0
11.9
37.9
18.2
13.9
10.8
BKT 57.1
37.9
20.9
10.8
56.6
47.0
33.7
11.4
52.0
36.1
22.3
11.8
41.8
17.6
13.9
14.2
54.2
28.3
14.2
11.2
52.3
33.4
21.0
11.9
41.8
17.6
13.9
10.8
FoundFL 17.9
20.6
23.8
23.3
17.8
18.8
24.0
9.8
18.7
18.6
19.1
16.9
17.0
20.0
16.6
20.6
18.1
9.9
13.5
9.2
17.9
17.6
19.4
16.0
17.0
9.9
13.5
9.2
NNM 67.7
50.2
16.7
48.4
68.0
63.7
57.6
27.4
66.5
56.6
39.5
15.6
41.6
22.6
16.4
14.4
67.4
59.5
36.6
14.6
62.3
50.5
33.4
24.1
41.6
22.6
16.4
14.4
WoLA 66.4
55.1
34.8
18.9
66.3
62.8
50.8
23.6
64.8
56.5
37.2
19.2
58.6
46.5
36.1
30.7
65.6
48.1
36.5
16.2
64.3
53.8
39.1
21.7
58.6
46.5
34.8
16.2
WoLA+NNM 67.0
58.9
38.7
43.0
66.9
65.1
59.0
24.8
67.8
67.6
66.8
56.8
59.5
48.3
41.5
31.3
66.4
61.4
49.0
34.5
65.5
60.3
51.0
38.1
59.5
48.3
38.7
24.8
WoLA 66.3
53.7
24.4
8.1
66.1
61.1
45.5
30.0
64.2
54.6
31.0
16.1
57.7
45.8
29.4
18.2
65.3
47.8
39.2
18.5
63.9
52.6
33.9
18.2
57.7
45.8
24.4
8.1
WoLA+NNM 66.5
58.1
30.7
15.3
66.5
63.5
53.3
25.1
67.5
66.3
63.6
30.7
58.9
48.6
26.5
19.6
65.9
59.1
44.9
19.2
65.1
59.1
43.8
22.0
58.9
48.6
26.5
15.3
MKrum - 65.7
42.4
17.4
5.0
66.3
61.3
57.2
22.6
61.9
53.8
37.9
22.8
46.3
22.5
18.0
13.6
64.5
50.8
13.2
7.7
60.9
46.2
28.7
14.3
46.3
22.5
13.2
5.0
BKT 53.9
37.6
17.4
5.0
53.2
48.8
57.2
22.6
48.7
38.2
37.9
22.8
35.5
13.4
18.0
13.6
51.0
35.8
13.2
7.7
48.5
34.8
28.7
14.3
35.5
13.4
13.2
5.0
FoundFL 16.8
18.8
21.5
18.4
19.2
21.4
26.5
12.0
19.1
16.8
17.8
12.3
18.7
18.9
16.4
17.9
21.3
10.0
13.4
9.2
19.0
17.2
19.1
14.0
16.8
10.0
13.4
9.2
NNM 67.5
49.4
18.1
5.9
67.7
63.0
54.2
27.2
65.8
54.8
37.7
15.3
49.4
16.0
16.0
12.4
67.2
58.3
34.2
9.0
63.5
48.3
32.0
14.0
49.4
16.0
16.0
5.9
WoLA 65.9
57.5
38.7
19.6
65.5
63.8
56.5
23.9
67.8
67.4
38.5
12.2
57.1
49.1
37.9
29.2
65.0
56.3
13.8
7.5
64.3
58.9
37.1
18.5
57.1
49.1
13.8
7.5
WoLA+NNM 66.8
58.1
38.9
22.2
66.6
64.6
55.1
23.4
67.7
67.7
66.9
14.2
59.7
49.6
41.2
33.2
66.4
60.0
35.2
7.5
65.4
60.0
47.4
20.1
59.7
49.6
35.2
7.5
WoLA 65.7
56.8
28.6
7.3
65.9
62.2
52.9
25.0
67.6
65.9
26.1
11.2
57.4
47.7
29.4
24.9
64.7
54.4
17.2
9.5
64.3
57.4
30.9
15.6
57.4
47.7
17.2
7.3
WoLA+NNM 66.4
57.4
30.0
6.9
66.4
63.4
52.1
25.2
67.6
66.3
63.2
11.0
58.9
44.8
29.9
21.8
65.9
58.2
37.8
9.4
65.0
58.0
42.6
14.9
58.9
44.8
29.9
6.9
GM - 58.1
39.8
28.1
6.1
60.6
47.9
31.0
10.5
66.0
56.5
31.1
8.2
45.5
17.7
18.6
13.5
61.0
52.5
39.2
12.2
58.2
42.9
29.6
10.1
45.5
17.7
18.6
6.1
BKT 61.7
48.2
28.1
6.1
61.4
62.0
31.0
10.5
56.7
54.4
31.1
8.2
46.1
21.5
18.6
13.5
60.1
54.5
39.2
12.2
57.2
48.1
29.6
10.1
46.1
21.5
18.6
6.1
FoundFL 16.6
16.8
21.9
21.8
22.9
14.9
14.1
10.9
15.9
17.4
19.3
16.4
16.7
16.6
19.0
18.9
32.0
13.7
13.5
13.1
20.8
15.9
17.6
16.2
15.9
13.7
13.5
10.9
NNM 67.5
48.1
20.6
53.1
67.7
63.7
58.4
35.7
66.4
56.6
38.9
16.8
46.8
15.1
13.4
13.3
67.3
60.1
32.1
19.8
63.1
48.7
32.7
27.7
46.8
15.1
13.4
13.3
WoLA 63.5
51.7
38.5
23.1
64.6
60.3
53.7
18.0
66.9
64.1
51.0
17.9
59.5
48.6
34.7
32.6
62.3
56.3
50.2
7.4
63.3
56.2
45.6
19.8
59.5
48.6
34.7
7.4
WoLA+NNM 66.9
58.8
39.9
46.8
67.0
65.0
59.3
31.7
67.9
67.5
66.1
54.7
59.3
50.7
42.9
35.1
66.4
59.9
37.3
32.4
65.5
60.4
49.1
40.2
59.3
50.7
37.3
31.7
WoLA 63.1
49.9
29.2
9.3
63.9
58.4
50.3
10.8
66.7
62.4
46.1
11.2
58.6
47.0
24.5
23.0
61.4
55.4
47.1
9.3
62.7
54.6
39.4
12.7
58.6
47.0
24.5
9.3
WoLA+NNM 66.4
58.0
30.8
15.7
66.4
63.7
54.3
27.6
67.6
66.1
62.6
38.1
58.7
46.0
35.0
20.1
66.1
58.4
35.7
21.0
65.0
58.4
43.7
24.5
58.7
46.0
30.8
15.7